### v1

# Decentralization of Ethereum’s Builder Market

Abstract.
Blockchains protect an ecosystem worth more than \$500bn with their strong security properties derived from the principle of decentralization. Is today’s blockchain really decentralized? In this paper, we empirically studied one of the least decentralized parts of Ethereum—the most used blockchain system in practice—and shed light on the decentralization issue from a new perspective.
To avoid centralization caused by Maximal Extractable Value (MEV), Ethereum adopts a novel mechanism that produces blocks through a builder market. After two years in operation, however, the builder market has evolved to a highly centralized one with three builders producing more than 90% of blocks. Why does the builder market centralize, given that it is permissionless and anyone can join? Moreover, what are the security implications of a centralized builder market to MEV-Boost auctions? Through a rigorous empirical study of the builder market’s core mechanism, MEV-Boost auctions, we answered these two questions using a large-scale auction dataset we curated since 2022.
Unlike previous works that focus on who wins the auctions, we focus on why they win, to shed light on the openness, competitiveness, and efficiency of MEV-Boost auctions. We show the access barriers around private order flows constitute a significant entry barrier to the builder market. A new builder needs to pay up to 1.4 ETH to access this private order flow, a significant amount considering their low initial profits. By computing the true value of bids, we investigated the inequality in block-building capabilities among builders and the competitiveness and efficiency of the MEV-Boost auctions. We observed that top, middle, and tail builders have significantly varying capabilities in extracting MEV, and the inequality worsens when MEV increases. 88% of the MEV-Boost auctions we studied were competitive, yet uncompetitive auctions still resulted in a total loss of 221.09 ETH for the proposers, which is 0.98% of their total gain. Meanwhile, only about 79% of the MEV-Boost auctions were efficient, and over half of the inefficient auctions were caused by block subsidization. Our findings also help identify directions for improving the decentralization of builder markets.
1. Introduction
At the core of any decentralized blockchain, a key assumption underpinning their strong security properties (integrity, availability, censorship resistance, incentive compatibility, etc) is “decentralization”, that the system is run by a large number of independent participants (Nakamoto, 2008; Eyal and Sirer, 2018; Schwarz-Schilling et al., 2022; Wahrstätter et al., 2023a; Bahrani et al., 2024; Roughgarden, 2020). It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) (Daian et al., 2020), profits that protocol participants, hereafter referred to as validators, can reap by manipulating the ordering of transactions. MEV can lead to centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have (Bahrani et al., 2024; Ethereum Foundation, 2024b; Pai and Resnick, 2023; Buterin, 2021b, a).
In an effort to level the playground for validators, the Ethereum community proposed Proposer [^1] A proposer is a validator chosen as the leader in a round of consensus to add a new block to the blockchain.-Builder Separation (PBS) (Ethereum Foundation, 2024b). The idea is to outsource the task of building profitable blocks to specialized entities called block builders. In the current implementation called MEV-Boost (Flashbots, 2024), the proposer runs an auction among builders and adds the block with the highest bid to the blockchain. MEV-Boost is permissionless for builders so any builder can join the auction. Ideally, builders compete with each other in auctions, and the auctioneer—the proposer—collects a significant portion of MEV in auction revenue without doing much work, eliminating the competitive advantage of sophisticated validators over regular ones.
Figure 1. Market share of builders. The builder market is arguably one of the least decentralized parts of Ethereum!
Does PBS achieve its goal? The answer is rather unclear. As shown in Fig. 1, the current market share is highly concentrated, with over 90% of the blocks produced through MEV-Boost being built by just three builders. However, the current concentration does not necessarily defeat the purpose of PBS for two reasons: first, builder markets may evolve and new builders may join and disrupt the concentration; second, centralization does not imply a lack of competition—it only takes a few non-colluding bidders to make the auction competitive and efficient.
Thus, while previous works (Flashbots, 2024; Wahrstätter, 2024; Labs, 2024; Heimbach et al., 2023; Wahrstätter et al., 2023b; Yang et al., 2022; Thiery, 2023) point out the centralization issue, they did not answer two key questions: First, how open is the builder market to new builders? Second, despite the centralization, are MEV auctions actually competitive and efficient? (We leave technical definitions to future sections, but intuitively, an auction is competitive if the winner outbids the second-best bidder; an auction is efficient if the bidder with the highest valuation wins.)
In this paper, we answer these questions through a rigorous empirical study of historical MEV auctions. The key ingredient of our study is a large-scale auction dataset we curated since 2022: we collected 5.57 billion partial bids [^2] Partial and full bids are to be defined in Sec. 3.1 of 3,810,630 auctions from September 2022 to March 2024; in addition, in collaboration with ultra sound relay (Money, 2023), we obtained full bids for a total of 147,926 auctions from April 2023 to August 2023.
Auction data gives us direct visibility into the openness, competitiveness, and efficiency of MEV-Boost auctions. For example, historical bids allow us to “simulate” past auctions assuming a different set of transactions were available to builders, allowing us to discern the importance of different transactions. Another example is that we can compute the true value of each bid (i.e., the actual worth of the block to a builder, which is usually higher than what they bid) from transactions. The competitiveness and efficiency of an auction follow directly from a comparison between the true values and bid values. This is in contrast to, say, eBay auctions, where the true value of an item is subjective and cannot be retroactively computed from data.
Now we present an overview of methods and findings.
Table 1. Summary of findings
Finding 1: Private order flows contribute to 60% of the MEV in over 50% of the blocks in MEV-Boost.
Finding 2: To access private order flows, new builders need to pay up to 1.4 ETH to enter the market.
Finding 3: The inequality in block building is highest among tail builders, followed by middle and least among top builders, and it worsens when MEV increases.
Finding 4: 88.84% of the MEV-Boost auctions are competitive, with proposers’ losses from uncompetitive auctions amounting to 0.98% of total gains.
Finding 5: 79.74% of the MEV-Boost auctions are efficient, and 51.4% of the inefficient auctions are caused by block subsidization.
1.1. Openness of the Builder Market
First, we investigate the reasons behind centralization by quantifying the market entry barrier for a new builder. One such barrier could be the ability to build effective MEV extraction algorithms, but we investigated a perhaps more basic barrier: the ability to access profitable transactions. Recall that MEV extraction is about manipulating the ordering of transactions, so having access to “extractable” transactions is the basis of any extraction algorithm.
As we will present in Sec. 4, private order flows (Kilbourn, 2022)—the transactions that bypass the public mempool—contribute an increasing portion to builder income. This means that it would be difficult for a builder to win the auction without access to private order flows. Moreover, not all private order flows are equally profitable for builders, so the ability to access “good” private order flows forms a barrier to entering the builder market.
Using the auction dataset, we identified the providers of profitable transactions, quantified their importance to the builders, and investigated their accessibility to new builders.
Identifying pivotal private order flow providers.   In today’s MEV ecosystem (to be presented in Sec. 2.2), private order flows come from various private order flow providers. Different providers impose different access barriers to their private order flows. Some providers (such as MEV-Share (Flashbots, 2024) and MEV Blocker (MEV Blocker, 2023)) have relatively clear criteria a builder must satisfy to access their flows. Still, some providers (e.g., most searchers) are almost completely opaque to the public. To identify important providers, we define the pivotal level of a provider $P$ as the percentage of MEV-Boost auctions in which the winner would have lost had transactions from $P$ were not available. As we will show in Sec. 4.1, we identified five pivotal providers whose pivotal level exceeds 50% over a period longer than two weeks. That is, if a builder cannot access the flows from any of these providers, it will lose the majority of the auctions during the period of our study.
The good news is that two of these pivotal providers are MEV Blocker (MEV Blocker, 2023) and MEV-Share (Flashbots, 2024), which are open to builders who meet a relatively clear set of criteria. The bad news is that satisfying the requirements set by MEV Blocker and MEV-Share is non-trivial for new builders, which we discuss next.
Quantifying the market entrance barrier.   Typically, providers share private order flows to reputable builders with a certain market share (Sec. 4.2). However, this presents a chicken-and-egg problem for new builders: to gain market shares, they need private order flows to win the auction, but starting out new, they do not have market share and thus cannot access private order flows. To navigate this dilemma, new builders resort to subsidization (Flashbots, 2024): bidding higher than their actual profits in MEV-Boost auctions, to establish the required market shares. Therefore a new builder needs to have enough capital to keep subsidizing until reaching a certain market share. To understand the significance of this entry barrier, we compute the minimal cost of subsidization required to maintain a 1% market share, which is the requirement of MEV Blocker. Specifically, we first calculate the minimal subsidy a builder needs to win an MEV-Boost auction without private order flows; then, we sum up the minimum subsidies over 1% of total auctions per day as the daily cost.
Our analysis highlights a non-trivial cost for new builders to establish a reputation through subsidization—they need to invest up to 1.4 ETH to subsidize their blocks and secure a minimal 1% market share during the period of our study. The cost is comparable to or even higher than the profits of small builders (Flashbots, 2024).
Trust barriers result in access barriers.   An important takeaway is that siloes of private order flows stem from the lack of a fair exchange mechanism between providers and builders. In the current design, builders are at an advantage and malicious builders can harm providers by, e.g., sandwich attacks (Zhou et al., 2021) and imitation attacks (Qin et al., 2023). Therefore, providers only share their order flows with reputable builders, even though such reputation-based protection is pretty weak. As entry barriers arise from the trust crisis, reducing or eliminating these trust barriers is essential for making the builder market more decentralized. We discuss several ideas in Sec. 7.2.
1.2. Competitiveness and Efficiency
The first question focuses on the evolution of builder markets. In the second question, we focus on the status quo and study the implications of a centralized builder market as we have today.
MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction (Besanko et al., 2009), with a 12-second deadline. English auctions have appealing features: ideally, assuming bidders do not collude and can react to each other instantaneously, the bidders who can extract most MEV will win (efficiency), paying at least the MEV extracted by the second-highest bidder (competitiveness).
However, given the small number of builders in today’s builder market, it is unclear if MEV-Boost auctions indeed induce the desired dynamics. We investigate whether the apparent centralization of the builder market indeed leads to inadequate competition and a loss of efficiency.
True values in MEV-Boost auctions.   In MEV-Boost auctions, for a block $B$, a bid value (BV) is the amount the builder is willing to pay for $B$ to be added to the blockchain. The total revenue gained by the bidder from $B$ is the true value (TV). Rational bidders would bid lower than the true value to get a non-zero profit. The true value of a block $B$ can computed from transactions in $B$.
Inequality in block building capabilities.   We first compare the true values across different builders for the same auction to understand the variations in builders’ abilities to extract values. A builder who consistently demonstrates a superior ability to extract value could dominate the market. We used the quartile coefficient of dispersion (QCD) (Wikipedia, 2022) to quantify the inequality of true values across builders. Our results in Sec. 5.1 reveal significant differences in the block-building capabilities of top, middle, and tail builders. The inequality worsens as the MEV of a slot increases, possibly because significant MEV opportunities are more likely to be shared with only a limited number of builders. Fortunately, the inequality among top builders is not significant, which supports the premise of competitiveness and efficiency.
More than 88% of auctions are competitive.   Using the true values, we propose two metrics, competitive index (CI) and efficient index (EI), to quantify whether an auction is competitive or efficient, respectively (see Sec. 5.2 for details.) Our results reveal that approximately 88.84% of the MEV-Boost auctions we analyzed were competitive. However, we also found that the auctions tend to be less competitive when the MEV of a slot increases. As discussed previously, the main concern is that a lack of competition can cause loss to proposers. We therefore quantify the loss for proposers. Our study shows that 16,498 of the 147,926 MEV-Boost auctions were uncompetitive, where proposers incurred total losses amounting to 221.09 ETH, which is 0.98% of the total gains of the proposers.
Only 79% of auctions are efficient.   An interesting finding is that more than 20% of MEV-Boost auctions are inefficient and the auctions tend to be inefficient when the MEV in an MEV-Boost auction is low. Further analysis reveals that over half of the inefficient auctions are caused by block subsidization. The winners of these inefficient auctions overbid to win, although they do not have the highest true value. Moreover, block subsidization is more common when the extractable value is small, because the cost will be relatively low. Although this may not benefit the overall welfare, it is advantageous for proposers as they earn more than expected.
Suggestions for future developments.   Since many identified issues stem from trust concerns, future developments should aim to design mechanisms that reduce or eliminate the reliance on trust. Possible directions include community supervision (Flashbots, 2023a) and the use of Trusted Execution Environments (TEEs) (Costan and Devadas, 2016) to ensure the integrity and confidentiality of private order flows (Flashbots, 2024c). Another idea is to implement a reliable leakage detection mechanism for private order flows, which we briefly discuss in Sec. 7.2. Such mechanisms would encourage private order flow providers to distribute their flows across multiple builders, thus increasing the equality of the builders’ ability to extract value.
Contribution
To the best of our knowledge, we are the first to use the rich auction data to analyze the decentralization of Ethereum’s builder market, while most previous works are limited to on-chain data. Our dataset of the true values of 147,926 MEV-Boost auctions from April to August 2023 may be of independent interest. Our dataset of 5.57 billion partial bids since 2022 is the largest-scale known to us.
We characterized the efficacy of PBS in terms of being open to new builders and inducing competitive and efficient auctions. Our key findings are highlighted in Tab. 1, which we summarize below.
  • Sec. 4: Private order flows contribute to more than 60% of the MEV in over 50% of blocks every day since July 2023. We identified five pivotal private order flow providers that had a significant and sustained impact on the winners of more than half of the MEV-Boost auctions. Two of them are MEV Blocker and MEV-Share.
  • Sec. 4.2: Private order flow providers impose market share requirements on builders, which costs new builders up to 1.4 ETH in subsidy to enter the market during the period of our study.
  • Sec. 5.1: The QCD shows a significant inequality in builders’ abilities to extract value across top, middle, and tail builders, but the inequality among the top builders is not significant.
  • Sec. 5.2: We find that 88.84% of the 147,926 MEV-Boost auctions are competitive. However, proposers incur losses of 221.09 ETH from the uncompetitive auctions, representing 0.98% of their total gain. These losses can still threaten the stable separation between proposers and builders.
  • Sec. 5.2: 79.74% of the MEV-Boost auctions in our study are efficient. A significant factor contributing to inefficiency is block subsidization, where winners overbid to secure the auctions, possibly driven by the need to meet market share requirements.
Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions. Currently, there is only limited auction theoretic analysis that considers the complexity of real-world MEV auctions (Gupta et al., 2023; Pai and Resnick, 2023; Wu et al., 2024). The findings in this paper can serve as empirical evidence for future theoretic analysis of MEV-Boost auctions.
Figure 2. Illustration of the MEV supply chain in the PBS scheme.
2. Background
2.1. MEV-Boost Auctions
MEV.   Miner/Maximal Extractable Value (MEV) (Daian et al., 2020) refers to the profit that privileged players (e.g., validators) can earn by including, excluding, and reordering the transactions. MEV poses a threat to decentralization, as explained in the introduction. There are two schools of thought on mitigating the negative impacts of MEV. The first one is referred to as MEV democratization (Flashbots, 2022b), the main idea of which is to facilitate MEV extraction and level the playground. MEV-Boost (Flashbots, 2024), to be introduced below, is the primary example. Another approach is to enforce certain ordering policies (e.g., FIFO) to prevent order manipulation, such as (Kelkar et al., 2020, 2023). We focus on MEV-Boost because it is widely used.
PoS Ethereum.   Before introducing background on MEV-Boost auctions, we first cover the relevant concepts in Ethereum’s Proof of Stake-based consensus protocol called Gasper (Buterin et al., 2020). Participants of Gasper are called validators. Each validator must deposit 32 ETH as collateral (Ethereum Foundation, 2024a). Time is divided into 12-second intervals called slots. This results in 7,200 slots in Ethereum per day. During each slot, a single validator, referred to as the proposer, proposes a block, while other validators vote for the proposed block if it extends the chain head in their local view.
Proposer-builder separation (PBS) and MEV-Boost.   Proposer-builder separation (PBS) can be conceptually viewed as a sidecar to the core consensus protocol (although some proposals are more tightly coupled with consensus) to allow the proposer to outsource the job of building blocks to builders.
MEV-Boost (Flashbots, 2024) is the current incarnation of the PBS idea. About 90% of all blocks on Ethereum are built by MEV-Boost (Wahrstätter, 2024). The basic idea is to auction off the right to propose a block to the builder who bids the highest in an open-bid ascending price auction (contributors, 2024). Because builders and proposers are mutually untrusted, the auction is facilitated by a (trusted) third party called a relay (Flashbots, 2022a). Here are the high-level steps (we refer readers to (Flashbots, 2024) for details):
  • Builders submit bids to relays throughout the slot. For the purpose of this paper, we denote a bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key. Bids are signed by the builder but we omit the signature for clarity. Relays verify that $B$ indeed pays the proposer at least $BV$.
  • Relays expose a public API that publishes partial bids in the form of $(\mathsf{Hash}(B),BV,metadata)$. $metadata$ includes parent hash, gas limit, gas used, etc. Partial bids do not include block content.
  • The proposer selects a block hash $h_{B}$, typically the one with the highest $BV$, and sends the relay a signature on $h_{B}$. This signature is a promise to propose $B$ in the consensus protocol (if it proposes another block, this signature is evidence of equivocation and will cause the proposer to be penalized (Buterin et al., 2020).) After verifying the signature, the relay returns $B$ to the proposer.
Two additional remarks. First, relays are trusted entities, but anyone can run a relay (currently there are ten (EthStaker Community, 2024)), and builders and proposers can connect to the relays they trust. Second, winning bids are recorded on-chain, and losing bids are discarded after the auction. We collected and archived the historical bids over time to enable this study, as we will detail in Sec. 3.
2.2. MEV Supply Chain
The ecosystem around discovering and extracting MEV forms an MEV supply chain, as depicted in Fig. 2. Builders, relays, and proposers (validators) have been introduced above. Now we introduce the “upstream” of builders.
Wallet.   Users use wallets to interact with blockchains. A user communicates her intent (e.g., transferring cryptocurrency) to the wallet, which generates the transaction to fulfill the user’s intent and sends it to the network. MEV may be generated during this process, e.g., when the user transaction creates an arbitrage opportunity. While the wallet is typically controlled by the users, it can also be managed by a third party, such as Telegram trading bots (Intelligence, 2024). These bots are automated programs integrated with the Telegram app (Telegram, 2024), automatically conducting cryptocurrency trades on behalf of users. Users’ transactions eventually reach builders or proposers, either through the peer-to-peer network (the mempool) or through third-party services called private channels.
Mempool.   Mempool is a temporary storage of pending transactions before they are added to the blockchain. All transactions stored in the mempool are publicly visible.
Channel.   A private channel, or channel for short, connects users directly with builders, bypassing the mempool. Channels receive transactions from users and then, forward them to the builders specified by the users, offering privacy and atomicity guarantees. E.g., a user can send a bundle (list) of transactions and request them not to be unbundled. Examples include MEV-Share (Flashbots, 2024), MEV Blocker (MEV Blocker, 2023), BloXroute (bloXroute Labs, 2024), etc.
Searcher.   Searchers are a special type of users. They run (typically proprietary) algorithms to construct profitable transaction sequences (bundles) based on the pending transactions and the blockchain state. Same as regular users, searchers may submit transactions to mempool or channels. In addition, most builders have public APIs that allow searchers to submit bundles directly. Some searchers may prefer channels for the privacy and atomicity benefits. A search is said to integrate with a builder if it sends almost all of its orders to that builder exclusively. For example, searcher 0xA69b… is considered an integrated searcher of Beaver Builder because Beaver Builder includes the majority of its transactions.
Some channels, such as MEV-Share and MEV Blocker, allow searchers to read pending transactions in the channel and extract MEV from them, provided that such extraction should not harm users (e.g., only backrunning is allowed).
Private order flow.   Order flow refers to a source of transactions that may carry extractable value and can be categorized into two categories: public order flow, originating from the public mempool, and private order flow, sent directly by users or searchers without being broadcast to the public mempool. The entities that can provide private order flows to the builders are referred to as private order flow providers, which includes searchers, channels, users, and Telegram trading bots as shown in Fig. 2.
2.3. Model and Assumptions.
Accounts and states.   We follow the model of (Babel et al., 2023a, b). We use $\mathcal{A}$ to denote the space of all possible accounts. For example, in Ethereum, both user-owned and contract-owned accounts are represented by 160-bit identifiers. For $a\in\mathcal{A}$, $\text{balance}(a)$ denotes the balance of all tokens held in account $a$, and $\text{balance}(a)[\mathbf{T}]$ specifies the balance of token $\mathbf{T}$. For simplicity, $\text{balance}(a)[0]$ is used to denote the balance of the primary token (e.g., ETH in Ethereum).
Transaction and block.   A transaction $tx$ is a string whose execution changes the system’s state. A block is an ordered sequence of transactions. Executing transactions in a block $B$ transforms an initial state $s$ into a new state $s^{\prime}$, represented by $s^{\prime}=\text{action}(B)(s)$.
Extracted value (EV).   The extracted value of a block $B$ in state $s$ for a participant $p$ is the changes of $p$’s balances after executing $B$. Assuming $p$ controls a set of accounts $A_{p}$, the extracted value of $B$ for $p$ is:
(1) $EV_{p}(B)=\sum_{a\in A_{p}}\text{balance}_{s^{\prime}}(a)[0]-\text{balance}_{s}(a)[0]$
where $s^{\prime}=action(B)(s)$.
Builder’s income.   We assume a builder $p$’s income from building a block $B$ is $EV_{p}(B)$, or $EV(B)$ for short since the builder identity is usually encoded in the block. We compute EV following a similar approach as that in MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023). Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees (Ethereum, 2024) and direct transfers (Flashbots, 2024a) to the builder’s addresses minus the refund to other addresses. We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
Note that $EV(B)$ is a lower bound of the builder’s actual income from $B$ because the builder may own accounts out of $A_{p}$ in Equation 1, or even receive payments off-chain. We discuss the implications in Sec. 7.
3. Dataset
We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to March 2024, and a dataset of full bids from April to August 2023. The full bids are obtained from the ultra sound relay (Money, 2023). Recall from Sec. 2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc.
To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. 2 summarizes the datasets.
Dataset Content # of entries Source
On-chain blocks 4,019,895 Reth node
transactions 606,484,921
Auction blocks produced by MEV-Boost 3,517,851 relay API
partial bids 5,577,318,196 relay API
full bids from ultra sound 164,700,566 ultra sound relay
Misc. builder public keys 191 see Sec. 3.1
builder addresses 125
searcher addresses 2,069
private transactions 49,081,209
transaction sources 33,010,466
Table 2. Overview of collected datasets.
3.1. Data Collection and Validation
On-chain data.   We run an Ethereum execution client Reth (paradigmxyz, 2024) along with a consensus client Lighthouse (Prime, 2024) to collect blocks and transactions from September 15, 2022, to March 31, 2024. The on-chain dataset includes 4,019,895 blocks and 606,484,921 transactions.
Data from relay APIs.   We connected to all 13 known relays333three relays were inactive during the period of our study and collected the following information from their public APIs: 1) the winning bid including the final block sent to the proposer; and 2) partial bids for blocks that did not win the auction. Recall that partial bids do not include the block body. We collected 3,517,851 winning bids and 5,577,318,196 partial bids from September 15, 2022, to March 31, 2024, in a total of 3,810,630 auctions. We use winning bids to identify blocks produced through MEV-Boost. We use partial bids to quantify the importance of private order flow providers in Sec. 4.
To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets (Network, 2023; Wahrstätter, 2023; DataAlways, 2024). Our dataset includes 5,382 more winning bids than other datasets combined. Upon comparison with on-chain blocks, we find they are all included on Ethereum, thereby indicating a high level of completeness.
Full bids from the ultra sound relay .   From the ultra sound relay (Money, 2023), we obtained the full bids in 147,926 historical auctions that took place between April and August 2023. Since the data volume is large (we record roughly 200 GB data per day), our dataset samples one week of each month: April 9-15, May 1-7, June 1-7, July 1-7, and August 1-7, 2023. The dataset consists of 164,700,566 full bids (block, bid value, and builder public key). We use full bids in Sec. 5 to quantify the competitiveness and efficacy of MEV-Boost auctions.
The original data from relays includes bids submitted after the deadline because of network latency. We removed them as we are only interested in the bids before the auction ends. On April 18, 2023, MEV-Boost introduced bid cancellation that allows a bidder to cancel a previous bid by submitting a lower one. However, proposers can query bids before they are canceled and store them locally (Ethereum Research, 2023), effectively bypassing cancellation. Therefore we do not take cancellation into consideration.
Not all builders connect to the ultra sound relay, so we need to validate that the full bids from ultra sound relay cover a significant subset of all builders. We compute the percentage of the builders that submit to ultra sound relay over all known builders. The set of all known builders is a union of all builders that appeared in our datasets, including winning bids, partial bids, and bids from the ultra sound relay. As shown in Fig. 10, bids from the ultra sound relay covered more than 80% of the builders in over 75% of the MEV-Boost auctions during our study period. Moreover, the top-5 builders by market share (Flashbots, Beaver, Rsync, Builder0x69, and Titan) almost always submitted to ultra sound relay.
Miscellaneous data.   In practice, builders are known by their common names (e.g., flashbots builder), but a builder may use multiple public keys to sign bids and use multiple addresses to receive and make payments. We collect builders’ public keys to attribute bids to builders. We collect builder’s addresses to calculate the true value of bids. We use private transaction hashes to identify private order flows and use the searchers’ addresses and sources of transactions to identify the private order flow providers in Sec. 4.
  • Builder public keys: We collect public keys controlled by each builder combining two data sources. First, some builders publish their public keys in official documents or code repositories. For example, Titan builder has disclosed 12 public keys in its official document (Builder, 2024). Second, some builders leave distinct marks in their blocks’ “extra field”. For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’. Given that such marks are unauthenticated and susceptible to impersonation, we ignore them unless they are used more than 100 times in the bids by the same builder public key.
    To validate the completeness of our dataset, we manually cross-check it with public datasets (Rated Network, 2023; Network, 2024) and related works (Heimbach et al., 2024, 2023), and the results show that our dataset includes 62 more public keys, so our dataset is the most complete one so far.
  • Builder addresses: We collect the builder addresses from the blocks’ last transactions combined with the public dataset (Etherscan, 2024). Similarly, to validate their completeness, we manually cross-check them with related works (Heimbach et al., 2024, 2023), and the results show that our dataset includes 113 more builder addresses, which is the most complete so far.
  • Searcher addresses: We obtained known searcher addresses from Etherscan (Etherscan, 2024), libMEV (libMEV, 2024), and related work (Heimbach et al., 2024).
  • Private transaction hashes: We use data from Mempool Guru (Yang et al., 2022) to label transactions that bypassed the public mempool as private. We believe these labels have already been cross-validated since the private transactions are collected by seven distributed nodes.
  • Transaction sources: we are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through.
    We identified transactions originated from Telegram trading bots, including Maestro (Maestro, 2023a), Banana Gun (Banana Gun, 2023), and Unibot (Unibot X, 2023), by searching for transactions that interacted with their official smart contracts.
    Two channels, MEV Blocker and MEV-Share publish data (Protocol, 2024; Flashbots, 2024b) about transactions submitted to them. Specifically, transactions submitted to MEV Blocker are publicly available on Dune (Protocol, 2024), from which we extracted 14,901,303 transaction hashes that fall within our study’s range. We obtained 13,992,371 transactions submitted to MEV-Share from the public Flashbots dataset (Flashbots, 2024b). These transaction labels are sourced from their official datasets and we view them as ground truth.
4. Openess of the builder market
The builder market is technically open to everyone, but the barrier to accessing profitable private order flows (Sec. 2.2) forms an implicit access barrier.
To illustrate the importance of private order flow to the success of builders, Fig. 3 plots the fraction of builder income from private order flows. Currently, approximately 60% of the block value comes from private order flows.
Figure 3. Fraction of builder income from private order flows.
Finding I: The private order flow is crucial for builders to win the MEV-Boost auction because it contributes to more than 60% of the block value in over 50% of blocks.
Where can new builders get private order flows? The entities who provide (or sell) private order flows to builders are called private order flow providers. Different providers impose different access barriers to their private order flows. Some providers are open to new builders who satisfy certain requirements (such as MEV Blocker or MEV-Share), while some (e.g., most searchers) are almost completely opaque to the public.
To understand if the builder market is open to new builders, we first identify important “pivotal” providers and then quantify the cost for a new builder to access private order flows from representative pivotal providers.
4.1. Pivotal Private Order Flow Providers
Prominent private order flow providers are listed in Tab. 6. Based on their roles in the MEV supply chain (see Fig. 2), we categorize them into three types: searcher, Telegram trading bot, and channel.
We now present our method to quantify the importance of individual providers. An existing approach (Lu, 2024; Lu, Angela and Sui, Danning and Durnford, Jaden, 2024) is to compute the contribution of each provider to the builders’ income (Sec. 2.3) during a specific period. However, this approach only considers the magnitude of impact but neglects the duration (in terms of time or the number of auctions). For example, a single private transaction, such as one444 0xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip, could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot. Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that can affect half of the MEV-Boost auctions in over two weeks. This captures their sustained impact on the system.
We define a provider $P$ as pivotal for an instance of MEV-Boost auction if removing $P$’s transactions from the winning block causes the winner to lose the auction. Intuitively, had the winner not accessed $P$’s order flow, the winner might lose. [^5] Note that the winner may or may not actually lose as it may be able to find an alternative pivotal provider, but this does not invalidate the fact that $P$ is pivotal. We identify pivotal providers in a given auction as follows: for each winning bid, we identify the providers of private order flow within the block and subsequently calculate the revised true value excluding each provider’s contribution. If the recalculated true value drops below the next highest bid (from the partial bids dataset Sec. 3.1), the involved private order flow provider is considered pivotal. Note that there can be multiple pivotal providers for an auction.
For each pivotal provider, we compute the fraction of auctions in which it is pivotal. We call this metric the pivotal level, a number from 0 to 1. Fig. 4 plots the pivotal level of the top-5 providers.
We can make a few observations from Fig. 4. First, a particular searcher, jaredfromsubway.eth, has been gaining prominence since February 2023 and reached its peak significance in May 2023. During this period, the winners of more than 70% of the MEV-Boost auctions would have failed had they not received order flow from jaredfromsubway.eth. Its influence has gradually decreased since then but still remains pivotal in 20%-40% of auctions since Nov 2023. Besides, the order flow from Telegram trading bots has become vital for builders with the rise of such bots. For instance, Banana Gun is pivotal in about 40% of the MEV-Boost auctions in September 2023.
Figure 4. A provider $P$ is pivotal if the winner of an auction could have lost without order flows from $P$. The pivotal level of top-5 pivotal providers over time. A line may not start from the beginning because the provider did not exist at that point. Maestro became an independent provider around December 2023.
Second, we can observe a positive trend from Fig. 4 that MEV Share and MEV Blocker have been increasingly pivotal since May 2023. These two providers positively impact decentralization because they have relatively clear requirements for accessing their order flows (the requirements of different providers will be discussed in Sec. 4.2).
Third, we observe a notable downtrend in the pivotal level of MEV-Share and MEV Blocker starting from November 2023, accompanied by an uptrend in Maestro, a popular Telegram trading bot (Maestro, 2023a), beginning in December 2023. This is because Maestro used to be an important upstream of MEV-Share and MEV Blocker but stopped in November 2023, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned. For example, Mastro may prefer to send flows to builders directly for a better price than what MEV-Share and MEV Blocker can offer. It is alarming to see that upstream providers can significantly swing the significance of platforms.
Finding II: We identified five pivotal providers (MEV-Share, MEV Blocker, jaredfromsubway.eth, Banana Gun, Maestro) who had a sustained influence on the outcome of more than half of MEV-Boost auctions. A positive trend towards decentralization is observed as the pivotal level of the searcher jaredfromsubway.eth decreased, while the pivotal level of MEV-Share and MEV Blocker increased. However, their reliance on upstream providers like Maestro is alarming.
Builder-specific Pivotal Level.    We compute each provider’s builder-specific pivotal level to understand the relationship between providers and builders. For a given provider $P$, its pivotal level specific to builder $B$ is the pivotal level of the MEV-Boost auctions where $B$ is the winner.
Fig. 5 plots selected providers’ builder-specific pivotal levels for the top-3 builders, ranked based on their weekly market share in April 2024. We include providers whose weekly pivotal levels exceed 25% for any of the three builders over two months to illustrate these builder-specific pivotal levels. We note a clear variation in the importance of providers for different builders. For example, the significance of MEV Blocker and Banana Gun to Titan Builder is greater than they are to other builders. Furthermore, we observe that some providers have a pivotal level exceeding 50% only with specific builders, suggesting a possible collaborative relationship. The searcher 0xA69b… consistently influences over 50% of the MEV-Boost auctions won by Beaver Builder, while the searchers 0x0087…, 0x2800…, and 0x51c7… successively affect those won by Rsync Builder. These pivotal providers are all integrated searchers of the corresponding builders that have been identified in previous works (Heimbach et al., 2024; Gupta et al., 2023). This observation corroborates findings from previous studies (Heimbach et al., 2024; Gupta et al., 2023; Thiery, 2023), confirming that the order flows from integrated searchers are crucial to those builders.
Figure 5. Weekly builder-specific pivotal level for top-3 builders.
4.2. Quantifying the Access Barrier
Pivotal providers have different degrees of openness to builders: searchers are generally not accessible, and platforms like MEV Share and MEV Blocker are more open, although they also set up certain barriers. In this section, we present more details about the access barrier set up by providers.
Searchers.   Since searchers are operated by anonymous entities without any public communication channels, there is no public statement about which builders may access their flows. Existing studies indicate that searchers tend to send their order flow to the top builders (Titan, 2023) or share it exclusively with vertically integrated builders (Titan, 2023; Thiery, 2023; Gupta et al., 2023).
Telegram trading bots.   The requirements for Telegram trading bots, in terms of sending order flow to builders, depend on whether they utilize channels like MEV Blocker (Maestro, 2023b; Labs, 2023). If so, their requirements are aligned with those of these channels. Banana Gun does not utilize these channels, nor does it explicitly state the requirements for builders.
Channels.   Since channels are public-facing services, we were able to extract their access requirements from the documentation and/or through their communication channels. The requirements of each provider are summarized in Tab. 6.
The channels impose various requirements on a builder’s reputation. Market share—the percentage of blocks built by a given builder—is a metric most relay providers use to evaluate builders’ reputations. Intuitively, a builder with a high market share, seen as reputable, is less likely to compromise private order flows as doing so will forgo their future revenue. MeowRPC (Meow RPC, 2024) sends the order flow to trusted builders with high market share. Similarly, MEV-Share (Flashbots, 2024) requires that a builder possesses a competitive market share. MEV Blocker (MEV Blocker, 2023) requires that a builder must maintain at least 1% market share over one week.
The cost of establishing minimal market share.   New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share. In practice, builders pay out of pocket to overbid in MEV Boost auctions, a practice known as block subsidization. This represents a capital entry barrier for new builders, which we are able to quantify using the auction dataset.
We compute the minimal subsidy needed to establish a 1% market share for one week (which is the requirement of MEV Blocker (MEV Blocker, 2023)) by replaying historical auctions with an added new builder who can only access the public order flow. We relegate the details to Sec. C.4.
Figure 6. Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
The result is plotted in Fig. 6. Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 1.4 ETH by March 2024. It is important to recognize that the costs derived from this method only represent the lower bound. Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory. Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share. In reality, the subsidies new builders need to provide are substantially higher. For example, tbuilder (Tbuilder, 2024), a new builder in the builder market, subsidized 6.8 ETH in total for 645 blocks in the week before March 21, 2024 (Flashbots, 2024).
Finding III: The private order flow providers impose reputation requirements on the builders, which are usually evaluated based on market share. The minimum entry cost for new builders required to establish a required market share increases over time.
4.3. Trust Crisis and Centralization
Our findings highlight a trust issue in today’s MEV supply chain: the lack of trustworthy fair exchange mechanisms between parties creates order flow silos and harms decentralization.
Private order flow providers are at a disadvantage when interacting with builders and must trust builders not to unbundle the transactions. As a result, providers rely on an informal reputation system for self-protection, but it is not hard to see that reputation mechanisms are easily bypassed, as blockchain is fundamentally anonymous. Once a reputable builder receives the bundle from the searcher, the builder can create a Sybil and cheat the searcher (e.g., by unbundling the transaction bundle) under the new name.
While such an attack has not been reported in practice, it is noteworthy that builders occasionally unbundle searcher’s bundles. For example, the f1b builder unbundled a bundle from the searcher jaredfromsubway.eth at block height 18476515. Although the searcher was compensated afterward, this incident underscores the fragile security foundation of the trust chain in the MEV-Boost auction. We also noticed that one anonymous builder666public key: 0xad5ce1019364de39…35af575554a7aa6f might be the Sybil of a reputable builder because this builder’s blocks included order flows from sources that are only accessible by reputable builders.
The malicious actors within the trust chain pose a dual threat. On the one hand, security concerns deter searchers from submitting their bundles to multiple builders, further entrenching private order flow as a barrier to entry for new builders. On the other hand, integrating the roles of searcher and builder offers a potential mitigation to the security issue. Integrated searcher-builder has other structural advantages (Gupta et al., 2023; Pai and Resnick, 2023) (such as low latency) that may be unfair to regular builders.
In Sec. 7.2, we discuss directions to bridge the trust gap between searchers and builders.
Finding IV: The requirements for builders stem from the barrier of trust, as illustrated by the hidden trust crisis confirmed through real-world examples. Lowering or removing trust requirements in the MEV-Boost auction can facilitate decentralization.
5. Competition and Efficiency in the MEV-Boost auctions
Sec. 4 focuses on the evolution of builder market. In this section, we focus on the status quo and study the implications of a centralized builder market as we have today. First, we investigate the (in)equality of builders’ block-building capability, measured by the true values of builders’ bids in MEV-Boost auctions. Then, we quantify the competitiveness and efficiency of historical MEV Boost auctions. These two studies complement each other. The former evaluates whether today’s builder market has the potential to be competitive, while the latter evaluates whether the current mechanism in fact induces the desired outcomes. Our study includes some MEV-Boost auctions, the highest bids of which are not included on Ethereum due to network latency. However, we do not omit them since the auctions themselves are still meaningful.
5.1. Inequality in Block-building Capacity
A primary metric to evaluate a builder’s block-building capacity is the MEV it extracts over time. In MEV Boost auctions, this quantity is captured by the true values of the builder’s bids.
In the context of auctions generally, the true value of the item being auctioned refers to its actual worth to the bidder. Recall that in MEV-Boost auctions (Sec. 2), a builder submits a block $B$ (a list of transactions) and a bid value $BV$ that the builder is willing to pay to the proposer if $B$ is proposed. The true value of $B$, denoted $TV(B)$, is typically higher than $BV$. One can think of $TV$ as the builder’s revenue and $BV$ as the cost. The difference $EV:=TV-BV$ is the profit pocketed by the builder, which we call extracted value (EV) — i.e., $EV(B)$ is the difference of the builder’s ETH balance after executing transactions in $B$. For example, suppose executing $B$ yields a net profit of $0.1$ ETH after paying $0.9$ ETH in bid value in the auction, then the true value (the revenue) is $1.0$ ETH.
To compute $TV(B)$ of a given block $B$, we follow the standard approach in open source MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023), though we built our own analysis toolchain and optimized its performance to handle the large volume of auction data. We compute $EV(B)$ using a modified Rust Ethereum execution client (Reth (paradigmxyz, 2024)) to execute transactions in $B$ and compute the net ETH balance changes of the bidder’s addresses. Most builders’ addresses are well-known partly because they need to build up a reputation. We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec. 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
In every slot $s$, bidders can submit multiple bids throughout the duration of the slot. We use $BV(p,s)$ to denote the highest bid value from bidder $p$ in slot $s$, and $TV(p,s)$ the corresponding true value.
Quantifying the inequality of block-building capacities.   We compare the highest true values of different builders in the same slot to understand the differences in builders’ ability to extract values. If a builder consistently has the highest true value, then she can potentially dominate the builder market. On the contrary, if many builders have similar true values, then it is plausible that a properly designed market could avoid a monopoly.
To quantify the disparity of block-building capacity, we compute the quartile coefficient of dispersion (QCD) (Wikipedia, 2022) of the highest true values in a given slot. Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$. Let $Q_{1}(\vec{x})$ and $Q_{3}(\vec{x})$ representing the first and third quartiles of $\vec{x}$ (i.e., 25% and 75% percentiles). The QCD of slot $s$ is
$QCD_{s}=QCD(\vec{tv}_{s})=\frac{Q_{3}(\vec{tv}_{s})-Q_{1}(\vec{tv}_{s})}{Q_{3}(\vec{tv}_{s})+Q_{1}(\vec{tv}_{s})}.$
QCD is a real value in $[0,1]$. A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$). Since it uses quartiles, QCD has the benefit of being robust to outliers (Wikipedia, 2022).
We now present the result given by QCD, noting that other metrics lead to a similar conclusion (see Sec. D.1).
Evaluation results.   We categorize the builders into three groups based on their market share from April to August 2023: top builders (top 5), middle builders (6-15), and tail builders (16-25). To understand how the magnitude of MEV in a slot affects the block-building capability, we categorize every slot into one of three tiers according to the final winning bid (as a proxy to the significance of the MEV opportunity in that slot): low MEV (below 0.0225 ETH, the 10th percentile of the winning bid values), high MEV (above 0.4586 ETH, 95th percentile) and medium MEV in between.
In Fig. 7, we plot the distribution of the QCD for MEV-Boost auctions from April 2023 to August 2023 for different builder groups and MEV tiers. (We sample the first week of each month. See Sec. 3.1 for details.) It reveals two trends. In each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail. That is, top builders have comparable block-building capabilities, but middle and tail builders have significantly different capabilities. This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of builder market.
Second, the inequality worsens as the MEV tier increases. One possible reason is that private order flow providers may tend to give significant MEV opportunities to select builders due to the trust barrier.
Figure 7. Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders. The QCD measures the inequality of block-build capacity across builders. This figure reveals that such inequality worsens as MEV increases. Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
Finding V: The inequality in block building is lowest among top builders, higher among middle builders, and most pronounced among tail builders, and the inequality worsens as a slot’s MEV increases.
5.2. Quantifying Competitiveness and Efficiency
The previous section shows that top-5 builders have comparable true values in most cases. Ideally, revenue from MEV-Boost auctions should be near optimal due to competition. However, whether real-world auctions induce such near-perfect competition is complicated by several confounding factors: builders may not be able to respond to bids due to network latency, they may collude implicitly or explicitly, or they may have incentives to overbid (e.g., to gain market share). In this section, we quantify the competitiveness and efficiency of the MEV-Boost auctions using what we call competitive index (CI) and efficient index (EI).
Competitive index.   In a given slot $s$, suppose builders $p_{1},\cdots,p_{n}$ are ordered by their true value from high to low, i.e., $TV(p_{1},s)\geq\dots\geq TV(p_{n},s)$, $BV_{w}(s)$ is the bid value of the winner and $TV_{w}(s)$ is the corresponding true value, the competitive index of slot $s$, $CI(s)$ is defined as:
$CI(s)=\frac{BV_{w}(s)-TV(p_{2},s)}{TV(p_{2},s)}\times 100\%.$
Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value. $CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, satisfying the definition of a competitive auction. Conversely, $CI(s)<0$ means that the competition is insufficient.
Efficient index.   Same as above, the efficient index, $EI(s)$, is defined as:
$EI(s)=\frac{TV_{w}(s)-TV(p_{2},s)}{TV(p_{2},s)}\times 100\%.$
$EI(s)>0$ indicates that the winner has the highest true value, whereas $EI(s)\leq 0$ implies that the bidder with the highest true value lost the auction.
Figure 8. The distribution of CI and EI, where each dot represents an auction. The color represents density when multiple dots overlap. 88.84% of the MEV-Boost auctions are competitive (sum of the slots having $CI\geq 0$), and 79.74% of them are efficient (sum of the slots with $EI>0$). Dots on the diagonal line ($CI=EI$) are auctions where the winners bid the true values. The dots above this line are auctions where the winners bid lower than true values.
Fig. 8 shows the distribution of the calculated CI and EI across all MEV-Boost auctions in the ultra sound relay dataset. Notably, 74.83% of the auctions are both competitive and efficient: the winner possesses the highest true value and bids higher than or equal to the second-highest true value. 88.84% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 79.74% (sum of the areas with $EI>0$) are efficient.
5.2.1. Implications of uncompetitive auctions
We first note Fig. 8 shows a somewhat strict definition of competitiveness (i.e., $CI\geq 0$). In reality, builders may bid slightly lower than their true value to ensure profits. Therefore, it is reasonable to relax and consider an auction with $CI\geq-\delta$ to be competitive for a small $\delta$. We call this notion $\delta$-competitive. As shown in Fig. 14, when we consider $\delta=1\%$ or $\delta=2\%$, the percentage of uncompetitive auctions reduces to about 6% and 5%, respectively. That is, under the slightly relaxed definition, approximately 95% auctions were competitive.
Quantifying proposers’ losses.   The concern with uncompetitive auctions lies in the potential loss to validators. If such losses are significant, top validators might be motivated to operate their own builders. Furthermore, vertical integration between top proposers and builders could adversely impact the decentralization of the validators (Bahrani et al., 2024). Thus, quantifying the exact loss validators incur due to insufficient competition is crucial.
For each uncompetitive auction, we consider the second-highest true value as the ideal winning bid value and calculate the difference between this ideal value and the actual winning bid value. This difference represents the loss a validator incurs, which they would have received had the auction been competitive.
Table 3. Validators’ losses due to insufficient competition
Time #Slots #Slots w/ CI ¡ 0 (%) Profits (ETH) Losses (ETH) (%)
April 9-15 28,376 1,437 (5.06%) 2,711.32 47.17 (2.02%)
May 1-7 30,279 8,449 (27.90%) 9,367.50 118.50 (1.36%)
June 1-7 35,414 3,137 (8.86%) 4,346.51 24.72 (0.62%)
July 1-7 36,032 2,051 (5.69%) 3,945.67 18.41 (0.48%)
August 1-7 17,825 1,424 (7.99%) 2,171.07 12.29 (0.58%)
As shown in Tab. 3, we note that the total losses caused by uncompetitive auctions amount to 221.09 ETH in our study, which represents 0.98% of the total gains of proposers, with an average loss of 0.0134 ETH per uncompetitive auction.
The losses are particularly significant during May 1-7, 2023, when a total of 118.50 ETH was lost due to insufficient competition, which is 1.27% of their total gain. These losses may closely correlate with MEV opportunities, given that proposers’ total profits were also notably high during this period. Consequently, even though the losses seem to decrease, this reduction should not be interpreted as an indicator of increasingly intense competition. The concern regarding potential centralization persists, given that the losses experienced by validators have not been fully mitigated.
Finding VI: About 11% of the MEV-Boost auctions in our study are uncompetitive, resulting in total losses of 221.09 ETH for the proposers. These losses represent 0.98% of the total gains of proposers.
5.2.2. Reasons of inefficiency
20.26% of MEV-Boost auctions were not efficient. Further investigation is required to understand why. In an inefficient auction, the builder with the highest true value does not win the auction. There are two possible reasons. First, block subsidization allows a builder to win the auction without possessing the highest true value. Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
We analyzed the underlying reasons for all inefficient auctions. In 51.5% of cases, winners win the auctions due to subsidization. In 52.6% of cases, the builder with the highest true value shaded too much. It is important to note that auctions may exhibit inefficiency for multiple reasons. For instance, the winner may subsidize her bid while simultaneously, the builder with the highest true value shades. Thus, the cumulative percentage exceeds 100%.
It is important to note that inefficiency does not necessarily disadvantage the validators. In fact, in inefficient scenarios involving subsidization, validators may receive more value.
Finding VII: MEV-Boost auctions are generally competitive, yet their efficiency falls short of expectations. Further analysis reveals that 51.5% of inefficiencies are caused by subsidization, which, interestingly, does not disadvantage proposers but leads to increased earnings for them in such scenarios.
5.2.3. MEV’s implication on auction outcomes
To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we follow the same methodology and categorize the auctions into three tiers: low MEV, medium MEV, and high MEV. The ratio of competitive and efficient auctions in each tier is shown in Fig. 9.
We note that competitiveness decreases as the MEV of a slot increases, which is consistent with the finding in Sec. 5.1 about the inequality of block-building capabilities in high-MEV slots.
The trend of efficiency is less expected. First, only 71.7% of the auctions within the low MEV tier are efficient. This can be attributed to the generally low overall bid values, rendering the auction outcomes vulnerable to alterations from minor subsidies. Second, while auctions tend to be more efficient in the medium MEV tier, we also observe that efficiencies worsen in the high MEV tier: only 77.2% of these auctions are efficient, and about 90% of inefficient cases result from bid shading, where the builders with the highest true values place their bids much lower than their true values.
Figure 9. The ratio of auctions that are competitive and efficient across various MEV tiers.
Finding VIII: The competitiveness of MEV-Boost auctions deteriorates as a slot’s MEV increases. Efficiency may decline due to block subsidization in the low MEV tier and bid shading in the high MEV tier.
6. Related Work
Empirical study on Ethereum’s builder market.   Several previous works studied the centralization of the builder market shares. An early exploration by Yang et al. (Yang et al., 2022) analyzed the market share of relays and builders. Their findings indicated a centralization of Ethereum’s builder market, with Flashbots and builder0x69 accounting for over 53% of the market share between September 15th, 2022, and November 30th, 2022. Similar studies on the builder market conducted by (Wahrstätter et al., 2023b; Heimbach et al., 2023) also confirmed the centralization in the builder market. Their observations were based on the market share and the distribution of total profits generated within the PBS ecosystem. Moreover, several online dashboards (Wahrstätter, 2024; Labs, 2024; Bitfly, 2024; Flashbots, 2024) offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics. However, these studies do not quantify the causes of centralization or the security implications of a centralized builder market.
Private order flow.   Many studies also discuss the importance of private order flows, particularly focusing on private order flows from integrated searchers. Thiery (Thiery, 2023) counted the total number of transactions, along with their values, from June 1 to July 15, 2023. He found that private order flows account for only 30% of the transactions yet represent 80% of the total value paid to builders. Our study highlights a similar result: in over 50% of blocks, private order flows account for more than 60% of the block value. Gupta et al. (Gupta et al., 2023) confirmed that builders are more likely to win the MEV auction when their integrated searchers can provide high-value exclusive order flow by effectively exploiting arbitrage between a centralized exchange (CEX) and a decentralized exchange (DEX). The dashboard by Winnie (Xiao, 2024) and the study by Titan (Titan, 2023) analyzed the vertical integration between searchers and builders through the quantity and value of transactions sent from searchers to builders. Heimbach et al. (Heimbach et al., 2024) observed that during their study, Beaver Builder received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH. This large gap suggests that Beaver Builder might have received significant off-chain profits from its integrated searchers. These studies show that three integrated builders (Beaver, Rsync, and Manta) gain a significant advantage in order flow from their integrated searchers. Our study made the same observation and further quantified the pivotal level.
Strategic behaviors in MEV-Boost auctions.   The strategies used by different entities in the MEV-Boost auctions and their potential impacts have attracted significant interest from researchers. The strategic behaviors of proposers were first studied by (Schwarz-Schilling et al., 2023; Wahrstätter et al., 2023b) focusing on the timing game where proposers strategically delay block proposals to maximize their profits. Öz et al. (Öz et al., 2023) further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it. A recent study (Always, 2024) empirically investigated timing games in the wild and how block proposal delays affect the block fork rate. The results indicate that three validator sets–P2P, Kiln, and Attestant, that are known or suspected of playing timing games–did not delay their block proposals enough to cause their blocks to be forked. Wu et al. (Wu et al., 2024) proposed a game-theoretic model for MEV-Boost auctions, using simulations to examine various builders’ bidding strategies, including naive, adaptive, last-minute, and bluff bidding. Their results highlighted the importance of latency and exclusive order flow in the effectiveness of bidding strategies. Pai et al. (Pai and Resnick, 2023) analyzed the latency advantage of searcher-builder integration in adjusting bidding values. The strategies discussed in these papers highlight the potential factors that can make MEV-Boost auctions uncompetitive or inefficient, while our paper quantifies these uncompetitive and inefficient auctions.
Causes and implications of centralized builder market.   A provider may exclusively share its order flow with a single builder. Previous works such as (Kilbourn, 2022; Frontier, 2023; Gupta et al., 2023) pointed out that exclusive order flows have potential centralizing effects on the builder market, because builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers and in winning order flow auctions. The decentralization of Ethereum’s builder market not only benefits itself but also positively impacts the decentralization of validators. A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization (Bahrani et al., 2024). We complement these theoretical analyses with empirical studies to analyze potential problems caused by the current centralized builder market.
Blockchain decentralization metrics.   Nakamoto coefficient was proposed to describe the minimum number of entities required to disrupt the blockchain’s subsystem (Srinivasan and Lee, 2017). Other metrics, such as the Gini coefficient (Dorfman, 1979), Shannon entropy (Lin, 1991), and Herfindahl–Hirschman index (Rhoades, 1993), have also been utilized to measure decentralization (or the lack of it) in staking distribution (Grandjean et al., 2023; Kwon et al., 2019), mining power (Lin et al., 2021; Wu et al., 2019; Kwon et al., 2019), and block building (Heimbach et al., 2023). An SoK paper (Zhang et al., 2022) comprehensively summarized the existing works on blockchain centralization; however, none explored decentralization from the perspective of entry barriers or auctions.
7. Discussion
7.1. Limitations of Our Approach and Data
True value.   We compute the true values of bids using a similar approach as that in MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023). This approach has a few limitations. First, it relies on the knowledge of the set of addresses controlled by each builder. Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses. We cross-validated our builder address dataset with other sources to ensure completeness (see Sec. 3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate. Secondly, this approach neglects off-chain profits, such as off-chain payments received from searchers. Off-chain payments are more likely to happen between integrated builder-searcher pairs. We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization. However, in our study, we did not observe any excessively large subsidies. Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness. We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value. However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value. It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
Missing pivotal providers.   Our results in Sec. 4.1 might omit some pivotal providers since the providers we can identify are a subset of all providers. For example, BloXroute (a widely used channel) does not publish the list of transactions they processed. Thus, we can’t evaluate its pivotal level. This limitation is common in existing studies (Lu, 2024; Lu, Angela and Sui, Danning and Durnford, Jaden, 2024). However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec. 3). Thus, the significance of each provider will only be underestimated rather than overestimated.
Our auction dataset might miss partial bids since they are self-reported by relays. Recall that we used the second-highest bid value in an MEV-Boost auction as the threshold to determine a pivotal provider. Therefore, underestimating the second-highest bid value might also lead us to underestimate the significance of each provider.
7.2. Future Works
Based on the findings, we now outline directions to enhance the decentralization of builder markets. These directions present interesting technical challenges that we leave for future work.
Several issues we identified boil down to the lack of a trustworthy mechanism to facilitate fair exchanges between private order flow providers and builders. In today’s ecosystem builders are at an advantage. To protect themselves, private order flow providers resort to reputation-based solutions such as requiring a minimal market share, which creates siloes.
One approach is building platforms to facilitate the said fair exchange. For example, MEV-Share is one such platform between searchers and builders. To receive order flows from MEV-Share, builders must agree to the Fair Market Principles set force by the community (Flashbots, 2023a). Should builders fail to adhere to these principles, they will be either removed from or flagged by the community.
However, in its current form, MEV-Share is a trusted party. One direction being explored (e.g., by SUAVE (Flashbots, 2024c)) is to use Trusted Execution Environments (TEEs) to run block-building algorithms without leaking private order flows to the builders, but standard challenges such as side channels and covert channels (Salles et al., 2023) apply.
Another unsolved problem is how to reliably detect private order flow leakage across builders. Currently, after receiving private orders, a builder can leak the orders to another builder (potentially a Sybil), bypassing any reputation or economic penalty. One idea is to embed watermarks in order flows to enable tracking. Watermarks must be invisible to the adversary and not interfere with transaction execution. Designing a secure watermarking scheme against a practical adversary model is an interesting future direction.
A reliable leakage detection mechanism (e.g., watermarking) will open up the design space of reputation systems. For example, one possibility is to ask builders to deposit a certain amount of collateral to access private order flows, and order flow leakage will be penalized economically.
8. Conclusion
In this paper, we characterized the efficacy of PBS by evaluating its openness to new builders and assessing the competitiveness and efficiency of the auctions. Our study identifies a significant entry barrier in the builder market: private order flow. We further quantify the barrier costs—a new builder must pay up to 1.4 ETH to access this flow. Given this barrier, it is not surprising that we observed varying capabilities among builders in extracting MEV, with inequalities among top, middle, and tail builders worsening as MEV increases. Our study shows that 88% of the MEV-Boost auctions were competitive, while 79% were efficient. The proposers’ losses due to uncompetitive auctions are 0.98% of their total gain. Based on these insights, we propose suggestions for future developments to mitigate the trust crisis in the MEV-Boost auctions and facilitate the decentralization of Ethereum’s builder market.
Acknowledgements
This work is partially supported by an Ethereum Foundation Grant. We thank ultra sound relay for sharing the full bid dataset. We also thank the teams of BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions about access requirements. We are grateful for the valuable discussions with Thomas Thiery and Barnabé Monnot.
Appendix A Visualization for distribution
Fig. 10 shows the distribution of the percentage of builders who submit to the ultra sound relay. Fig. 11 shows the CDF of the winning bid values of MEV-Boost auctions.
Figure 10. Distribution of the percentage of builders who submit to the ultra sound relay.
Figure 11. CDF of the winning bid value of auction dataset.
Appendix B List of Searchers and Builders
In Tab. 4, we provide the list of searchers and their addresses mentioned in our study. In Tab. 5, we have counted the top 25 builders in the market, ranked by their market share from April 2023 to August 2023.
Table 4. The list of searchers
Searcher Address
jaredfromsubway.eth 0xae2Fc483527B8EF99EB5D9B44875F005ba1FaE13
Searcher: 0xA69b… 0xA69babEF1cA67A37Ffaf7a485DfFF3382056e78C
Searcher: 0x0087… 0x0087BB802D9C0e343F00510000729031ce00bf27
Searcher: $0$x280027dd00eE0050d3F9d168EFD6B40090009246
Searcher: 0x51c7… 0x51c72848c68a965f66fa7a88855f9f7784502a7f
Table 5. The percentages of blocks and MEV-Boost blocks built by each builder from April to August 2023.
Builder % of total blocks % of MEV-Boost blocks
Beaver 22.418739 24.719523
Rsync 18.611899 20.521996
Builder0x69 17.799124 19.625807
Flashbots 13.276622 14.639172
Titan 7.105763 7.835011
BloXroute 2.384068 2.628740
Blocknative 2.267458 2.500162
F1b 2.158849 2.380407
eth-builder 1.086828 1.198366
BuildAI 0.894899 0.986741
payload.de 0.645310 0.711536
lightspeedbuilder 0.321781 0.354805
boba-builder 0.257682 0.284128
Manta 0.254004 0.280072
Eden 0.202596 0.223388
0x3be…436 0.186227 0.205339
It’s Free Real Estate 0.133072 0.146728
Gambit Labs 0.117806 0.129896
bobTheBuilder 0.102815 0.113367
Manifold 0.088929 0.098056
blockbeelder 0.084239 0.092884
finest artisanal blocks 0.074307 0.081933
nfactorial 0.072375 0.079803
0x563…cf3 0.041660 0.045935
uwu.builders 0.016370 0.018050
Appendix C Openness of the builder market
C.1. Importance of private order flows
To quantify the importance of private order flows for building valuable blocks, we compute the ratio of block values derived from private transactions. We first identify all private order flows using methods described in Sec. 3.1. Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total transaction priority fees (Ethereum, 2024) and direct transfers (Flashbots, 2024a) to the block builder) to the total values extracted from all transactions in the block.
C.2. Case Study for Maestro
To delve deeper into the relationship between these two trends, we plot Maestro’s contribution to the total daily value of transactions sent through MEV-Share and MEV Blocker during our study in Fig. 12. Surprisingly, during the time periods when MEV-Share and MEV Blocker were highly pivotal in Fig. 4, i.e., from May to November, Maestro’s contribution accounts for approximately 80%! This means Maestro used to be an important upstream of MEV-Share and MEV Blocker. In November 2023, Maestro stopped sending transactions to MEV-Share and MEV Blocker, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
Figure 12. Daily percentage of the value of the transactions originating from Maestro for both MEV-Share and MEV Blocker.
C.3. Requirements of Providers
Tab. 6 summarizes the providers we identify and their requirements for accessing their private order flows.
Type Provider Requirements
Searcher individual searchers Requirements are unclear.
Telegram trading bot Banana Gun (Banana Gun, 2023), Maestro (Maestro, 2023a), Unibot (Unibot X, 2023) The requirements are the same as the channels they use or unclear.
Channel BloXroute (bloXroute Labs, 2024) Builders are granted access to different tiers, each providing varying privileges and permissions, based on an assessment.
Merkle (Merkle Software Inc., 2023) Builders must have a proven track record of accurately and honestly computing bundles, or they should consult with the Merkle team for consideration.
MeowRPC (Meow RPC, 2024) The transactions are sent to the top builders which MeowRPC’s team trusts.
MEV Blocker (MEV Blocker, 2023) Builders must maintain at least 1% market share over one week.
MEV-Share (Flashbots, 2024) Builders must demonstrate a competitive market share for consideration.
Transaction Boost (Blocknative, 2024) Consideration for partnership is based on the Blocknative team’s evaluation of the builder’s reputation.
  • $\star$
    We do not claim that the list of the private order flow providers is complete since there is no ground truth.
Table 6. Collections of the private order flow providers and their requirements.
C.4. Minimal Cost for 1% Market Share
We now present how we compute the minimal subsidy for a builder to account for 1% market share over one week. We assume that all builders have access to public order flows but the new builder cannot access any of the private order flows. Furthermore, this new builder can win the auction if their bid value is the same as other builders. Other builders will not submit bids beyond their true value, defined as the total revenue of their blocks. Or they will not increase subsidies if they already subsidize.
Under our assumptions, a new builder must submit a bid at least equal to the auction winner’s maximum–the higher of its bid value or true value. The revenue from the new builder’s block equals the value contributed by the public order flows to the winning bid. Thus, the subsidy needed for winning the auction is the difference between the revenue from public order flows and the required bid value. Then, as we know the subsidy for proposing one block, we can further calculate the minimum overall subsidy a new builder would need to cover to propose 72 blocks each day over one week (Ethereum produces 7200 blocks a day.)
Appendix D Competition and Efficiency in the MEV-Boost auctions
D.1. Measuring block-building capabilities
Coefficient of variation.   We use $\mathsf{CV}(\vec{v})\to\mathbb{R}_{\geq 0}$ to denote the coefficient of variation (CV), a real value in $\mathbb{R}_{\geq 0}$, for values $\vec{v}=(v_{1},\dots,v_{n})$. Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$. The CV of slot $s$ is computed as $CV(\vec{tv}_{s})$.
Figure 13. Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of block-build capacity across builders.
As shown in Fig. 13, we can make the same finding as Fig. 7. It reveals that such inequality worsens as MEV increases. Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
D.2. Implications of uncompetitive auctions
Fig. 14 shows the distribution of the competitive index (CI) of the MEV-Boost auctions in our study.
Figure 14. The distribution of the CIs of MEV-Boost auctions

### v3

@@
 Decentralization of Ethereum’s Builder Market
-Abstract.
-Blockchains protect an ecosystem worth more than \$500bn with their strong security properties derived from the principle of decentralization.
-Is today’s blockchain really decentralized?
-In this paper, we empirically studied one of the least decentralized parts of Ethereum—the most used blockchain system in practice—and shed light on the decentralization issue from a new perspective.
-To avoid centralization caused by Maximal Extractable Value (MEV), Ethereum adopts a novel mechanism that produces blocks through a builder market.
-After two years in operation, however, the builder market has evolved to a highly centralized one with three builders producing more than 90% of blocks.
-Why does the builder market centralize, given that it is permissionless and anyone can join?
-Moreover, what are the security implications of a centralized builder market to MEV-Boost auctions?
-Through a rigorous empirical study of the builder market’s core mechanism, MEV-Boost auctions, we answered these two questions using a large-scale auction dataset we curated since 2022.
-Unlike previous works that focus on who wins the auctions, we focus on why they win, to shed light on the openness, competitiveness, and efficiency of MEV-Boost auctions.
-We show the access barriers around private order flows constitute a significant entry barrier to the builder market.
-A new builder needs to pay up to 1.4 ETH to access this private order flow, a significant amount considering their low initial profits.
-By computing the true value of bids, we investigated the inequality in block-building capabilities among builders and the competitiveness and efficiency of the MEV-Boost auctions.
-We observed that top, middle, and tail builders have significantly varying capabilities in extracting MEV, and the inequality worsens when MEV increases.
-88% of the MEV-Boost auctions we studied were competitive, yet uncompetitive auctions still resulted in a total loss of 221.09 ETH for the proposers, which is 0.98% of their total gain.
-Meanwhile, only about 79% of the MEV-Boost auctions were efficient, and over half of the inefficient auctions were caused by block subsidization.
-Our findings also help identify directions for improving the decentralization of builder markets.
-1. Introduction
-At the core of any decentralized blockchain, a key assumption underpinning their strong security properties (integrity, availability, censorship resistance, incentive compatibility, etc) is “decentralization”, that the system is run by a large number of independent participants (Nakamoto, 2008; Eyal and Sirer, 2018; Schwarz-Schilling et al., 2022; Wahrstätter et al., 2023a; Bahrani et al., 2024; Roughgarden, 2020).
-It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) (Daian et al., 2020), profits that protocol participants, hereafter referred to as validators, can reap by manipulating the ordering of transactions.
-MEV can lead to centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have (Bahrani et al., 2024; Ethereum Foundation, 2024b; Pai and Resnick, 2023; Buterin, 2021b, a).
-In an effort to level the playground for validators, the Ethereum community proposed Proposer [^1] A proposer is a validator chosen as the leader in a round of consensus to add a new block to the blockchain.-Builder Separation (PBS) (Ethereum Foundation, 2024b).
-The idea is to outsource the task of building profitable blocks to specialized entities called block builders.
-In the current implementation called MEV-Boost (Flashbots, 2024), the proposer runs an auction among builders and adds the block with the highest bid to the blockchain.
-MEV-Boost is permissionless for builders so any builder can join the auction.
-Ideally, builders compete with each other in auctions, and the auctioneer—the proposer—collects a significant portion of MEV in auction revenue without doing much work, eliminating the competitive advantage of sophisticated validators over regular ones.
-Figure 1. Market share of builders. The builder market is arguably one of the least decentralized parts of Ethereum!
-Does PBS achieve its goal?
-The answer is rather unclear.
+Abstract
+Blockchains protect an ecosystem worth more than \$500bn with strong security properties derived from the principle of decentralization.
+Is today’s blockchain decentralized?
+In this paper, we empirically studied one of the least decentralized parts of Ethereum, its builder market.
+The builder market was introduced to fairly distribute Maximal Extractable Values (MEV) among validators and avoid validator centralization.
+As of the time of writing, three builders produced the vast majority (more than 80%) of blocks in Ethereum, creating a concerning centralization factor.
+However, the community believes that such centralization “is okay,” arguing that builder centralization will not lead to validator centralization.
+In this empirical study, we interrogate the causes and implications of builder centralization and challenge this belief that it is acceptable.
+Our main finding is that builder centralization has led to a significant loss by validators and, if left uncontrolled, could lead to validator centralization.
+Moreover, MEV mitigation solutions slated for adoption are affected too because they rely on the builder market as an “MEV oracle,” which is made inaccurate by centralization.
+Our investigation revealed two reasons behind builder centralization.
+We propose a structural change to the existing MEV supply chain and a solution based on the new supply chain structure.
+However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
+1 Introduction
+At the core of any decentralized blockchain, a key assumption underpinning their strong security properties (integrity, availability, censorship resistance, incentive compatibility, etc.) is “decentralization”, that the system is run by a large number of independent participants [63, 28, 75, 83, 4, 74].
+It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) [16], profits that protocol participants, known as validators, can reap by manipulating the ordering of transactions.
+MEV can lead to centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have [4, 23, 67, 12, 11].
+To distribute MEV fairly among validators and avoid validator centralization, Ethereum adopted a solution called Proposer-Builder Separation (PBS) [23].
+A proposer is the validator chosen to add a new block to the blockchain.
+In the current implementation called MEV-Boost [35], the proposer runs an auction among builders (anyone can be a builder) to solicit the best block.
+Ideally, builders compete with each other, and the proposer collects a significant portion of MEV in auction revenue without doing much work, eliminating the advantage of sophisticated validators over regular ones.
+Figure 1: Market share of builders. The builder market is one of the least decentralized parts of Ethereum!
+Is the builder market working as intended?
+The answer is rather unclear because the builder market is highly centralized.
 As shown in Fig.
-1, the current market share is highly concentrated, with over 90% of the blocks produced through MEV-Boost being built by just three builders.
-However, the current concentration does not necessarily defeat the purpose of PBS for two reasons: first, builder markets may evolve and new builders may join and disrupt the concentration; second, centralization does not imply a lack of competition—it only takes a few non-colluding bidders to make the auction competitive and efficient.
-Thus, while previous works (Flashbots, 2024; Wahrstätter, 2024; Labs, 2024; Heimbach et al., 2023; Wahrstätter et al., 2023b; Yang et al., 2022; Thiery, 2023) point out the centralization issue, they did not answer two key questions: First, how open is the builder market to new builders?
-Second, despite the centralization, are MEV auctions actually competitive and efficient?
-(We leave technical definitions to future sections, but intuitively, an auction is competitive if the winner outbids the second-best bidder; an auction is efficient if the bidder with the highest valuation wins.)
-In this paper, we answer these questions through a rigorous empirical study of historical MEV auctions.
-The key ingredient of our study is a large-scale auction dataset we curated since 2022: we collected 5.57 billion partial bids [^2] Partial and full bids are to be defined in Sec.
-3.1 of 3,810,630 auctions from September 2022 to March 2024; in addition, in collaboration with ultra sound relay (Money, 2023), we obtained full bids for a total of 147,926 auctions from April 2023 to August 2023.
-Auction data gives us direct visibility into the openness, competitiveness, and efficiency of MEV-Boost auctions.
-For example, historical bids allow us to “simulate” past auctions assuming a different set of transactions were available to builders, allowing us to discern the importance of different transactions.
-Another example is that we can compute the true value of each bid (i.e., the actual worth of the block to a builder, which is usually higher than what they bid) from transactions.
-The competitiveness and efficiency of an auction follow directly from a comparison between the true values and bid values.
-This is in contrast to, say, eBay auctions, where the true value of an item is subjective and cannot be retroactively computed from data.
-Now we present an overview of methods and findings.
-Table 1. Summary of findings
-Finding 1: Private order flows contribute to 60% of the MEV in over 50% of the blocks in MEV-Boost.
-Finding 2: To access private order flows, new builders need to pay up to 1.4 ETH to enter the market.
-Finding 3: The inequality in block building is highest among tail builders, followed by middle and least among top builders, and it worsens when MEV increases.
-Finding 4: 88.84% of the MEV-Boost auctions are competitive, with proposers’ losses from uncompetitive auctions amounting to 0.98% of total gains.
-Finding 5: 79.74% of the MEV-Boost auctions are efficient, and 51.4% of the inefficient auctions are caused by block subsidization.
-1.1. Openness of the Builder Market
-First, we investigate the reasons behind centralization by quantifying the market entry barrier for a new builder.
-One such barrier could be the ability to build effective MEV extraction algorithms, but we investigated a perhaps more basic barrier: the ability to access profitable transactions.
-Recall that MEV extraction is about manipulating the ordering of transactions, so having access to “extractable” transactions is the basis of any extraction algorithm.
-As we will present in Sec.
-4, private order flows (Kilbourn, 2022)—the transactions that bypass the public mempool—contribute an increasing portion to builder income.
-This means that it would be difficult for a builder to win the auction without access to private order flows.
-Moreover, not all private order flows are equally profitable for builders, so the ability to access “good” private order flows forms a barrier to entering the builder market.
-Using the auction dataset, we identified the providers of profitable transactions, quantified their importance to the builders, and investigated their accessibility to new builders.
+1, three builders won over 90% of auctions, raising questions about the competitiveness of the market.
+However, the community believes that “it is OK to centralize block building” [23], arguing that block centralization will not lead to validator centralization.
+We challenge this belief through an empirical study of the past MEV-Boost auctions.
+As will be shown in detail, builder centralization has caused a significant loss to the proposers and, if left uncontrolled, could lead to undesired consequences including validator centralization.
+1.1 Implications of Builder Centralization
+Proposer loss and its consequences.
+Intuitively, a proposer’s loss is the difference between the MEV available in a block and the auction revenue.
+Builder centralization can cause proposer loss in two ways.
+First, auctions may not be competitive enough due to, e.g., collusion or other limitations of the auction mechanism.
+Second, even if builders compete with each other, if a single builder has a unique advantage, it will not be incentivized to bid on all of the MEV it can extract.
+Instead, it only needs to outbid the second-best bidder slightly to win.
+We say that the proposer incurs a loss from uncompetitiveness in the first case and a loss from inequality in the second case.
+If proposers suffer significant losses, they would be incentivized to extract MEV by themselves (instead of using the builder market).
+This is undesirable because it re-introduces the dynamics that PBS wants to avoid: sophisticated validators would have an advantage over small ones in extracting MEV.
+In other words, significant proposer losses would make PBS “unstable”.
+Moreover, several MEV mitigation solutions (e.g., MEV-burn [19]) rely on the MEV-Boost auctions as an “oracle” to know how much MEV has been extracted from the system (e.g., MEV-burn proposed to remove that amount from circulation, effectively redistributing them to all holders of ETH through deflation.) The proposer’s loss translates to inaccuracy in the MEV oracle.
+Next, we introduce the auction mechanism and present our method to quantify the proposer losses using auction data.
+Modeling MEV-Boost auctions.
+MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction [6], with a 12-second deadline.
+For each auction instance, builders submit bids in the form of $(B,BV)$ specifying a block $B$ and a bid value (BV) the builder is willing to pay for $B$ to be added to the blockchain.
+The revenue a bidder gains from $B$ is called the true value ($TV(B)$), which can be estimated from transactions in $B$.
+At the end of an auction, the block $B$ with the highest bid is added to the blockchain, the builder pays $BV$ to the proposer and earns $TV$ when $B$ is executed (i.e., the builder’s profit is $TV-BV$).
+Recall that the proposer’s loss is the difference between the auction revenue and the MEV available in a block.
+The latter can be represented by the builders’ highest true values.
+With that, the two types of losses can be computed from auction data, as illustrated in Fig.
+2.
+Specifically, suppose the true values of the bids from a given auction instance are $TV_{1}>TV_{2}\geq\cdots\geq TV_{N}$.
+The loss from inequality is $TV_{2}-TV_{1}$ because a competitive auction can, at most, generate revenue of $TV_{2}$.
+The loss from uncompetitiveness is $\max(TV_{2}-BV,0)$ (if this value is negative, the auction is considered competitive, and there is no loss from uncompetitiveness.)
+Figure 2: Definition of proposer losses.
+Auction data.
+A practical challenge in quantifying the proposer loss in past MEV-Boost auctions is the lack of data.
+Only the winning bids are recorded in the blockchain, but to compute the losses, we also need losing bids.
+The key ingredient that enabled our study is a large-scale auction dataset we curated since 2022: we collected 8.79 billion partial bids [^1] Partial and full bids are to be defined in Sec.
+3.1 of auctions from September 2022 to July 2024; in addition, in collaboration with ultra sound relay [62], we obtained full bids for a total of 301,479 auctions from April 2023 to December 2023.
+From the auction dataset, we can compute the true value of all bids.
+Results.
+91.8% of the MEV-Boost auctions we analyzed were competitive.
+However, auctions tend to be less competitive when the MEV of a slot increases.
+The loss from uncompetitiveness turns out rather insignificant, ranging from 0.5% to 1.7% during the periods of our study.
+To understand the variations in builders’ abilities to extract values, we statistically compared the true values across different builders in the same auction.
+We used several metrics, including the quartile coefficient of dispersion (QCD) [88], to quantify the inequality of true values across builders.
+We found significant differences in the block-building capacities of top, middle, and tail builders.
+The inequality worsens as the MEV of a slot increases, likely because significant MEV opportunities are shared with only a limited number of builders.
+Inequality of block-building capacities has led to a proposer loss ranging from 5.6% to 11.5%.
+I.e., had there been no inequality, the proposer would have earned roughly 10% more.
+This indicates that inequality is the primary driver of the proposer’s losses in the MEV-Boost auctions.
+1.2 What caused Inequality in Block-building Capacities
+Given that proposers suffer significant losses due to the builders’ block-building capacity inequality, we now investigate the underlying causes of this inequality.
+In the current MEV ecosystem (to be presented in Sec.
+2.2), builders receive order flows from order flow providers and compute the optimal ordering of received order flows (see Appx.
+E for a reference builder algorithm).
+Thus, a builder’s block-building capacity is determined by the “quality” of the order flows it receives.
+At the same time, not all order flows are accessible to all builders.
+The difference in builder’s ability to access important private order flows drives the inequality in block-building capacity.
+This also explains builder centralization: only a small set of builders can access these profitable order flows.
+We proposed a metric to identify important order flows, and we investigated their accessibility by different builders and the underlying reasons.
 Identifying pivotal private order flow providers.
-In today’s MEV ecosystem (to be presented in Sec.
-2.2), private order flows come from various private order flow providers.
-Different providers impose different access barriers to their private order flows.
-Some providers (such as MEV-Share (Flashbots, 2024) and MEV Blocker (MEV Blocker, 2023)) have relatively clear criteria a builder must satisfy to access their flows.
-Still, some providers (e.g., most searchers) are almost completely opaque to the public.
-To identify important providers, we define the pivotal level of a provider $P$ as the percentage of MEV-Boost auctions in which the winner would have lost had transactions from $P$ were not available.
-As we will show in Sec.
-4.1, we identified five pivotal providers whose pivotal level exceeds 50% over a period longer than two weeks.
-That is, if a builder cannot access the flows from any of these providers, it will lose the majority of the auctions during the period of our study.
-The good news is that two of these pivotal providers are MEV Blocker (MEV Blocker, 2023) and MEV-Share (Flashbots, 2024), which are open to builders who meet a relatively clear set of criteria.
-The bad news is that satisfying the requirements set by MEV Blocker and MEV-Share is non-trivial for new builders, which we discuss next.
-Quantifying the market entrance barrier.
-Typically, providers share private order flows to reputable builders with a certain market share (Sec.
-4.2).
-However, this presents a chicken-and-egg problem for new builders: to gain market shares, they need private order flows to win the auction, but starting out new, they do not have market share and thus cannot access private order flows.
-To navigate this dilemma, new builders resort to subsidization (Flashbots, 2024): bidding higher than their actual profits in MEV-Boost auctions, to establish the required market shares.
-Therefore a new builder needs to have enough capital to keep subsidizing until reaching a certain market share.
-To understand the significance of this entry barrier, we compute the minimal cost of subsidization required to maintain a 1% market share, which is the requirement of MEV Blocker.
-Specifically, we first calculate the minimal subsidy a builder needs to win an MEV-Boost auction without private order flows; then, we sum up the minimum subsidies over 1% of total auctions per day as the daily cost.
-Our analysis highlights a non-trivial cost for new builders to establish a reputation through subsidization—they need to invest up to 1.4 ETH to subsidize their blocks and secure a minimal 1% market share during the period of our study.
-The cost is comparable to or even higher than the profits of small builders (Flashbots, 2024).
-Trust barriers result in access barriers.
-An important takeaway is that siloes of private order flows stem from the lack of a fair exchange mechanism between providers and builders.
-In the current design, builders are at an advantage and malicious builders can harm providers by, e.g., sandwich attacks (Zhou et al., 2021) and imitation attacks (Qin et al., 2023).
-Therefore, providers only share their order flows with reputable builders, even though such reputation-based protection is pretty weak.
-As entry barriers arise from the trust crisis, reducing or eliminating these trust barriers is essential for making the builder market more decentralized.
-We discuss several ideas in Sec.
-7.2.
-1.2. Competitiveness and Efficiency
-The first question focuses on the evolution of builder markets. In the second question, we focus on the status quo and study the implications of a centralized builder market as we have today.
-MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction (Besanko et al., 2009), with a 12-second deadline.
-English auctions have appealing features: ideally, assuming bidders do not collude and can react to each other instantaneously, the bidders who can extract most MEV will win (efficiency), paying at least the MEV extracted by the second-highest bidder (competitiveness).
-However, given the small number of builders in today’s builder market, it is unclear if MEV-Boost auctions indeed induce the desired dynamics.
-We investigate whether the apparent centralization of the builder market indeed leads to inadequate competition and a loss of efficiency.
-True values in MEV-Boost auctions.
-In MEV-Boost auctions, for a block $B$, a bid value (BV) is the amount the builder is willing to pay for $B$ to be added to the blockchain.
-The total revenue gained by the bidder from $B$ is the true value (TV).
-Rational bidders would bid lower than the true value to get a non-zero profit.
-The true value of a block $B$ can computed from transactions in $B$.
-Inequality in block building capabilities.
-We first compare the true values across different builders for the same auction to understand the variations in builders’ abilities to extract values.
-A builder who consistently demonstrates a superior ability to extract value could dominate the market.
-We used the quartile coefficient of dispersion (QCD) (Wikipedia, 2022) to quantify the inequality of true values across builders.
-Our results in Sec.
-5.1 reveal significant differences in the block-building capabilities of top, middle, and tail builders.
-The inequality worsens as the MEV of a slot increases, possibly because significant MEV opportunities are more likely to be shared with only a limited number of builders.
-Fortunately, the inequality among top builders is not significant, which supports the premise of competitiveness and efficiency.
-More than 88% of auctions are competitive.
-Using the true values, we propose two metrics, competitive index (CI) and efficient index (EI), to quantify whether an auction is competitive or efficient, respectively (see Sec.
-5.2 for details.) Our results reveal that approximately 88.84% of the MEV-Boost auctions we analyzed were competitive.
-However, we also found that the auctions tend to be less competitive when the MEV of a slot increases.
-As discussed previously, the main concern is that a lack of competition can cause loss to proposers.
-We therefore quantify the loss for proposers.
-Our study shows that 16,498 of the 147,926 MEV-Boost auctions were uncompetitive, where proposers incurred total losses amounting to 221.09 ETH, which is 0.98% of the total gains of the proposers.
-Only 79% of auctions are efficient.
-An interesting finding is that more than 20% of MEV-Boost auctions are inefficient and the auctions tend to be inefficient when the MEV in an MEV-Boost auction is low.
-Further analysis reveals that over half of the inefficient auctions are caused by block subsidization.
-The winners of these inefficient auctions overbid to win, although they do not have the highest true value.
-Moreover, block subsidization is more common when the extractable value is small, because the cost will be relatively low.
-Although this may not benefit the overall welfare, it is advantageous for proposers as they earn more than expected.
-Suggestions for future developments.
-Since many identified issues stem from trust concerns, future developments should aim to design mechanisms that reduce or eliminate the reliance on trust.
-Possible directions include community supervision (Flashbots, 2023a) and the use of Trusted Execution Environments (TEEs) (Costan and Devadas, 2016) to ensure the integrity and confidentiality of private order flows (Flashbots, 2024c).
-Another idea is to implement a reliable leakage detection mechanism for private order flows, which we briefly discuss in Sec.
-7.2.
-Such mechanisms would encourage private order flow providers to distribute their flows across multiple builders, thus increasing the equality of the builders’ ability to extract value.
+To formally capture the intuition of “important providers”, we defined pivotal level to quantify the sustained impact of a given provider.
+A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ were not available.
+In other words, pivotal providers are necessary for the winner to win.
+The pivotal level of a provider $P$ is the percentage of MEV-Boost auctions in which $P$ is pivotal.
+Using the auction dataset, we identified five pivotal providers whose pivotal level exceeds 50% over a period longer than two weeks.
+That is, if a builder cannot access the flows from any of these providers, it will lose the majority of the auctions during our study period.
+These order flows are not equally inaccessible to all builders for two reasons we now introduce.
+Inequality due to trust barriers.
+Order flow providers do not openly share private order flows with any builder because a malicious builder can harm the provider by, e.g., sandwich attacks [94] and imitation attacks [71].
+Providers set up different access barriers for self-protection.
+Most MEV searchers (a specific type of order flow providers) resort to “obscurity for security” in that they are not publicly reachable; thus, we could not learn about their access barriers and how to overcome them.
+On the other hand, MEV channels (another type of provider) typically publish a clear access policy, which allowed us to gain some insight into the nature of their access barrier.
+For example, MEV-Blocker only shares order flows with builders with a market share of at least 1%, using market share as a proxy for reputation and trustworthiness.
+This requirement introduces a chicken-and-egg problem for new builders: to gain market share, they need access to private order flows to win auctions, but as newcomers, they lack the required market share to access these flows.
+As a result, new builders often resort to subsidization: bidding higher than their actual profits in MEV-Boost auctions to establish the required market shares.
+Therefore, a new builder needs to keep subsidizing until reaching a certain market share to overcome the access barriers set by MEV channels.
+The cost to do so continues to increase over time and can reach up to 7.4 ETH, according to our quantification.
+In summary, with a moderate cost, builders can overcome the access barrier of one type of order flow provider (MEV channels).
+However, how to access order flows from other providers, particularly important searchers, remains unclear.
+Inequality due to integration.   A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders, causing an inequality in block-building capacities.
+To identify such provider-builder pairs, we introduced builder-specific pivotal level of a provider $P$ for a builder $B$ as the pivotal level of $P$ in auctions won by $B$.
+With this metric, we rediscovered known exclusive relationships (called “integration” [40]), but also found unreported integrated providers for top builders (Titan and Beaver respectively).
+This finding is interesting because Titan has been the only top builder without a well-known integrated provider.
+Our findings show that all three top builders, accounting for 90% of market share, have integrated providers.
+Integration poses a market entry barrier as new builders likely lack integrated providers.
+Moreover, it offers a unique competitive edge and will likely cause dominance.
+Finally, there is a strong incentive to integrate: compared to sending valuable order flows to multiple builders, sending them to a single builder enables this builder to win the auction with higher profit; the provider and the builder can share the profit, which means integration improves their total utility.
+Integration is detrimental to proposers, as it will reduce their profits. The strong incentive for integration makes it increasingly difficult to address builder centralization.
+1.3 Mitigation Ideas
+To summarize, we found that trust barriers and incentives for integration contribute to builder centralization, leading proposers to incur significant losses.
+The consequences include the instability of PBS and the inaccuracy of MEV oracles.
+We propose a structural change to the existing MEV supply chain (Sec.
+2.2) and a solution based on the new supply chain structure.
+However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
+The new supply chain has only one builder to which the order flow providers send flows and from which the proposers accept blocks.
+In this setting, a recent work [57] suggests a truthful mechanism that incentivizes order flow providers to report the MEV in their order flows truthfully.
+Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15], hardware-based enclaves that guarantee the confidentiality and integrity of programs running inside.
+Order flow providers verify that the correct mechanism is running in the TEE using remote attestation, then send order flows to the TEE along with a bid indicating the MEV in their order flow.
+Then, the TEE computes a fair refund to order flow providers.
+The refund mechanism is designed so that bidding truthfully is a dominant strategy [57].
+(While oversimplifying, one can make a loose analogy to sealed-bid second-price auctions among order flow providers.) Then, the TEE can compute the total MEV in a block from the providers’ bids, solving the problem of inaccurate MEV oracles.
+This also addresses the instability of PBS because the builder program in the TEE can be programmed to pay the proposer fairly, leaving just enough for its operation.
+Since this builder is fair to order flow providers, it is conceivable that it will be preferred by order flow providers over other builders.
+However, as mentioned above, it is unclear if having a single builder is an equilibrium.
+Also, the proposers may be incentivized to become order flow providers, which may re-introduce the concern similar to PBS instability.
+We elaborate on the idea in Sec.
+7.1.
 Contribution
 To the best of our knowledge, we are the first to use the rich auction data to analyze the decentralization of Ethereum’s builder market, while most previous works are limited to on-chain data.
-Our dataset of the true values of 147,926 MEV-Boost auctions from April to August 2023 may be of independent interest.
-Our dataset of 5.57 billion partial bids since 2022 is the largest-scale known to us.
-We characterized the efficacy of PBS in terms of being open to new builders and inducing competitive and efficient auctions. Our key findings are highlighted in Tab. 1, which we summarize below.
-  • Sec.
-4: Private order flows contribute to more than 60% of the MEV in over 50% of blocks every day since July 2023.
-We identified five pivotal private order flow providers that had a significant and sustained impact on the winners of more than half of the MEV-Boost auctions.
-Two of them are MEV Blocker and MEV-Share.
-  • Sec. 4.2: Private order flow providers impose market share requirements on builders, which costs new builders up to 1.4 ETH in subsidy to enter the market during the period of our study.
-  • Sec. 5.1: The QCD shows a significant inequality in builders’ abilities to extract value across top, middle, and tail builders, but the inequality among the top builders is not significant.
-  • Sec.
-5.2: We find that 88.84% of the 147,926 MEV-Boost auctions are competitive.
-However, proposers incur losses of 221.09 ETH from the uncompetitive auctions, representing 0.98% of their total gain.
-These losses can still threaten the stable separation between proposers and builders.
-  • Sec.
-5.2: 79.74% of the MEV-Boost auctions in our study are efficient.
-A significant factor contributing to inefficiency is block subsidization, where winners overbid to secure the auctions, possibly driven by the need to meet market share requirements.
+We summarize our contribution below.
+  • We found that builder centralization has led to significant proposer losses, which could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
+We challenge the community’s belief that “builder centralization is okay".
+  • We attribute proposer losses to two factors: auctions are not competitive, builders’ block-building capacities are unequal, and inequality is the primary cause.
+  • We identify two reasons for inequality in block-building capacity: providers’ trust concerns and the incentive for integration.
+  • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest.
+Our dataset of 8.79 billion partial bids since 2022 is the largest scale known to us.
   • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions.
-Currently, there is only limited auction theoretic analysis that considers the complexity of real-world MEV auctions (Gupta et al., 2023; Pai and Resnick, 2023; Wu et al., 2024).
-The findings in this paper can serve as empirical evidence for future theoretic analysis of MEV-Boost auctions.
-Figure 2. Illustration of the MEV supply chain in the PBS scheme.
-2. Background
-2.1. MEV-Boost Auctions
+Currently, there is only limited auction theoretic analysis that considers the complexity of real-world MEV auctions [40, 67, 90].
+The findings in this paper can serve as empirical evidence for future theoretical interrogation.
+Figure 3: Illustration of the MEV supply chain in the PBS scheme.
+2 Background
+2.1 MEV-Boost Auctions
 MEV.
-Miner/Maximal Extractable Value (MEV) (Daian et al., 2020) refers to the profit that privileged players (e.g., validators) can earn by including, excluding, and reordering the transactions.
+Miner/Maximal Extractable Value (MEV) [16] refers to the profit that privileged players (e.g., validators) can earn by including, excluding, and reordering the transactions.
 MEV poses a threat to decentralization, as explained in the introduction.
 There are two schools of thought on mitigating the negative impacts of MEV.
-The first one is referred to as MEV democratization (Flashbots, 2022b), the main idea of which is to facilitate MEV extraction and level the playground.
-MEV-Boost (Flashbots, 2024), to be introduced below, is the primary example.
-Another approach is to enforce certain ordering policies (e.g., FIFO) to prevent order manipulation, such as (Kelkar et al., 2020, 2023).
+The first one is referred to as MEV democratization [30], the main idea of which is to facilitate MEV extraction and level the playground.
+MEV-Boost [35], to be introduced below, is the primary example.
+Another approach is to enforce certain ordering policies (e.g., FIFO) to prevent order manipulation, such as [45, 44].
 We focus on MEV-Boost because it is widely used.
 PoS Ethereum.
-Before introducing background on MEV-Boost auctions, we first cover the relevant concepts in Ethereum’s Proof of Stake-based consensus protocol called Gasper (Buterin et al., 2020).
+Before introducing background on MEV-Boost auctions, we first cover the relevant concepts in Ethereum’s Proof of Stake-based consensus protocol called Gasper [13].
 Participants of Gasper are called validators.
-Each validator must deposit 32 ETH as collateral (Ethereum Foundation, 2024a).
+Each validator must deposit 32 ETH as collateral [22].
 Time is divided into 12-second intervals called slots.
 This results in 7,200 slots in Ethereum per day.
-During each slot, a single validator, referred to as the proposer, proposes a block, while other validators vote for the proposed block if it extends the chain head in their local view.
+During each slot, a single validator, referred to as the proposer, proposes a block, while other validators vote for the chain head in their local view.
 Proposer-builder separation (PBS) and MEV-Boost.
-Proposer-builder separation (PBS) can be conceptually viewed as a sidecar to the core consensus protocol (although some proposals are more tightly coupled with consensus) to allow the proposer to outsource the job of building blocks to builders.
-MEV-Boost (Flashbots, 2024) is the current incarnation of the PBS idea.
-About 90% of all blocks on Ethereum are built by MEV-Boost (Wahrstätter, 2024).
-The basic idea is to auction off the right to propose a block to the builder who bids the highest in an open-bid ascending price auction (contributors, 2024).
-Because builders and proposers are mutually untrusted, the auction is facilitated by a (trusted) third party called a relay (Flashbots, 2022a).
-Here are the high-level steps (we refer readers to (Flashbots, 2024) for details):
+PBS can be conceptually viewed as a sidecar to the core consensus protocol (although some proposals are more tightly coupled with consensus) to allow the proposer to outsource the job of building blocks to builders.
+MEV-Boost [35] is the current incarnation of the PBS idea.
+About 90% of all blocks on Ethereum are built by MEV-Boost [87].
+The basic idea is to auction off the right to propose a block to the builder who bids the highest in an open-bid ascending price auction [14].
+Because builders and proposers are mutually untrusted, the auction is facilitated by a (trusted) third party called a relay [29].
+Here are the high-level steps (we refer readers to [35] for details):
   • Builders submit bids to relays throughout the slot.
-For the purpose of this paper, we denote a bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key.
+For this paper, we denote a bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key.
 Bids are signed by the builder but we omit the signature for clarity.
@@
   • The proposer selects a block hash $h_{B}$, typically the one with the highest $BV$, and sends the relay a signature on $h_{B}$.
-This signature is a promise to propose $B$ in the consensus protocol (if it proposes another block, this signature is evidence of equivocation and will cause the proposer to be penalized (Buterin et al., 2020).) After verifying the signature, the relay returns $B$ to the proposer.
+This signature is a promise to propose $B$ in the consensus protocol (if it proposes another block, this signature is evidence of equivocation and will cause the proposer to be penalized [13].) After verifying the signature, the relay returns $B$ to the proposer.
 Two additional remarks.
-First, relays are trusted entities, but anyone can run a relay (currently there are ten (EthStaker Community, 2024)), and builders and proposers can connect to the relays they trust.
+First, relays are trusted entities, but anyone can run a relay (currently ten [27]), and builders and proposers can connect to the relays they trust.
 Second, winning bids are recorded on-chain, and losing bids are discarded after the auction.
@@
 3.
-2.2. MEV Supply Chain
+2.2 MEV Supply Chain
 The ecosystem around discovering and extracting MEV forms an MEV supply chain, as depicted in Fig.
-2.
+3.
 Builders, relays, and proposers (validators) have been introduced above.
 Now we introduce the “upstream” of builders.
-Wallet.
+Wallet, users, trading bots.
 Users use wallets to interact with blockchains.
@@
 MEV may be generated during this process, e.g., when the user transaction creates an arbitrage opportunity.
-While the wallet is typically controlled by the users, it can also be managed by a third party, such as Telegram trading bots (Intelligence, 2024).
-These bots are automated programs integrated with the Telegram app (Telegram, 2024), automatically conducting cryptocurrency trades on behalf of users.
+While the wallet is typically controlled by the users, it can also be managed by a third party, such as Telegram trading bots [43].
+These bots are automated programs integrated with the Telegram app [79], automatically conducting cryptocurrency trades on behalf of users.
 Users’ transactions eventually reach builders or proposers, either through the peer-to-peer network (the mempool) or through third-party services called private channels.
@@
 A private channel, or channel for short, connects users directly with builders, bypassing the mempool.
-Channels receive transactions from users and then, forward them to the builders specified by the users, offering privacy and atomicity guarantees.
+Channels receive transactions from users and then, forward them to the user-specified builders, offering privacy and atomicity guarantees.
 E.g., a user can send a bundle (list) of transactions and request them not to be unbundled.
-Examples include MEV-Share (Flashbots, 2024), MEV Blocker (MEV Blocker, 2023), BloXroute (bloXroute Labs, 2024), etc.
+Examples include MEV-Share [37], MEV Blocker [60], BloXroute [9], etc.
 Searcher.
@@
 Some searchers may prefer channels for the privacy and atomicity benefits.
-A search is said to integrate with a builder if it sends almost all of its orders to that builder exclusively.
+A searcher is said to integrate with a builder if it sends almost all of its orders to that builder exclusively.
 For example, searcher 0xA69b… is considered an integrated searcher of Beaver Builder because Beaver Builder includes the majority of its transactions.
@@
 Private order flow.
-Order flow refers to a source of transactions that may carry extractable value and can be categorized into two categories: public order flow, originating from the public mempool, and private order flow, sent directly by users or searchers without being broadcast to the public mempool.
-The entities that can provide private order flows to the builders are referred to as private order flow providers, which includes searchers, channels, users, and Telegram trading bots as shown in Fig.
-2.
-2.3. Model and Assumptions.
+Order flow refers to a stream of transactions that may carry extractable value and can be categorized into two categories: public order flow from the public mempool, and private order flow, sent directly by users or searchers without being broadcast to the mempool.
+The entities that can provide private order flows to the builders are referred to as private order flow providers, which include searchers, channels, users, and Telegram trading bots as shown in Fig.
+3.
+2.3 Model and Assumptions.
 Accounts and states.
-We follow the model of (Babel et al., 2023a, b).
+We follow the model of [2, 3].
 We use $\mathcal{A}$ to denote the space of all possible accounts.
@@
 Assuming $p$ controls a set of accounts $A_{p}$, the extracted value of $B$ for $p$ is:
-(1) $EV_{p}(B)=\sum_{a\in A_{p}}\text{balance}_{s^{\prime}}(a)[0]-\text{balance}_{s}(a)[0]$
+$EV_{p}(B)=\sum_{a\in A_{p}}\text{balance}_{s^{\prime}}(a)[0]-\text{balance}_{s}(a)[0]$ (1)
 where $s^{\prime}=action(B)(s)$.
@@
 We assume a builder $p$’s income from building a block $B$ is $EV_{p}(B)$, or $EV(B)$ for short since the builder identity is usually encoded in the block.
-We compute EV following a similar approach as that in MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023).
-Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees (Ethereum, 2024) and direct transfers (Flashbots, 2024a) to the builder’s addresses minus the refund to other addresses.
+We compute EV following a similar approach as that in MEV analysis tools [20, 31] and previous works [41, 80].
+Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees [21] and direct transfers [32] to the builder’s addresses minus the refund to other addresses.
 We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
@@
 7.
-3. Dataset
-We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to March 2024, and a dataset of full bids from April to August 2023.
-The full bids are obtained from the ultra sound relay (Money, 2023).
+3 Dataset
+We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to July 2024, and a dataset of full bids from April to December 2023.
+The full bids are obtained from the ultra sound relay [62].
 Recall from Sec.
 2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc.
-To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. 2 summarizes the datasets.
+To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below.
+Tab.
+1 summarizes the datasets.
 Dataset Content # of entries Source
-On-chain blocks 4,019,895 Reth node
-transactions 606,484,921
-Auction blocks produced by MEV-Boost 3,517,851 relay API
-partial bids 5,577,318,196 relay API
-full bids from ultra sound 164,700,566 ultra sound relay
-Misc. builder public keys 191 see Sec. 3.1
-builder addresses 125
-searcher addresses 2,069
-private transactions 49,081,209
-transaction sources 33,010,466
-Table 2. Overview of collected datasets.
-3.1. Data Collection and Validation
+On-chain blocks 4,892,579 Reth node
+transactions 748,859,795
+Auction blocks produced by MEV-Boost 4,305,571 relay API
+partial bids 8,789,557,915 relay API
+full bids from ultra sound 326,574,676 ultra sound relay
+Misc. builder public keys 248 see Sec. 3.1
+builder addresses 102
+searcher addresses 2,140
+private transactions 77,747,623
+transaction sources 60,345,066
+Table 1: Overview of collected datasets.
+3.1 Data Collection and Validation
 On-chain data.
-We run an Ethereum execution client Reth (paradigmxyz, 2024) along with a consensus client Lighthouse (Prime, 2024) to collect blocks and transactions from September 15, 2022, to March 31, 2024.
-The on-chain dataset includes 4,019,895 blocks and 606,484,921 transactions.
+We run an Ethereum execution client Reth [68] along with a consensus client Lighthouse [69] to collect blocks and transactions from September 15, 2022, to July 31, 2024.
+The on-chain dataset includes 4,892,579 blocks and 748,859,795 transactions.
 Data from relay APIs.
-We connected to all 13 known relays333three relays were inactive during the period of our study and collected the following information from their public APIs: 1) the winning bid including the final block sent to the proposer; and 2) partial bids for blocks that did not win the auction.
+We connected to all 13 known relays [^2] Three relays were inactive during the period of our study.
+and collected the following information from their public APIs: 1) the winning bid including the final block sent to the proposer; and 2) partial bids for blocks that did not win the auction.
 Recall that partial bids do not include the block body.
-We collected 3,517,851 winning bids and 5,577,318,196 partial bids from September 15, 2022, to March 31, 2024, in a total of 3,810,630 auctions.
+We collected 4,305,571 winning bids and 8,789,557,915 partial bids from September 15, 2022, to July 31, 2024.
 We use winning bids to identify blocks produced through MEV-Boost.
 We use partial bids to quantify the importance of private order flow providers in Sec.
-4.
-To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets (Network, 2023; Wahrstätter, 2023; DataAlways, 2024).
-Our dataset includes 5,382 more winning bids than other datasets combined.
+5.
+To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets [64, 85, 17].
+Our dataset includes 720 more winning bids than other datasets combined.
 Upon comparison with on-chain blocks, we find they are all included on Ethereum, thereby indicating a high level of completeness.
 Full bids from the ultra sound relay .
-From the ultra sound relay (Money, 2023), we obtained the full bids in 147,926 historical auctions that took place between April and August 2023.
-Since the data volume is large (we record roughly 200 GB data per day), our dataset samples one week of each month: April 9-15, May 1-7, June 1-7, July 1-7, and August 1-7, 2023.
-The dataset consists of 164,700,566 full bids (block, bid value, and builder public key).
+From the ultra sound relay [62], we obtained the full bids in 301,479 historical auctions that took place between April and December 2023 (ultra sound relay stopped collecting bids at the end of 2023).
+Since the data volume is large (we record roughly 200 GB of data per day), our dataset samples one week from each month: April 9-15 and May 1-7 to December 1-7, 2023.
+The dataset comprises 326,574,676 full bids (block, bid value, and builder public key).
 We use full bids in Sec.
-5 to quantify the competitiveness and efficacy of MEV-Boost auctions.
+4 to quantify the competitiveness and efficiency of MEV-Boost auctions.
 The original data from relays includes bids submitted after the deadline because of network latency.
@@
 On April 18, 2023, MEV-Boost introduced bid cancellation that allows a bidder to cancel a previous bid by submitting a lower one.
-However, proposers can query bids before they are canceled and store them locally (Ethereum Research, 2023), effectively bypassing cancellation.
-Therefore we do not take cancellation into consideration.
+However, proposers can query bids before they are canceled and store them locally [24], effectively bypassing cancellation.
+Therefore we do not consider cancellation.
 Not all builders connect to the ultra sound relay, so we need to validate that the full bids from ultra sound relay cover a significant subset of all builders.
@@
 We use private transaction hashes to identify private order flows and use the searchers’ addresses and sources of transactions to identify the private order flow providers in Sec.
-4.
-  • Builder public keys: We collect public keys controlled by each builder combining two data sources.
-First, some builders publish their public keys in official documents or code repositories.
-For example, Titan builder has disclosed 12 public keys in its official document (Builder, 2024).
-Second, some builders leave distinct marks in their blocks’ “extra field”.
-For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’.
-Given that such marks are unauthenticated and susceptible to impersonation, we ignore them unless they are used more than 100 times in the bids by the same builder public key.
-    To validate the completeness of our dataset, we manually cross-check it with public datasets (Rated Network, 2023; Network, 2024) and related works (Heimbach et al., 2024, 2023), and the results show that our dataset includes 62 more public keys, so our dataset is the most complete one so far.
-  • Builder addresses: We collect the builder addresses from the blocks’ last transactions combined with the public dataset (Etherscan, 2024).
-Similarly, to validate their completeness, we manually cross-check them with related works (Heimbach et al., 2024, 2023), and the results show that our dataset includes 113 more builder addresses, which is the most complete so far.
-  • Searcher addresses: We obtained known searcher addresses from Etherscan (Etherscan, 2024), libMEV (libMEV, 2024), and related work (Heimbach et al., 2024).
-  • Private transaction hashes: We use data from Mempool Guru (Yang et al., 2022) to label transactions that bypassed the public mempool as private.
+5.
+  • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx.
+B for details).
+    To validate the completeness of our dataset, we manually cross-check it with public datasets [72, 65] and related works [42, 41], and the results show that our dataset includes 119 more public keys, so our dataset is the most complete one so far.
+  • Builder addresses: We collect the builder addresses from the blocks’ last transactions combined with the public dataset [26].
+Similarly, to validate their completeness, we manually cross-check them with related works [42, 41], and the results show that our dataset includes 86 more builder addresses, which is the most complete so far.
+  • Searcher addresses: We obtained known searcher addresses from Etherscan [25], libMEV [50], and related work [42].
+  • Private transaction hashes: We use data from Mempool Guru [92] to label transactions that bypassed the public mempool as private.
 We believe these labels have already been cross-validated since the private transactions are collected by seven distributed nodes.
   • Transaction sources: we are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through.
-    We identified transactions originated from Telegram trading bots, including Maestro (Maestro, 2023a), Banana Gun (Banana Gun, 2023), and Unibot (Unibot X, 2023), by searching for transactions that interacted with their official smart contracts.
-    Two channels, MEV Blocker and MEV-Share publish data (Protocol, 2024; Flashbots, 2024b) about transactions submitted to them.
-Specifically, transactions submitted to MEV Blocker are publicly available on Dune (Protocol, 2024), from which we extracted 14,901,303 transaction hashes that fall within our study’s range.
-We obtained 13,992,371 transactions submitted to MEV-Share from the public Flashbots dataset (Flashbots, 2024b).
-These transaction labels are sourced from their official datasets and we view them as ground truth.
-4. Openess of the builder market
-The builder market is technically open to everyone, but the barrier to accessing profitable private order flows (Sec. 2.2) forms an implicit access barrier.
+We identified transaction sources using ground truth data (as described in Appx.
+B).
+4 Implications of Builder Centralization
+In this section, we focus on the status quo and study the implications of a centralized builder market as we have today.
+In particular, we are interested in the losses incurred by proposers, as significant losses could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
+We quantify proposers’ losses in historical MEV-Boost auctions considering two factors: the competitiveness of the auctions and the inequality in block-building capacity.
+Our study includes some MEV-Boost auctions where the highest bids were not included on Ethereum due to network latency.
+However, we do not omit these auctions since the auctions themselves remain meaningful.
+4.1 Competitiveness of MEV-Boost Auctions
+Ideally, revenue from MEV-Boost auctions should be near optimal due to competition.
+However, whether real-world auctions induce such near-perfect competition is complicated by several confounding factors: builders may not be able to respond to bids due to network latency, or they may collude implicitly or explicitly.
+In this section, we quantify the competitiveness of the MEV-Boost auctions using what we call competitive index (CI).
+We also look at the allocation efficiency as it evaluates whether the current mechanism induces the desired outcomes.
+True value.
+In the context of auctions generally, the true value of the item being auctioned refers to its actual worth to the bidder.
+Recall that in MEV-Boost auctions (Sec.
+2), a builder submits a block $B$ (a list of transactions) and a bid value $BV$ that the builder is willing to pay to the proposer if $B$ is proposed.
+The true value of $B$, denoted $TV(B)$, is typically higher than $BV$.
+One can think of $TV$ as the builder’s revenue and $BV$ as the cost.
+The difference $EV:=TV-BV$ is the profit pocketed by the builder, which we call extracted value (EV) — i.e., $EV(B)$ is the difference of the builder’s ETH balance after executing transactions in $B$.
+For example, suppose executing $B$ yields a net profit of $0.1$ ETH after paying $0.9$ ETH in bid value in the auction, then the true value (the revenue) is $1.0$ ETH.
+To compute $TV(B)$ of a given block $B$, we follow the standard approach in open source MEV analysis tools [20, 31] and previous works [41, 80], though we built our own analysis toolchain and optimized its performance to handle the large volume of auction data.
+We compute $EV(B)$ using a modified Ethereum execution client (Reth [68]) to execute transactions in $B$ and compute the net ETH balance changes of the bidder’s addresses.
+Most builders’ addresses are well-known partly because they need to build up a reputation.
+We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec.
+3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
+In every slot $s$, bidders can submit multiple bids throughout the duration of the slot.
+We use $BV(p,s)$ to denote the highest bid value from bidder $p$ in slot $s$, and $TV(p,s)$ the corresponding true value.
+Competitive index.
+In a given slot $s$, suppose builders $p_{1},\cdots,p_{n}$ are ordered by their true value from high to low, i.e., $TV(p_{1},s)\geq\dots\geq TV(p_{n},s)$, $BV_{w}(s)$ is the bid value of the winner and $TV_{w}(s)$ is the corresponding true value, the competitive index of slot $s$, $CI(s)$ is defined as:
+$CI(s)=\frac{BV_{w}(s)-TV(p_{2},s)}{TV(p_{2},s)}\times 100\%.$
+Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value.
+$CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, satisfying the definition of a competitive auction, while $CI(s)<0$ means insufficient competition.
+Efficient index.   Same as above, the efficient index, $EI(s)$, is defined as:
+$EI(s)=\frac{TV_{w}(s)-TV(p_{2},s)}{TV(p_{2},s)}\times 100\%.$
+$EI(s)>0$ indicates that the winner has the highest true value, whereas $EI(s)\leq 0$ implies that the bidder with the highest true value lost the auction.
+Figure 4: The distribution of CI and EI, where each dot represents an auction.
+The color represents density when multiple dots overlap.
+91.8% of the MEV-Boost auctions are competitive (sum of the slots having $CI\geq 0$), and 75.1% of them are efficient (sum of the slots with $EI>0$).
+Dots on the diagonal line ($CI=EI$) are auctions where the winners bid the true values.
+The dots above this line are auctions where the winners bid lower than true values.
+Analysis of uncompetitiveness and inefficiency.
+Fig.
+4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the ultra sound relay dataset.
+Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
+We note Fig.
+4 shows a somewhat strict definition of competitiveness (i.e., $CI\geq 0$).
+In reality, builders may bid slightly lower than their true value to ensure profits.
+Therefore, it is reasonable to relax and consider an auction with $CI\geq-\delta$ to be competitive for a small $\delta$.
+We call this notion $\delta$-competitive.
+As shown in Fig.
+13, when we consider $\delta=1\%$ or $\delta=2\%$, the percentage of uncompetitive auctions reduces to about 6% and 5%, respectively.
+That is, under the slightly relaxed definition, approximately 95% auctions were competitive.
+In the interest of space, we leave the analysis of efficiency and the implication of MEV on efficiency to Appx. D.1.
+MEV’s implication on auction outcomes.
+To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we categorize every slot into one of three tiers according to the winning bid (as a proxy to the significance of the MEV opportunity in that slot): low MEV (below 0.0215 ETH, the 10th percentile of the winning bid values), high MEV (above 0.3933 ETH, 95th percentile) and medium MEV in between.
+The ratio of competitive auctions in each tier is shown in Fig.
+5.
+We note that competitiveness decreases as the MEV of a slot increases.
+Figure 5: The ratio of auctions that are competitive and efficient across various MEV tiers.
+Finding I: MEV-Boost auctions are generally competitive, while the competitiveness of MEV-Boost auctions deteriorates as a slot’s MEV increases.
+Loss from uncompetiveness.
+For an uncompetitive auction, the loss from uncompetitiveness is the difference between the second-highest true value and the actual winning bid value, representing the loss a validator incurs compared to what it would have received in a competitive auction.
+We computed the loss from uncompetitiveness using our dataset, and the results are shown in Tab.
+2.
+The loss from uncompetitiveness ranges from 0.5% to 1.7% during the periods covered.
+Table 2: Losses from uncompetitiveness and inequality.
+Time Slots Profits (ETH) Losses-un (ETH) (%) Losses-in (ETH) (%)
+April 9-15 28,385 2,704.4 46.9 (1.7) 312.1 (11.5)
+May 1-7 30,300 9,331.7 115.8 (1.2) 518.6 (5.6)
+June 1-7 35,443 4,341.8 25.1 (0.6) 342.2 (7.9)
+July 1-7 36,040 3,938.8 19.1 (0.5) 246.1 (6.3)
+August 1-7 17,831 2,135.5 12.5 (0.6) 146.6 (6.9)
+September 1-7 35,815 3,939.5 48.5 (1.2) 247.7 (6.3)
+October 1-7 39,131 3,427.0 28.9 (0.8) 234.1 (6.8)
+November 1-7 39,848 4,736.2 45.6 (1.0) 305.0 (6.4)
+December 1-7 38,686 5,884.4 80.0 (1.4) 365.8 (6.2)
+4.2 Inequality in Block-building Capacities
+The previous section shows that the MEV-Boost auctions are competitive in most cases.
+However, a proposer may still incur loss due to a significant block-building capacity gap between builders.
+In MEV Boost auctions, this quantity is captured by the true values of the builder’s bids.
+Quantifying the inequality of block-building capacities.
+We compare the highest true values of different builders in the same slot to understand the differences in builders’ ability to extract values.
+If a builder consistently has the highest true value, then she can potentially dominate the builder market.
+On the contrary, if many builders have similar true values, then it is plausible that a properly designed market could avoid a monopoly.
+To quantify the disparity of block-building capacity, we compute the quartile coefficient of dispersion (QCD) [88] of the highest true values in a given slot.
+Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$.
+Let $Q_{1}(\vec{x})$ and $Q_{3}(\vec{x})$ representing the first and third quartiles of $\vec{x}$ (i.e., 25% and 75% percentiles).
+The QCD of slot $s$ is
+$QCD_{s}=QCD(\vec{tv}_{s})=\frac{Q_{3}(\vec{tv}_{s})-Q_{1}(\vec{tv}_{s})}{Q_{3}(\vec{tv}_{s})+Q_{1}(\vec{tv}_{s})}.$
+QCD is a real value in $[0,1]$.
+A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$).
+Since it uses quartiles, QCD is robust to outliers [88].
+We now present the result given by QCD, noting that other metrics lead to a similar conclusion (see Sec. D.2).
+Evaluation results.
+We categorize the builders into three groups based on their market share from April to December 2023: top builders (top 5), middle builders (6-15), and tail builders (16-25).
+To understand how the magnitude of MEV in a slot affects the block-building capability, we follow the same methodology in Sec.
+4.1 and categorize the auctions into three tiers: low MEV, medium MEV, and high MEV.
+In Fig.
+6, we plot the distribution of the QCD for MEV-Boost auctions from April 2023 to December 2023 for different builder groups and MEV tiers.
+(We sample the first week of each month.
+See Sec.
+3.1 for details.) It reveals two trends.
+First, within each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail.
+This suggests that while top builders have comparable block-building capacities, middle and tail builders have significantly different capacitiesx.
+This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of the builder market.
+Second, the inequality worsens as the MEV tier increases.
+One possible reason is that providers may give significant MEV opportunities to select builders, as we will discuss in Sec.
+5.3.
+Figure 6: Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders.
+The QCD measures the inequality of block-build capacity across builders.
+This figure reveals that such inequality worsens as MEV increases.
+Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
+Finding II: The inequality in block building is lowest among top builders, higher among middle builders, and most pronounced among tail builders, and the inequality worsens as a slot’s MEV increases.
+Loss from inequality.
+For each auction, the loss from inequality is the difference between the highest and second-highest true values, reflecting the potential loss a validator may face due to unequal block-building capacities, as a competitive auction only ensures the winning bid slightly exceeds the second-highest true value.
+As shown in Tab.
+2, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset.
+Comparing the two types of loss, we find that inequality is the proposers’ primary source of loss.
+Note that the time range of the full bids is from April to December 2023, and the inequality in the current builder market may worsen (as discussed in Sec.
+5.3).
+5 What Caused Inequality in Block-building Capacities
+Given the significant losses proposers incur due to the inequality in builders’ block-building capacity, we further investigate the causes of this inequality in this section.
+Although the builder market is technically open to all, the barrier to accessing profitable private order flows (Sec. 2.2) forms an implicit access barrier.
 To illustrate the importance of private order flow to the success of builders, Fig.
 7 plots the fraction of builder income from private order flows.
-Currently, approximately 60% of the block value comes from private order flows.
-Figure 3. Fraction of builder income from private order flows.
-Finding I: The private order flow is crucial for builders to win the MEV-Boost auction because it contributes to more than 60% of the block value in over 50% of blocks.
+Currently, approximately 80% of the block value comes from private order flows.
+Figure 7: Fraction of builder income from private order flows.
+Finding III: The private order flow is crucial for builders to win the MEV-Boost auction because it contributes to more than 80% of the block value in over 50% of blocks.
 Where can new builders get private order flows?
@@
 Different providers impose different access barriers to their private order flows.
-Some providers are open to new builders who satisfy certain requirements (such as MEV Blocker or MEV-Share), while some (e.g., most searchers) are almost completely opaque to the public.
-To understand if the builder market is open to new builders, we first identify important “pivotal” providers and then quantify the cost for a new builder to access private order flows from representative pivotal providers.
-4.1. Pivotal Private Order Flow Providers
+Some providers, like MEV Blocker or MEV-Share, are open to new builders who satisfy certain requirements, while some (e.g., most searchers) are almost opaque to the public.
+To understand what causes the access barrier for new builders, we first identify important “pivotal” providers and then analyze the requirements to access private order flows from these representative pivotal providers.
+5.1 Pivotal Private Order Flow Providers
 Prominent private order flow providers are listed in Tab.
-6.
+5.
 Based on their roles in the MEV supply chain (see Fig.
-2), we categorize them into three types: searcher, Telegram trading bot, and channel.
+3), we categorize them into three types: searcher, Telegram trading bot, and channel.
 We now present our method to quantify the importance of individual providers.
-An existing approach (Lu, 2024; Lu, Angela and Sui, Danning and Durnford, Jaden, 2024) is to compute the contribution of each provider to the builders’ income (Sec.
+An existing approach [53, 54] is to compute the contribution of each provider to the builders’ income (Sec.
 2.3) during a specific period.
 However, this approach only considers the magnitude of impact but neglects the duration (in terms of time or the number of auctions).
-For example, a single private transaction, such as one444 0xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip, could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot.
-Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that can affect half of the MEV-Boost auctions in over two weeks.
-This captures their sustained impact on the system.
+For example, a single private transaction, such as one3330xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip, could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot.
+Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that can affect half of the MEV-Boost auctions in over two weeks, thereby capturing their sustained impact.
 We define a provider $P$ as pivotal for an instance of MEV-Boost auction if removing $P$’s transactions from the winning block causes the winner to lose the auction.
@@
 Note that there can be multiple pivotal providers for an auction.
+Figure 8: A provider $P$ is pivotal if the winner of an auction could have lost without order flows from $P$.
+The pivotal level of top-5 pivotal providers over time.
+A line may not start from the beginning because the provider did not exist at that point.
+Maestro became an independent provider around Dec 2023.
 For each pivotal provider, we compute the fraction of auctions in which it is pivotal.
@@
 8 plots the pivotal level of the top-5 providers.
-We can make a few observations from Fig.
-4.
-First, a particular searcher, jaredfromsubway.eth, has been gaining prominence since February 2023 and reached its peak significance in May 2023.
-During this period, the winners of more than 70% of the MEV-Boost auctions would have failed had they not received order flow from jaredfromsubway.eth.
-Its influence has gradually decreased since then but still remains pivotal in 20%-40% of auctions since Nov 2023.
-Besides, the order flow from Telegram trading bots has become vital for builders with the rise of such bots.
-For instance, Banana Gun is pivotal in about 40% of the MEV-Boost auctions in September 2023.
-Figure 4.
-A provider $P$ is pivotal if the winner of an auction could have lost without order flows from $P$.
-The pivotal level of top-5 pivotal providers over time.
-A line may not start from the beginning because the provider did not exist at that point.
-Maestro became an independent provider around December 2023.
-Second, we can observe a positive trend from Fig.
-4 that MEV Share and MEV Blocker have been increasingly pivotal since May 2023.
-These two providers positively impact decentralization because they have relatively clear requirements for accessing their order flows (the requirements of different providers will be discussed in Sec.
-4.2).
-Third, we observe a notable downtrend in the pivotal level of MEV-Share and MEV Blocker starting from November 2023, accompanied by an uptrend in Maestro, a popular Telegram trading bot (Maestro, 2023a), beginning in December 2023.
-This is because Maestro used to be an important upstream of MEV-Share and MEV Blocker but stopped in November 2023, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
-Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned.
-For example, Mastro may prefer to send flows to builders directly for a better price than what MEV-Share and MEV Blocker can offer.
-It is alarming to see that upstream providers can significantly swing the significance of platforms.
-Finding II: We identified five pivotal providers (MEV-Share, MEV Blocker, jaredfromsubway.eth, Banana Gun, Maestro) who had a sustained influence on the outcome of more than half of MEV-Boost auctions.
-A positive trend towards decentralization is observed as the pivotal level of the searcher jaredfromsubway.eth decreased, while the pivotal level of MEV-Share and MEV Blocker increased.
-However, their reliance on upstream providers like Maestro is alarming.
-Builder-specific Pivotal Level.
-We compute each provider’s builder-specific pivotal level to understand the relationship between providers and builders.
-For a given provider $P$, its pivotal level specific to builder $B$ is the pivotal level of the MEV-Boost auctions where $B$ is the winner.
-Fig.
-5 plots selected providers’ builder-specific pivotal levels for the top-3 builders, ranked based on their weekly market share in April 2024.
-We include providers whose weekly pivotal levels exceed 25% for any of the three builders over two months to illustrate these builder-specific pivotal levels.
-We note a clear variation in the importance of providers for different builders.
-For example, the significance of MEV Blocker and Banana Gun to Titan Builder is greater than they are to other builders.
-Furthermore, we observe that some providers have a pivotal level exceeding 50% only with specific builders, suggesting a possible collaborative relationship.
-The searcher 0xA69b… consistently influences over 50% of the MEV-Boost auctions won by Beaver Builder, while the searchers 0x0087…, 0x2800…, and 0x51c7… successively affect those won by Rsync Builder.
-These pivotal providers are all integrated searchers of the corresponding builders that have been identified in previous works (Heimbach et al., 2024; Gupta et al., 2023).
-This observation corroborates findings from previous studies (Heimbach et al., 2024; Gupta et al., 2023; Thiery, 2023), confirming that the order flows from integrated searchers are crucial to those builders.
-Figure 5. Weekly builder-specific pivotal level for top-3 builders.
-4.2. Quantifying the Access Barrier
-Pivotal providers have different degrees of openness to builders: searchers are generally not accessible, and platforms like MEV Share and MEV Blocker are more open, although they also set up certain barriers.
-In this section, we present more details about the access barrier set up by providers.
+A builder who lacks access to any of their order flows will lose the majority of auctions.
+However, we found that order flows from these providers are inaccessible to all builders, which requires further investigation.
+Other observations from Fig.
+8 are provided in Appx.
+F.2.
+5.2 Trust Barrier
+A direct reason for the access barrier is that private order flow providers are at a disadvantage when interacting with builders and must trust builders not to unbundle the transactions.
+However, in practice, builders occasionally unbundle searcher’s bundles.
+For example, the f1b builder unbundled a bundle from the searcher jaredfromsubway.eth at block height 18476515.
+Although the searcher was compensated afterward, this incident highlights the fragile security foundation of the MEV-Boost auction.
+Due to the lack of trustworthy fair exchange mechanisms, providers have to rely on an informal reputation system for self-protection, sending order flows only to builders who meet their requirements.
+To understand the access barrier caused by the trust crisis, we investigate different providers and their requirements for accessing order flows.
 Searchers.
 Since searchers are operated by anonymous entities without any public communication channels, there is no public statement about which builders may access their flows.
-Existing studies indicate that searchers tend to send their order flow to the top builders (Titan, 2023) or share it exclusively with vertically integrated builders (Titan, 2023; Thiery, 2023; Gupta et al., 2023).
+Existing studies indicate that searchers tend to send their order flow to the top builders [81] or share it exclusively with vertically integrated builders [81, 80, 40].
 Telegram trading bots.
-The requirements for Telegram trading bots, in terms of sending order flow to builders, depend on whether they utilize channels like MEV Blocker (Maestro, 2023b; Labs, 2023).
+The requirements for Telegram trading bots, in terms of sending order flow to builders, depend on whether they utilize channels like MEV Blocker [56, 48].
 If so, their requirements are aligned with those of these channels.
-Banana Gun does not utilize these channels, nor does it explicitly state the requirements for builders.
+Banana Gun does not explicitly state the requirements for builders but we observe its preference for Titan, which we will discuss later.
 Channels.
@@
 The requirements of each provider are summarized in Tab.
-6.
+5.
 The channels impose various requirements on a builder’s reputation.
@@
 Intuitively, a builder with a high market share, seen as reputable, is less likely to compromise private order flows as doing so will forgo their future revenue.
-MeowRPC (Meow RPC, 2024) sends the order flow to trusted builders with high market share.
-Similarly, MEV-Share (Flashbots, 2024) requires that a builder possesses a competitive market share.
-MEV Blocker (MEV Blocker, 2023) requires that a builder must maintain at least 1% market share over one week.
-The cost of establishing minimal market share.
+MeowRPC [58] sends the order flow to trusted builders with high market share.
+Similarly, MEV-Share [37] requires that a builder possesses a competitive market share.
+MEV Blocker [60] requires that a builder must maintain at least 1% market share over each week of one month.
+Subsidization.
 New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share.
-In practice, builders pay out of pocket to overbid in MEV Boost auctions, a practice known as block subsidization.
-This represents a capital entry barrier for new builders, which we are able to quantify using the auction dataset.
-We compute the minimal subsidy needed to establish a 1% market share for one week (which is the requirement of MEV Blocker (MEV Blocker, 2023)) by replaying historical auctions with an added new builder who can only access the public order flow.
-We relegate the details to Sec.
-C.4.
-Figure 6. Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
-The result is plotted in Fig.
-6.
-Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 1.4 ETH by March 2024.
-It is important to recognize that the costs derived from this method only represent the lower bound.
-Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory.
-Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share.
-In reality, the subsidies new builders need to provide are substantially higher.
-For example, tbuilder (Tbuilder, 2024), a new builder in the builder market, subsidized 6.8 ETH in total for 645 blocks in the week before March 21, 2024 (Flashbots, 2024).
-Finding III: The private order flow providers impose reputation requirements on the builders, which are usually evaluated based on market share.
-The minimum entry cost for new builders required to establish a required market share increases over time.
-4.3. Trust Crisis and Centralization
-Our findings highlight a trust issue in today’s MEV supply chain: the lack of trustworthy fair exchange mechanisms between parties creates order flow silos and harms decentralization.
-Private order flow providers are at a disadvantage when interacting with builders and must trust builders not to unbundle the transactions.
-As a result, providers rely on an informal reputation system for self-protection, but it is not hard to see that reputation mechanisms are easily bypassed, as blockchain is fundamentally anonymous.
-Once a reputable builder receives the bundle from the searcher, the builder can create a Sybil and cheat the searcher (e.g., by unbundling the transaction bundle) under the new name.
-While such an attack has not been reported in practice, it is noteworthy that builders occasionally unbundle searcher’s bundles.
-For example, the f1b builder unbundled a bundle from the searcher jaredfromsubway.eth at block height 18476515.
-Although the searcher was compensated afterward, this incident underscores the fragile security foundation of the trust chain in the MEV-Boost auction.
-We also noticed that one anonymous builder666public key: 0xad5ce1019364de39…35af575554a7aa6f might be the Sybil of a reputable builder because this builder’s blocks included order flows from sources that are only accessible by reputable builders.
-The malicious actors within the trust chain pose a dual threat.
-On the one hand, security concerns deter searchers from submitting their bundles to multiple builders, further entrenching private order flow as a barrier to entry for new builders.
-On the other hand, integrating the roles of searcher and builder offers a potential mitigation to the security issue.
-Integrated searcher-builder has other structural advantages (Gupta et al., 2023; Pai and Resnick, 2023) (such as low latency) that may be unfair to regular builders.
-In Sec. 7.2, we discuss directions to bridge the trust gap between searchers and builders.
-Finding IV: The requirements for builders stem from the barrier of trust, as illustrated by the hidden trust crisis confirmed through real-world examples.
-Lowering or removing trust requirements in the MEV-Boost auction can facilitate decentralization.
-5. Competition and Efficiency in the MEV-Boost auctions
-Sec.
-4 focuses on the evolution of builder market.
-In this section, we focus on the status quo and study the implications of a centralized builder market as we have today.
-First, we investigate the (in)equality of builders’ block-building capability, measured by the true values of builders’ bids in MEV-Boost auctions.
-Then, we quantify the competitiveness and efficiency of historical MEV Boost auctions.
-These two studies complement each other.
-The former evaluates whether today’s builder market has the potential to be competitive, while the latter evaluates whether the current mechanism in fact induces the desired outcomes.
-Our study includes some MEV-Boost auctions, the highest bids of which are not included on Ethereum due to network latency.
-However, we do not omit them since the auctions themselves are still meaningful.
-5.1. Inequality in Block-building Capacity
-A primary metric to evaluate a builder’s block-building capacity is the MEV it extracts over time. In MEV Boost auctions, this quantity is captured by the true values of the builder’s bids.
-In the context of auctions generally, the true value of the item being auctioned refers to its actual worth to the bidder.
-Recall that in MEV-Boost auctions (Sec.
-2), a builder submits a block $B$ (a list of transactions) and a bid value $BV$ that the builder is willing to pay to the proposer if $B$ is proposed.
-The true value of $B$, denoted $TV(B)$, is typically higher than $BV$.
-One can think of $TV$ as the builder’s revenue and $BV$ as the cost.
-The difference $EV:=TV-BV$ is the profit pocketed by the builder, which we call extracted value (EV) — i.e., $EV(B)$ is the difference of the builder’s ETH balance after executing transactions in $B$.
-For example, suppose executing $B$ yields a net profit of $0.1$ ETH after paying $0.9$ ETH in bid value in the auction, then the true value (the revenue) is $1.0$ ETH.
-To compute $TV(B)$ of a given block $B$, we follow the standard approach in open source MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023), though we built our own analysis toolchain and optimized its performance to handle the large volume of auction data.
-We compute $EV(B)$ using a modified Rust Ethereum execution client (Reth (paradigmxyz, 2024)) to execute transactions in $B$ and compute the net ETH balance changes of the bidder’s addresses.
-Most builders’ addresses are well-known partly because they need to build up a reputation.
-We cross-validated it with other sources to ensure that our dataset was the most complete one to date (see Sec.
-3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
-In every slot $s$, bidders can submit multiple bids throughout the duration of the slot.
-We use $BV(p,s)$ to denote the highest bid value from bidder $p$ in slot $s$, and $TV(p,s)$ the corresponding true value.
-Quantifying the inequality of block-building capacities.
-We compare the highest true values of different builders in the same slot to understand the differences in builders’ ability to extract values.
-If a builder consistently has the highest true value, then she can potentially dominate the builder market.
-On the contrary, if many builders have similar true values, then it is plausible that a properly designed market could avoid a monopoly.
-To quantify the disparity of block-building capacity, we compute the quartile coefficient of dispersion (QCD) (Wikipedia, 2022) of the highest true values in a given slot.
-Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$.
-Let $Q_{1}(\vec{x})$ and $Q_{3}(\vec{x})$ representing the first and third quartiles of $\vec{x}$ (i.e., 25% and 75% percentiles).
-The QCD of slot $s$ is
-$QCD_{s}=QCD(\vec{tv}_{s})=\frac{Q_{3}(\vec{tv}_{s})-Q_{1}(\vec{tv}_{s})}{Q_{3}(\vec{tv}_{s})+Q_{1}(\vec{tv}_{s})}.$
-QCD is a real value in $[0,1]$.
-A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$).
-Since it uses quartiles, QCD has the benefit of being robust to outliers (Wikipedia, 2022).
-We now present the result given by QCD, noting that other metrics lead to a similar conclusion (see Sec. D.1).
-Evaluation results.
-We categorize the builders into three groups based on their market share from April to August 2023: top builders (top 5), middle builders (6-15), and tail builders (16-25).
-To understand how the magnitude of MEV in a slot affects the block-building capability, we categorize every slot into one of three tiers according to the final winning bid (as a proxy to the significance of the MEV opportunity in that slot): low MEV (below 0.0225 ETH, the 10th percentile of the winning bid values), high MEV (above 0.4586 ETH, 95th percentile) and medium MEV in between.
-In Fig.
-7, we plot the distribution of the QCD for MEV-Boost auctions from April 2023 to August 2023 for different builder groups and MEV tiers.
-(We sample the first week of each month.
-See Sec.
-3.1 for details.) It reveals two trends.
-In each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail.
-That is, top builders have comparable block-building capabilities, but middle and tail builders have significantly different capabilities.
-This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of builder market.
-Second, the inequality worsens as the MEV tier increases.
-One possible reason is that private order flow providers may tend to give significant MEV opportunities to select builders due to the trust barrier.
-Figure 7.
-Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders.
-The QCD measures the inequality of block-build capacity across builders.
-This figure reveals that such inequality worsens as MEV increases.
-Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
-Finding V: The inequality in block building is lowest among top builders, higher among middle builders, and most pronounced among tail builders, and the inequality worsens as a slot’s MEV increases.
-5.2. Quantifying Competitiveness and Efficiency
-The previous section shows that top-5 builders have comparable true values in most cases.
-Ideally, revenue from MEV-Boost auctions should be near optimal due to competition.
-However, whether real-world auctions induce such near-perfect competition is complicated by several confounding factors: builders may not be able to respond to bids due to network latency, they may collude implicitly or explicitly, or they may have incentives to overbid (e.g., to gain market share).
-In this section, we quantify the competitiveness and efficiency of the MEV-Boost auctions using what we call competitive index (CI) and efficient index (EI).
-Competitive index.
-In a given slot $s$, suppose builders $p_{1},\cdots,p_{n}$ are ordered by their true value from high to low, i.e., $TV(p_{1},s)\geq\dots\geq TV(p_{n},s)$, $BV_{w}(s)$ is the bid value of the winner and $TV_{w}(s)$ is the corresponding true value, the competitive index of slot $s$, $CI(s)$ is defined as:
-$CI(s)=\frac{BV_{w}(s)-TV(p_{2},s)}{TV(p_{2},s)}\times 100\%.$
-Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value.
-$CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, satisfying the definition of a competitive auction.
-Conversely, $CI(s)<0$ means that the competition is insufficient.
-Efficient index.   Same as above, the efficient index, $EI(s)$, is defined as:
-$EI(s)=\frac{TV_{w}(s)-TV(p_{2},s)}{TV(p_{2},s)}\times 100\%.$
-$EI(s)>0$ indicates that the winner has the highest true value, whereas $EI(s)\leq 0$ implies that the bidder with the highest true value lost the auction.
-Figure 8.
-The distribution of CI and EI, where each dot represents an auction.
-The color represents density when multiple dots overlap.
-88.84% of the MEV-Boost auctions are competitive (sum of the slots having $CI\geq 0$), and 79.74% of them are efficient (sum of the slots with $EI>0$).
-Dots on the diagonal line ($CI=EI$) are auctions where the winners bid the true values.
-The dots above this line are auctions where the winners bid lower than true values.
+In practice, builders pay out of pocket to overbid in MEV-Boost auctions, a practice known as block subsidization [36], representing an entry barrier for new builders.
+We quantify the minimal cost for a new builder to meet the MEV Blocker’s requirements (Appx.
+F.5), which reflects the costs of building a reputation.
+The minimal cost continues to increase and can be up to 7.4 ETH.
+This cost represents a lower bound since we do not have the corresponding numbers for other providers.
+Note that although subsidization can help meet the requirements set by channels, it cannot ensure that searchers’ trust concerns are addressed, allowing builders to receive their order flows.
+5.3 Incentive for Integration
+Another access barrier to private order flows is that providers exclusively share order flows with specific builders, known as integration [40].
+We compute each provider’s builder-specific pivotal level to understand the relationship between providers and builders.
+For a given provider $P$, its pivotal level specific to builder $B$ is the pivotal level of the MEV-Boost auctions where $B$ is the winner.
+Figure 9: Weekly builder-specific pivotal level for top-3 builders. MEV-Blocker and MEV-Share are hidden for clarity.
 Fig.
-8 shows the distribution of the calculated CI and EI across all MEV-Boost auctions in the ultra sound relay dataset.
-Notably, 74.83% of the auctions are both competitive and efficient: the winner possesses the highest true value and bids higher than or equal to the second-highest true value.
-88.84% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 79.74% (sum of the areas with $EI>0$) are efficient.
-5.2.1. Implications of uncompetitive auctions
-We first note Fig.
-8 shows a somewhat strict definition of competitiveness (i.e., $CI\geq 0$).
-In reality, builders may bid slightly lower than their true value to ensure profits.
-Therefore, it is reasonable to relax and consider an auction with $CI\geq-\delta$ to be competitive for a small $\delta$.
-We call this notion $\delta$-competitive.
-As shown in Fig.
-14, when we consider $\delta=1\%$ or $\delta=2\%$, the percentage of uncompetitive auctions reduces to about 6% and 5%, respectively.
-That is, under the slightly relaxed definition, approximately 95% auctions were competitive.
-Quantifying proposers’ losses.
-The concern with uncompetitive auctions lies in the potential loss to validators.
-If such losses are significant, top validators might be motivated to operate their own builders.
-Furthermore, vertical integration between top proposers and builders could adversely impact the decentralization of the validators (Bahrani et al., 2024).
-Thus, quantifying the exact loss validators incur due to insufficient competition is crucial.
-For each uncompetitive auction, we consider the second-highest true value as the ideal winning bid value and calculate the difference between this ideal value and the actual winning bid value.
-This difference represents the loss a validator incurs, which they would have received had the auction been competitive.
-Table 3. Validators’ losses due to insufficient competition
-Time #Slots #Slots w/ CI ¡ 0 (%) Profits (ETH) Losses (ETH) (%)
-April 9-15 28,376 1,437 (5.06%) 2,711.32 47.17 (2.02%)
-May 1-7 30,279 8,449 (27.90%) 9,367.50 118.50 (1.36%)
-June 1-7 35,414 3,137 (8.86%) 4,346.51 24.72 (0.62%)
-July 1-7 36,032 2,051 (5.69%) 3,945.67 18.41 (0.48%)
-August 1-7 17,825 1,424 (7.99%) 2,171.07 12.29 (0.58%)
-As shown in Tab.
-3, we note that the total losses caused by uncompetitive auctions amount to 221.09 ETH in our study, which represents 0.98% of the total gains of proposers, with an average loss of 0.0134 ETH per uncompetitive auction.
-The losses are particularly significant during May 1-7, 2023, when a total of 118.50 ETH was lost due to insufficient competition, which is 1.27% of their total gain.
-These losses may closely correlate with MEV opportunities, given that proposers’ total profits were also notably high during this period.
-Consequently, even though the losses seem to decrease, this reduction should not be interpreted as an indicator of increasingly intense competition.
-The concern regarding potential centralization persists, given that the losses experienced by validators have not been fully mitigated.
-Finding VI: About 11% of the MEV-Boost auctions in our study are uncompetitive, resulting in total losses of 221.09 ETH for the proposers. These losses represent 0.98% of the total gains of proposers.
-5.2.2. Reasons of inefficiency
-20.26% of MEV-Boost auctions were not efficient.
-Further investigation is required to understand why.
-In an inefficient auction, the builder with the highest true value does not win the auction.
-There are two possible reasons.
-First, block subsidization allows a builder to win the auction without possessing the highest true value.
-Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
-We analyzed the underlying reasons for all inefficient auctions.
-In 51.5% of cases, winners win the auctions due to subsidization.
-In 52.6% of cases, the builder with the highest true value shaded too much.
-It is important to note that auctions may exhibit inefficiency for multiple reasons.
-For instance, the winner may subsidize her bid while simultaneously, the builder with the highest true value shades.
-Thus, the cumulative percentage exceeds 100%.
-It is important to note that inefficiency does not necessarily disadvantage the validators. In fact, in inefficient scenarios involving subsidization, validators may receive more value.
-Finding VII: MEV-Boost auctions are generally competitive, yet their efficiency falls short of expectations.
-Further analysis reveals that 51.5% of inefficiencies are caused by subsidization, which, interestingly, does not disadvantage proposers but leads to increased earnings for them in such scenarios.
-5.2.3. MEV’s implication on auction outcomes
-To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we follow the same methodology and categorize the auctions into three tiers: low MEV, medium MEV, and high MEV.
-The ratio of competitive and efficient auctions in each tier is shown in Fig.
-9.
-We note that competitiveness decreases as the MEV of a slot increases, which is consistent with the finding in Sec. 5.1 about the inequality of block-building capabilities in high-MEV slots.
-The trend of efficiency is less expected.
-First, only 71.7% of the auctions within the low MEV tier are efficient.
-This can be attributed to the generally low overall bid values, rendering the auction outcomes vulnerable to alterations from minor subsidies.
-Second, while auctions tend to be more efficient in the medium MEV tier, we also observe that efficiencies worsen in the high MEV tier: only 77.2% of these auctions are efficient, and about 90% of inefficient cases result from bid shading, where the builders with the highest true values place their bids much lower than their true values.
-Figure 9. The ratio of auctions that are competitive and efficient across various MEV tiers.
-Finding VIII: The competitiveness of MEV-Boost auctions deteriorates as a slot’s MEV increases.
-Efficiency may decline due to block subsidization in the low MEV tier and bid shading in the high MEV tier.
-6. Related Work
+9 plots selected providers’ builder-specific pivotal levels for the top three builders, ranked based on their weekly market share in July 2024.
+We include providers whose weekly pivotal levels exceed 25% for any of the three builders over two months to illustrate these builder-specific pivotal levels.
+MEV Blocker and MEV-Share are excluded from this analysis, as all builders connected to them reportedly receive the same order flows.
+We note that some providers have a pivotal level exceeding 50% only with specific builders, suggesting a possible collaborative relationship.
+The searcher 0xA69b… consistently influences over 50% of the MEV-Boost auctions won by Beaver, while the searchers 0x0087…, 0x2800…, and 0x51c7… successively affect those won by Rsync.
+These pivotal providers are all integrated searchers of the corresponding builders that have been identified in previous works [42, 40].
+This observation corroborates findings from previous studies [42, 40, 80], confirming that the order flows from integrated searchers are crucial to those builders.
+Furthermore, we observe a clear variation in the importance of providers for different builders.
+For example, the significance of Banana Gun to Titan is greater than it is to other builders.
+We further investigate the relationship between Banana Gun and Titan and suppose that Banana Gun only sends valuable order flow to Titan when MEV is significant.
+A case in this point is slot 8614161, where only Titan received the valuable order flow from Banana Gun.
+The much lower bids from other builders suggest they didn’t receive this order flow; otherwise, they could have outbid Titan.
+We illustrate this new integration pattern in Appx.
+F.3.
+The results in Fig.
+15 indicate that this integration pattern also applies to jaredfromsubway.eth and Beaver.
+Both integrations emerged around January 2024.
+There is a clear incentive for integration between providers and builders.
+If MEV from a provider can determine the auction outcome, sending order flow to an integrated builder is better than sending it to multiple builders.
+The integrated builder only needs to bid slightly higher to win, and they can share the surplus MEV.
+However, this integration is detrimental to proposers, as it exacerbates the inequality in builders’ block-building capacities and can eventually lead to significant losses for them.
+Finding IV: Two providers (Banana Gun and jaredfromsubway.eth) exclusively share high-value order flows with specific builders (Titan and Beaver) respectively, suggesting a new integration pattern.
+6 Related Work
 Empirical study on Ethereum’s builder market.
@@
 An early exploration by Yang et al.
-(Yang et al., 2022) analyzed the market share of relays and builders.
+[92] analyzed the market share of relays and builders.
 Their findings indicated a centralization of Ethereum’s builder market, with Flashbots and builder0x69 accounting for over 53% of the market share between September 15th, 2022, and November 30th, 2022.
-Similar studies on the builder market conducted by (Wahrstätter et al., 2023b; Heimbach et al., 2023) also confirmed the centralization in the builder market.
+Similar studies on the builder market conducted by [84, 41] also confirmed the centralization in the builder market.
 Their observations were based on the market share and the distribution of total profits generated within the PBS ecosystem.
-Moreover, several online dashboards (Wahrstätter, 2024; Labs, 2024; Bitfly, 2024; Flashbots, 2024) offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics.
+Moreover, several online dashboards [87, 49, 7, 36] offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics.
 However, these studies do not quantify the causes of centralization or the security implications of a centralized builder market.
@@
 Many studies also discuss the importance of private order flows, particularly focusing on private order flows from integrated searchers.
-Thiery (Thiery, 2023) counted the total number of transactions, along with their values, from June 1 to July 15, 2023.
+Thiery [80] counted the total number of transactions, along with their values, from June 1 to July 15, 2023.
 He found that private order flows account for only 30% of the transactions yet represent 80% of the total value paid to builders.
@@
 Gupta et al.
-(Gupta et al., 2023) confirmed that builders are more likely to win the MEV auction when their integrated searchers can provide high-value exclusive order flow by effectively exploiting arbitrage between a centralized exchange (CEX) and a decentralized exchange (DEX).
-The dashboard by Winnie (Xiao, 2024) and the study by Titan (Titan, 2023) analyzed the vertical integration between searchers and builders through the quantity and value of transactions sent from searchers to builders.
+[40] confirmed that builders are more likely to win the MEV auction when their integrated searchers can provide high-value exclusive order flow by effectively exploiting arbitrage between a centralized exchange (CEX) and a decentralized exchange (DEX).
+The dashboard by Winnie [91] and the study by Titan [81] analyzed the vertical integration between searchers and builders through the quantity and value of transactions sent from searchers to builders.
 Heimbach et al.
-(Heimbach et al., 2024) observed that during their study, Beaver Builder received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH.
-This large gap suggests that Beaver Builder might have received significant off-chain profits from its integrated searchers.
+[42] observed that during their study, Beaver Builder received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH.
+This large gap suggests that Beaver might have received significant off-chain profits from its integrated searchers.
 These studies show that three integrated builders (Beaver, Rsync, and Manta) gain a significant advantage in order flow from their integrated searchers.
@@
 The strategies used by different entities in the MEV-Boost auctions and their potential impacts have attracted significant interest from researchers.
-The strategic behaviors of proposers were first studied by (Schwarz-Schilling et al., 2023; Wahrstätter et al., 2023b) focusing on the timing game where proposers strategically delay block proposals to maximize their profits.
+The strategic behaviors of proposers were first studied by [76, 84] focusing on the timing game where proposers strategically delay block proposals to maximize their profits.
 Öz et al.
-(Öz et al., 2023) further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it.
-A recent study (Always, 2024) empirically investigated timing games in the wild and how block proposal delays affect the block fork rate.
+[66] further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it.
+A recent study [1] empirically investigated timing games in the wild and how block proposal delays affect the block fork rate.
 The results indicate that three validator sets–P2P, Kiln, and Attestant, that are known or suspected of playing timing games–did not delay their block proposals enough to cause their blocks to be forked.
 Wu et al.
-(Wu et al., 2024) proposed a game-theoretic model for MEV-Boost auctions, using simulations to examine various builders’ bidding strategies, including naive, adaptive, last-minute, and bluff bidding.
+[90] proposed a game-theoretic model for MEV-Boost auctions, using simulations to examine various builders’ bidding strategies, including naive, adaptive, last-minute, and bluff bidding.
 Their results highlighted the importance of latency and exclusive order flow in the effectiveness of bidding strategies.
 Pai et al.
-(Pai and Resnick, 2023) analyzed the latency advantage of searcher-builder integration in adjusting bidding values.
+[67] analyzed the latency advantage of searcher-builder integration in adjusting bidding values.
 The strategies discussed in these papers highlight the potential factors that can make MEV-Boost auctions uncompetitive or inefficient, while our paper quantifies these uncompetitive and inefficient auctions.
@@
 A provider may exclusively share its order flow with a single builder.
-Previous works such as (Kilbourn, 2022; Frontier, 2023; Gupta et al., 2023) pointed out that exclusive order flows have potential centralizing effects on the builder market, because builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers and in winning order flow auctions.
+Previous works such as [46, 38, 40] pointed out that exclusive order flows have potential centralizing effects on the builder market, because builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers and in winning order flow auctions.
 The decentralization of Ethereum’s builder market not only benefits itself but also positively impacts the decentralization of validators.
-A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization (Bahrani et al., 2024).
+A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization [4].
 We complement these theoretical analyses with empirical studies to analyze potential problems caused by the current centralized builder market.
 Blockchain decentralization metrics.
-Nakamoto coefficient was proposed to describe the minimum number of entities required to disrupt the blockchain’s subsystem (Srinivasan and Lee, 2017).
-Other metrics, such as the Gini coefficient (Dorfman, 1979), Shannon entropy (Lin, 1991), and Herfindahl–Hirschman index (Rhoades, 1993), have also been utilized to measure decentralization (or the lack of it) in staking distribution (Grandjean et al., 2023; Kwon et al., 2019), mining power (Lin et al., 2021; Wu et al., 2019; Kwon et al., 2019), and block building (Heimbach et al., 2023).
-An SoK paper (Zhang et al., 2022) comprehensively summarized the existing works on blockchain centralization; however, none explored decentralization from the perspective of entry barriers or auctions.
-7. Discussion
-7.1. Limitations of Our Approach and Data
+Nakamoto coefficient was proposed to describe the minimum number of entities required to disrupt the blockchain’s subsystem [77].
+Other metrics, such as the Gini coefficient [18], Shannon entropy [51], and Herfindahl–Hirschman index [73], have also been utilized to measure decentralization (or the lack of it) in staking distribution [39, 47], mining power [52, 89, 47], and block building [41].
+An SoK paper [93] comprehensively summarized the existing works on blockchain centralization; however, none explored decentralization from the perspective of entry barriers or auctions.
+7 Discussion
+7.1 Mitigation
+Our results suggest that proposers suffer from significant losses because of builder centralization, which can lead to two undesirable outcomes: First, PBS could become unstable.
+Second, MEV measured by the auctions may be inaccurate.
+In this section, we propose a structural change to the existing MEV supply chain Sec. 2.2 and a solution based on the new supply chain structure to address these two problems.
+In the new supply chain, only one builder receives order flows from providers, and proposers can only accept blocks from this builder.
+Protocol.   Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15] The pseudocode for the TEE builderis described in Fig. 17.
+The builder provides two functions.
+First, it opens RPC endpoints to receive the private order flows from the providers.
+Providers send their order flows along with a bid indicating the MEV in their order flows.
+Second, when it receives a signal for block building (e.g., a new slot starts in Ethereum), it will run an ordering algorithm to build a block with optimal MEV based on the received order flows (a reasonably performant ordering algorithm is described in Appx.
+E).
+The TEE builder runs the truthful mechanism in [57]: it refunds all providers for their marginal value contribution to the block.
+For each included order flow, the builder builds an alternative block without that specific order flow to calculate its marginal value contribution by comparing the MEV extracted between these two blocks.
+Next, we outline the key advantages of this solution.
+Security guarantees.
+TEEs offer a secure and isolated environment, ensuring that private order flows remain confidential and protected from external tampering.
+Moreover, order flow providers can utilize remote attestation to verify that the builder running within the TEE is indeed the intended one, thereby effectively mitigating trust concerns.
+Incentive compatibility.
+[57] shows that for providers, truthfully bidding is a dominant strategy when they are refunded based on their marginal value contribution.
+The formal proof and the required assumptions are in [57].
+Here, we offer an intuitive explanation: a provider effectively pays the second price for their order flows when refunding their marginal value contribution.
+Analysis.
+Since truthfully bidding is an incentive-compatible dominant strategy for all providers, the TEE builder can compute the total MEV in a block from the providers’ bids, achieving an accurate MEV oracle.
+Moreover, because the builder program in the TEE can be programmed to pay the proposer fairly, the instability of PBS is also addressed.
+Future works.
+We leave the analysis of whether the new supply chain structure is sustainable in the long term for future work.
+Intuitively, since providers are incentivized to send order flows only to the builder due to refunds, other builders will naturally be driven out of the market, leaving proposers to only accept blocks from the TEE builder.
+Furthermore, since real-world proposers might behave rationally, exploring the new equilibrium that emerges among a rational provider, builder, and proposer is another intriguing question to be addressed.
+Finally, this solution provides an accurate MEV oracle and transparent MEV information.
+These insights open up new possibilities for future MEV mitigation directions.
+7.2 Limitations of Our Approach and Data
 True value.
-We compute the true values of bids using a similar approach as that in MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023).
+We compute the true values of bids using a similar approach as that in MEV analysis tools [20, 31] and previous works [41, 80].
 This approach has a few limitations.
@@
 Thus, we can’t evaluate its pivotal level.
-This limitation is common in existing studies (Lu, 2024; Lu, Angela and Sui, Danning and Durnford, Jaden, 2024).
+This limitation is common in existing studies [53, 54].
 However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec.
@@
 Therefore, underestimating the second-highest bid value might also lead us to underestimate the significance of each provider.
-7.2. Future Works
-Based on the findings, we now outline directions to enhance the decentralization of builder markets. These directions present interesting technical challenges that we leave for future work.
-Several issues we identified boil down to the lack of a trustworthy mechanism to facilitate fair exchanges between private order flow providers and builders.
-In today’s ecosystem builders are at an advantage.
-To protect themselves, private order flow providers resort to reputation-based solutions such as requiring a minimal market share, which creates siloes.
-One approach is building platforms to facilitate the said fair exchange.
-For example, MEV-Share is one such platform between searchers and builders.
-To receive order flows from MEV-Share, builders must agree to the Fair Market Principles set force by the community (Flashbots, 2023a).
-Should builders fail to adhere to these principles, they will be either removed from or flagged by the community.
-However, in its current form, MEV-Share is a trusted party.
-One direction being explored (e.g., by SUAVE (Flashbots, 2024c)) is to use Trusted Execution Environments (TEEs) to run block-building algorithms without leaking private order flows to the builders, but standard challenges such as side channels and covert channels (Salles et al., 2023) apply.
-Another unsolved problem is how to reliably detect private order flow leakage across builders.
-Currently, after receiving private orders, a builder can leak the orders to another builder (potentially a Sybil), bypassing any reputation or economic penalty.
-One idea is to embed watermarks in order flows to enable tracking.
-Watermarks must be invisible to the adversary and not interfere with transaction execution.
-Designing a secure watermarking scheme against a practical adversary model is an interesting future direction.
-A reliable leakage detection mechanism (e.g., watermarking) will open up the design space of reputation systems.
-For example, one possibility is to ask builders to deposit a certain amount of collateral to access private order flows, and order flow leakage will be penalized economically.
-8. Conclusion
-In this paper, we characterized the efficacy of PBS by evaluating its openness to new builders and assessing the competitiveness and efficiency of the auctions.
-Our study identifies a significant entry barrier in the builder market: private order flow.
-We further quantify the barrier costs—a new builder must pay up to 1.4 ETH to access this flow.
-Given this barrier, it is not surprising that we observed varying capabilities among builders in extracting MEV, with inequalities among top, middle, and tail builders worsening as MEV increases.
-Our study shows that 88% of the MEV-Boost auctions were competitive, while 79% were efficient.
-The proposers’ losses due to uncompetitive auctions are 0.98% of their total gain.
-Based on these insights, we propose suggestions for future developments to mitigate the trust crisis in the MEV-Boost auctions and facilitate the decentralization of Ethereum’s builder market.
+7.3 Security Problems in PBS
+As we mentioned in Sec.
+5.2, one reason for the access barrier to private order flows is the difficulty of establishing trust.
+A malicious builder, relay, or proposer may attack providers.
+A real-world incident involved a malicious proposer who double-signed a header to unbundle a builder block and further stole MEV from searchers by proposing a different block [61].
+Another security concern is censorship.
+Ethereum’s censorship resistance currently depends heavily on whether builders choose to censor, particularly the actions of the top builders.
+For example, before the top builder, Beaver, stopped censoring in July 2024, over 50% of blocks were censoring [86].
+Our study reveals that builder centralization remains difficult to overcome, potentially exacerbating the censorship problem in Ethereum.
+8 Conclusions
+In this paper, we interrogate the causes and implications of builder centralization in Ethereum and challenge the community’s belief that it is acceptable.
+Our empirical study shows that the current blocker market has led to significant profit losses for validators.
+These significant losses could lead to PBS’s instability and MEV oracles’ inaccuracy and, if left uncontrolled, can lead to validator centralization.
+We attribute the loss to two factors: insufficient competition and inequality of block-building capacities, with the latter being the dominant reason.
+We further investigated the trust and incentive issues of the current MEV supply chain that contribute to builder centralization.
+To address these two problems, we propose a structural change to the current MEV supply chain and a solution based on this new structure.
 Acknowledgements
@@
 We also thank the teams of BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions about access requirements.
-We are grateful for the valuable discussions with Thomas Thiery and Barnabé Monnot.
+We are grateful for the insightful comments by Prof.
+Campbell Harvey and the valuable discussions with Thomas Thiery, Barnabé Monnot, Danning Sui, Burak Öz, Data Always, and Georgios Konstantopoulos.
 Appendix A Visualization for Distribution
 Fig. 10 shows the distribution of the percentage of builders who submit to the ultra sound relay. Fig. 11 shows the CDF of the winning bid values of MEV-Boost auctions.
-Figure 10. Distribution of the percentage of builders who submit to the ultra sound relay.
-Figure 11. CDF of the winning bid value of auction dataset.
-Appendix B List of Searchers and Builders
+Figure 10: Distribution of the percentage of builders who submit to the ultra sound relay.
+Figure 11: CDF of the winning bid value of auction dataset.
+Appendix B Data Collection
+Builder public keys.
+We collect builder public keys from two data sources.
+First, builders often disclose their public keys in official documents.
+For example, Titan builder has disclosed 15 public keys in its official document [10].
+Second, some builders leave their unique marks in the blocks’ “extra data” field.
+For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’.
+Given that such marks are unauthenticated and susceptible to impersonation, we ignore them unless they are used more than 100 times in the bids by the same builder public key.
+Transaction sources.
+We identified transactions originated from Telegram trading bots, including Maestro [55], Banana Gun [5], and Unibot [82], by searching for transactions that interacted with their official smart contracts.
+Two channels, MEV Blocker and MEV-Share publish data [70, 33] about transactions submitted to them.
+Specifically, transactions submitted to MEV Blocker are publicly available on Dune [70], from which we extracted 22,819,437 transaction hashes that fall within our study’s range.
+We obtained 20,334,513 transactions submitted to MEV-Share from the public Flashbots dataset [33].
+These transaction labels are sourced from their official datasets and we view them as ground truth.
+Appendix C List of Searchers and Builders
 In Tab.
-4, we provide the list of searchers and their addresses mentioned in our study.
+3, we provide the list of searchers and their addresses mentioned in our study.
 In Tab.
-5, we have counted the top 25 builders in the market, ranked by their market share from April 2023 to August 2023.
-Table 4. The list of searchers
+4, we have counted the top 25 builders in the market, ranked by their market share from April 2023 to December 2023.
+Table 3: The list of searchers
 Searcher Address
@@
 Searcher: 0x51c7… 0x51C72848c68a965f66FA7a88855F9f7784502a7F
-Table 5. The percentages of blocks and MEV-Boost blocks built by each builder from April to August 2023.
+Table 4: The percentages of blocks and MEV-Boost blocks built by each builder from April to December 2023.
 Builder % of total blocks % of MEV-Boost blocks
-Beaver 22.418739 24.719523
-Rsync 18.611899 20.521996
-Builder0x69 17.799124 19.625807
-Flashbots 13.276622 14.639172
-Titan 7.105763 7.835011
-BloXroute 2.384068 2.628740
-Blocknative 2.267458 2.500162
-F1b 2.158849 2.380407
-eth-builder 1.086828 1.198366
-BuildAI 0.894899 0.986741
-payload.de 0.645310 0.711536
-lightspeedbuilder 0.321781 0.354805
-boba-builder 0.257682 0.284128
-Manta 0.254004 0.280072
-Eden 0.202596 0.223388
-0x3be…436 0.186227 0.205339
-It’s Free Real Estate 0.133072 0.146728
-Gambit Labs 0.117806 0.129896
-bobTheBuilder 0.102815 0.113367
-Manifold 0.088929 0.098056
-blockbeelder 0.084239 0.092884
-finest artisanal blocks 0.074307 0.081933
-nfactorial 0.072375 0.079803
-0x563…cf3 0.041660 0.045935
-uwu.builders 0.016370 0.018050
-Appendix C Openness of the builder market
-C.1. Importance of private order flows
+Beaver 22.53 24.69
+Rsync 20.75 22.74
+builder0x69 12.65 13.86
+Titan 12.35 13.53
+Flashbots 12.24 13.42
+BloXroute 2.14 2.34
+f1b.io 1.66 1.81
+Blocknative 1.40 1.53
+eth-builder 0.77 0.84
+BuildAI 0.73 0.80
+Gambit Labs 0.64 0.70
+Jetbldr 0.47 0.51
+Manta 0.46 0.51
+payload.de 0.39 0.42
+boba-builder.com 0.27 0.30
+Penguinbuild 0.23 0.25
+Eden 0.22 0.24
+lightspeedbuilder 0.21 0.23
+0x3be…436 0.16 0.18
+lokibuilder.xyz 0.12 0.13
+bobTheBuilder.xyz 0.11 0.13
+blockbeelder.com 0.11 0.12
+finest artisanal blocks 0.09 0.10
+It’s Free Real Estate 0.08 0.09
+nfactorial.xyz 0.06 0.07
+Appendix D Implications of Builder Centralization
+D.1 Efficiency of MEV-Boost Auctions
+Analysis of inefficiency.
+Fig.
+4 shows that 24.9% of MEV-Boost auctions were not efficient.
+In an inefficient auction, the builder with the highest true value does not win the auction.
+There are two possible reasons.
+First, a builder can win the auction without possessing the highest true value by overbidding, known as block subsidization [36].
+Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
+We analyzed the underlying reasons for all inefficient auctions.
+In 61.0% of cases, winners win the auctions due to subsidization.
+In 48.0% of cases, the builder with the highest true value shaded too much.
+It is important to note that auctions may exhibit inefficiency for multiple reasons.
+For instance, the winner may subsidize her bid while simultaneously, the builder with the highest true value shades.
+Thus, the cumulative percentage exceeds 100%.
+It is important to note that inefficiency does not necessarily disadvantage the validators. In fact, validators may receive more revenues in inefficient scenarios involving subsidization.
+MEV’s implication on efficiency.
+The ratio of efficient auctions in each tier of MEV is shown in Fig.
+5.
+We note that the trend of efficiency is less expected.
+First, only 56.6% of the auctions within the low MEV tier are efficient.
+This can be attributed to the generally low overall bid values, rendering the auction outcomes vulnerable to alterations from minor subsidies.
+Second, while auctions tend to be more efficient as MEV increases, we still observe that 77.1% of auctions in the high MEV tier are efficient, and about 86% of inefficient cases result from bid shading, where the builders with the highest true values place their bids much lower than their true values.
+D.2 Measuring Block-building Capacities
+Coefficient of variation.
+We use $\mathsf{CV}(\vec{v})\to\mathbb{R}_{\geq 0}$ to denote the coefficient of variation (CV), a real value in $\mathbb{R}_{\geq 0}$, for values $\vec{v}=(v_{1},\dots,v_{n})$.
+Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$.
+The CV of slot $s$ is computed as $CV(\vec{tv}_{s})$.
+Figure 12: Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of block-build capacity across builders.
+As shown in Fig.
+12, we can make the same finding as Fig.
+6.
+It reveals that such inequality worsens as MEV increases.
+Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
+D.3 Implications of Uncompetitive Auctions
+Fig. 13 shows the distribution of the competitive index (CI) of the MEV-Boost auctions in our study.
+Figure 13: The distribution of the CIs of MEV-Boost auctions
+Appendix E Reference Builder Algorithm
+Flashbots has open-sourced a Rust implementation of the MEV-Boost builder [34], which serves as a reasonably performant reference for builder algorithms.
+We will briefly introduce their approach.
+The builder processes each bundle (where a single transaction is considered a bundle containing one transaction) by computing its effective gas price and total profit.
+Subsequently, the builder sorts all received bundles in descending order based on either effective gas price or total profit.
+Employing a greedy algorithm, the builder then appends bundles from this sorted list.
+1 presents the pseudocode of their algorithm.
+Listing 1: Flashbots block-building algorithm
+# Compute the effective gas price and profit for a given bundle
+# A single transaction is recognized as a bundle of size 1
+def receive(bundle):
+effective_gas_price, profit = compute_price_and_profit(bundle)
+bundle.egp = effective_gas_price
+bundle.profit = profit
+return bundle
+# Building a block
+def build(bundles, sorting_key):
+# Sort bundles by price or profit
+if sorting_key == ’egp’:
+bundles.sort(key=lambda b: b.egp)
+else:
+bundles.sort(key=lambda b: b.profit)
+# Initialize
+B = []
+# Get the current state
+state = get_current_state()
+# Greedy algorithm
+while not bundles.is_empty():
+# Get the first bundle from the queue
+bundle = bundles.pop(0)
+# Simulate the execution of the bundle
+success = simulate_bundle(state, bundle)
+# Update with the successful bundle
+if success:
+B.append(bundle)
+state.update(bundle)
+# Compute payment based on the final state
+payment = compute_payment(state)
+# Append the payment to the block
+B.append(payment)
+# Return the valid block
+return B
+Appendix F What Caused Inequality in Block-building Capacities
+F.1 Importance of Private Order Flows
 To quantify the importance of private order flows for building valuable blocks, we compute the ratio of block values derived from private transactions.
@@
 3.1.
-Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total transaction priority fees (Ethereum, 2024) and direct transfers (Flashbots, 2024a) to the block builder) to the total values extracted from all transactions in the block.
-C.2. Case Study for Maestro
+Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total transaction priority fees [21] and direct transfers [32] to the block builder) to the total values extracted from all transactions in the block.
+F.2 Pivotal Private Order Flow Providers
+We can make a few observations from Fig.
+8.
+First, a particular searcher, jaredfromsubway.eth, has been gaining prominence since February 2023 and reached its peak significance in May 2023.
+During this period, the winners of more than 70% of the MEV-Boost auctions would have failed had they not received order flow from jaredfromsubway.eth.
+Its influence has gradually decreased since then but still remains pivotal in 20%-40% of auctions since Nov 2023.
+Besides, the order flow from Telegram trading bots has become vital for builders with the rise of such bots.
+For instance, Banana Gun is pivotal in about 40% of the MEV-Boost auctions in September 2023.
+Second, we can observe a positive trend from Fig.
+8 that MEV Share and MEV Blocker have been increasingly pivotal since May 2023.
+These two providers positively impact decentralization because they have relatively clear requirements for accessing their order flows (the requirements of different providers will be discussed in Sec.
+5.2).
+Third, we observe a notable downtrend in the pivotal level of MEV-Share and MEV Blocker starting from November 2023, accompanied by an uptrend in Maestro, a popular Telegram trading bot [55], beginning in December 2023.
+This is because Maestro used to be an important upstream of MEV-Share and MEV Blocker but stopped in November 2023, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
+Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned.
+For example, Mastro may prefer to send flows to builders directly for a better price than what MEV-Share and MEV Blocker can offer.
+It is surprising to see that upstream providers can significantly swing the significance of platforms.
+Finding V: We identified five pivotal providers (MEV-Share, MEV Blocker, jaredfromsubway.eth, Banana Gun, Maestro) who had a sustained influence on the outcome of more than half of MEV-Boost auctions.
+A positive trend towards decentralization is observed as the pivotal level of the searcher jaredfromsubway.eth decreased, while the pivotal level of MEV-Share and MEV Blocker increased.
+However, their reliance on upstream providers like Maestro is alarming.
+Case study for maestro.
 To delve deeper into the relationship between these two trends, we plot Maestro’s contribution to the total daily value of transactions sent through MEV-Share and MEV Blocker during our study in Fig.
-12.
+14.
 Surprisingly, during the time periods when MEV-Share and MEV Blocker were highly pivotal in Fig.
-4, i.e., from May to November, Maestro’s contribution accounts for approximately 80%!
+8, i.e., from May to November, Maestro’s contribution accounts for approximately 80%!
 This means Maestro used to be an important upstream of MEV-Share and MEV Blocker.
 In November 2023, Maestro stopped sending transactions to MEV-Share and MEV Blocker, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
-Figure 12. Daily percentage of the value of the transactions originating from Maestro for both MEV-Share and MEV Blocker.
-C.3. Requirements of Providers
+Figure 14: Daily percentage of the value of the transactions originating from Maestro for both MEV-Share and MEV Blocker.
+Figure 15: Weekly builder-specific pivotal levels for the top-3 builders for slots with winning bid values of no less than 0.5 ETH.
+F.3 New Integration Pattern
+To illustrate the new integration pattern between providers and builders, we consider the MEV significant in a slot if its winning bid value is no less than 0.5 ETH.
+Additionally, we focus only on the provider that is the sole pivotal provider in a slot, meaning this provider provides significant MEV to the builder, enabling the builder to win the auction.
+We apply the above adjustments to builder-specific pivotal levels, and the results are shown in Fig.
+15.
+Compared to Fig.
+9, we observe that Banana Gun and Titan, as well as jaredfromsubway.eth and Beaver, follow this integration pattern: a provider selectively shares valuable order flows with specific builders when MEV is significant.
+F.4 Requirements of Providers
 Tab. 5 summarizes the providers we identify and their requirements for accessing their private order flows.
@@
 Searcher Individual searchers Requirements are unclear.
-Telegram trading bot Banana Gun (Banana Gun, 2023), Maestro (Maestro, 2023a), Unibot (Unibot X, 2023) The requirements are the same as the channels they use or unclear.
-Channel BloXroute (bloXroute Labs, 2024) Builders are granted access to different tiers, each providing varying privileges and permissions, based on an assessment.
-Merkle (Merkle Software Inc., 2023) Builders must have a proven track record of accurately and honestly computing bundles, or they should consult with the Merkle team for consideration.
-MeowRPC (Meow RPC, 2024) The transactions are sent to the top builders which MeowRPC’s team trusts.
-MEV Blocker (MEV Blocker, 2023) Builders must maintain at least 1% market share over one week.
-MEV-Share (Flashbots, 2024) Builders must demonstrate a competitive market share for consideration.
-Transaction Boost (Blocknative, 2024) Consideration for partnership is based on the Blocknative team’s evaluation of the builder’s reputation.
+Telegram trading bot Banana Gun [5], Maestro [55], Unibot [82] The requirements are the same as the channels they use or unclear.
+Channel BloXroute [9] Builders are granted access to different tiers, each providing varying privileges and permissions, based on an assessment.
+Merkle [59] Builders must have a proven track record of accurately and honestly computing bundles, or they should consult with the Merkle team for consideration.
+MeowRPC [58] The transactions are sent to the top builders which MeowRPC’s team trusts.
+MEV Blocker [60] Builders must maintain at least 1% market share each week of the previous month.
+MEV-Share [37] Builders must demonstrate a competitive market share for consideration.
+Transaction Boost [8] Consideration for partnership is based on the Blocknative team’s evaluation of the builder’s reputation.
   • $\star$
     We do not claim that the list of the private order flow providers is complete since there is no ground truth.
-Table 6. Collections of the private order flow providers and their requirements.
-C.4. Minimal Cost for 1% Market Share
-We now present how we compute the minimal subsidy for a builder to account for 1% market share over one week.
+Table 5: Collections of the private order flow providers and their requirements.
+F.5 Minimal Cost for 1% Market Share
+We compute the minimal subsidy needed to establish a 1% market share for one month (which is the requirement of MEV Blocker [60]) by replaying historical auctions with an added new builder who can only access the public order flow.
 We assume that all builders have access to public order flows but the new builder cannot access any of the private order flows.
@@
 Or they will not increase subsidies if they already subsidize.
+For simplicity of computation, we assume that one month has 30 days.
 Under our assumptions, a new builder must submit a bid at least equal to the auction winner’s maximum–the higher of its bid value or true value.
@@
 Then, as we know the subsidy for proposing one block, we can further calculate the minimum overall subsidy a new builder would need to cover to propose 72 blocks each day over one week (Ethereum produces 7200 blocks a day.)
-Appendix D Competition and Efficiency in the MEV-Boost auctions
-D.1. Measuring block-building capabilities
-Coefficient of variation.
-We use $\mathsf{CV}(\vec{v})\to\mathbb{R}_{\geq 0}$ to denote the coefficient of variation (CV), a real value in $\mathbb{R}_{\geq 0}$, for values $\vec{v}=(v_{1},\dots,v_{n})$.
-Specifically, suppose the true values of different builders in slot $s$ are $\vec{tv}_{s}$.
-The CV of slot $s$ is computed as $CV(\vec{tv}_{s})$.
-Figure 13. Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of block-build capacity across builders.
-As shown in Fig.
-13, we can make the same finding as Fig.
-7.
-It reveals that such inequality worsens as MEV increases.
-Within a given MEV tier, inequality is greatest among tail builders, followed by middle builders, and is least among top builders.
-D.2. Implications of uncompetitive auctions
-Fig. 14 shows the distribution of the competitive index (CI) of the MEV-Boost auctions in our study.
-Figure 14. The distribution of the CIs of MEV-Boost auctions
+The result is plotted in Fig.
+16.
+Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 7.4 ETH by July 2024.It is important to recognize that the costs derived from this method only represent the lower bound.
+Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory.
+Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share.
+In reality, the subsidies new builders need to provide are substantially higher.
+For example, tbuilder [78], a new builder in the builder market, subsidized 6.8 ETH in total for 645 blocks in the week before March 21, 2024 [36].
+Figure 16: Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
+Appendix G Pseudocode of TEE Builder
+The pseudocode for the builder running in TEEs with a refund mechanism is shown in Fig. 17.
+[…]
+$orders\leftarrow[]$ // Pending transactions.
+$sources\leftarrow\{\}$ // Record the order flow providers.
+// Receive order flows from providers.
+[…]
+$orders\leftarrow orders+[order]$
+$sources[order]\leftarrow provider$
+// Build best block and refund providers.
+On building:
+$v,B\leftarrow\textsc{build}(orders)$
+$payments\leftarrow[]$
+// For each provider, refund its marginal contribution,
+ For every $order$ in $B$:
+  $provider\leftarrow sources[order]$
+  $v^{\prime},\_\leftarrow\textsc{build}(orders-[order])$
+  $payment\leftarrow\textsc{transfer}(provider,v-v^{\prime})$
+  $payments\leftarrow payments+[payment]$
+// Remove all transactions included on chain.
+ For every $order$ in $B$:
+  $orders\leftarrow orders-[order]$
+  Delete $sources[order]$
+$B\leftarrow B+payments$
+[…]
+Figure 17: The pseudocode of TEE builder.

### v4

@@
 In this paper, we empirically studied one of the least decentralized parts of Ethereum, its builder market.
-The builder market was introduced to fairly distribute Maximal Extractable Values (MEV) among validators and avoid validator centralization.
-As of the time of writing, three builders produced the vast majority (more than 80%) of blocks in Ethereum, creating a concerning centralization factor.
-However, the community believes that such centralization “is okay,” arguing that builder centralization will not lead to validator centralization.
-In this empirical study, we interrogate the causes and implications of builder centralization and challenge this belief that it is acceptable.
-Our main finding is that builder centralization has led to a significant loss by validators and, if left uncontrolled, could lead to validator centralization.
+The builder market was introduced to fairly distribute Maximal Extractable Value (MEV) among validators and avoid validator centralization.
+As of the time of writing, two builders produced more than 85% of blocks in Ethereum, creating a concerning centralization factor.
+However, a common belief is that such centralization “is okay,” arguing that builder centralization will not lead to validator centralization.
+In this empirical study, we quantify the significant proposer losses within the centralized builder market and challenge the belief that this is acceptable.
+The significant proposer losses, if left uncontrolled, could undermine the goal of PBS.
 Moreover, MEV mitigation solutions slated for adoption are affected too because they rely on the builder market as an “MEV oracle,” which is made inaccurate by centralization.
-Our investigation revealed two reasons behind builder centralization.
-We propose a structural change to the existing MEV supply chain and a solution based on the new supply chain structure.
-However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
+Our investigation reveals the incentive issue within the current MEV supply chain and its implications for builder centralization and proposer losses.
+Finally, we analyze why the proposed mitigation cannot work and highlight two properties essential for effective solutions.
 1 Introduction
@@
 It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) [7], profits that protocol participants, known as validators, can reap by manipulating the ordering of transactions.
-MEV can lead to centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have [4, 23, 67, 12, 11].
+MEV can lead to validator centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have [5, 8, 9, 10, 11].
 To distribute MEV fairly among validators and avoid validator centralization, Ethereum adopted a solution called Proposer-Builder Separation (PBS) [8].
@@
 Ideally, builders compete with each other, and the proposer collects a significant portion of MEV in auction revenue without doing much work, eliminating the advantage of sophisticated validators over regular ones.
-Figure 1: Market share of builders. The builder market is one of the least decentralized parts of Ethereum!
+Figure 1: Market share of builders (the source of the data is introduced in Sec. 3). The builder market is one of the least decentralized parts of Ethereum!
 Is the builder market working as intended?
@@
 As shown in Fig.
-1, three builders won over 90% of auctions, raising questions about the competitiveness of the market.
-However, the community believes that “it is OK to centralize block building” [23], arguing that block centralization will not lead to validator centralization.
+1, two builders won over 90% of auctions.
+The builder centralization creates immediate security problems like censorship [4].
+However, the community still believes that “it is OK to centralize block building” [8], arguing that block centralization will not lead to validator centralization.
 We challenge this belief through an empirical study of the past MEV-Boost auctions.
-As will be shown in detail, builder centralization has caused a significant loss to the proposers and, if left uncontrolled, could lead to undesired consequences including validator centralization.
-1.1 Implications of Builder Centralization
+As will be shown in detail, proposers have incurred significant losses in a centralized builder market and, if left uncontrolled, could lead to undesired consequences including undermining PBS.
+1.1 Proposer Loss in the MEV-Boost Auctions
 Proposer loss and its consequences.
 Intuitively, a proposer’s loss is the difference between the MEV available in a block and the auction revenue.
-Builder centralization can cause proposer loss in two ways.
+Proposers may incur losses in a centralized builder market in two ways.
 First, auctions may not be competitive enough due to, e.g., collusion or other limitations of the auction mechanism.
@@
 This is undesirable because it re-introduces the dynamics that PBS wants to avoid: sophisticated validators would have an advantage over small ones in extracting MEV.
-In other words, significant proposer losses would make PBS “unstable”.
-Moreover, several MEV mitigation solutions (e.g., MEV-burn [19]) rely on the MEV-Boost auctions as an “oracle” to know how much MEV has been extracted from the system (e.g., MEV-burn proposed to remove that amount from circulation, effectively redistributing them to all holders of ETH through deflation.) The proposer’s loss translates to inaccuracy in the MEV oracle.
-Next, we introduce the auction mechanism and present our method to quantify the proposer losses using auction data.
+Moreover, several MEV mitigation solutions rely on the MEV-Boost auctions as an “oracle” to know how much MEV has been extracted from the system (e.g., MEV-burn [13] proposed to remove that amount from circulation, effectively redistributing them to all holders of ETH through deflation.) The proposer’s loss translates to inaccuracy in the MEV oracle.
+Given the undesired consequences of proposer losses, it is important to understand the extent of losses incurred by proposers in the current builder market.
+Next, we introduce the auction mechanism and present how we quantify proposer losses using auction data.
 Modeling MEV-Boost auctions.
 MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction [14], with a 12-second deadline.
-For each auction instance, builders submit bids in the form of $(B,BV)$ specifying a block $B$ and a bid value (BV) the builder is willing to pay for $B$ to be added to the blockchain.
-The revenue a bidder gains from $B$ is called the true value ($TV(B)$), which can be estimated from transactions in $B$.
+For each auction instance, builders continuously submit bids in the form of $(B,BV)$ specifying a block $B$ and a bid value (BV) the builder is willing to pay for $B$ to be added to the blockchain.
+The revenue a bidder gains from $B$ is called the true value ($TV(B)$), which we estimate from transactions in $B$.
 At the end of an auction, the block $B$ with the highest bid is added to the blockchain, the builder pays $BV$ to the proposer and earns $TV$ when $B$ is executed (i.e., the builder’s profit is $TV-BV$).
-Recall that the proposer’s loss is the difference between the auction revenue and the MEV available in a block.
-The latter can be represented by the builders’ highest true values.
-With that, the two types of losses can be computed from auction data, as illustrated in Fig.
+The two types of losses defined above can be computed from auction bids, as illustrated in Fig.
 2.
@@
 The loss from inequality is $TV_{2}-TV_{1}$ because a competitive auction can, at most, generate revenue of $TV_{2}$.
-The loss from uncompetitiveness is $\max(TV_{2}-BV,0)$ (if this value is negative, the auction is considered competitive, and there is no loss from uncompetitiveness.)
+The loss from uncompetitiveness is $\max(TV_{2}-BV,0)$; if $TV_{2}-BV$ is negative, the auction is considered competitive without this loss.
 Figure 2: Definition of proposer losses.
@@
 This indicates that inequality is the primary driver of the proposer’s losses in the MEV-Boost auctions.
-1.2 What caused Inequality in Block-building Capacities
+1.2 Private Order Flows
 Given that proposers suffer significant losses due to the builders’ block-building capacity inequality, we now investigate the underlying causes of this inequality.
-In the current MEV ecosystem (to be presented in Sec.
-2.2), builders receive order flows from order flow providers and compute the optimal ordering of received order flows (see Appx.
-E for a reference builder algorithm).
+In the current MEV ecosystem (c.f.
+Sec.
+2.2), builders receive order flows from order flow providers and compute the optimal ordering of received order flows (see [17] for a reference builder algorithm).
 Thus, a builder’s block-building capacity is determined by the “quality” of the order flows it receives.
 At the same time, not all order flows are accessible to all builders.
-The difference in builder’s ability to access important private order flows drives the inequality in block-building capacity.
-This also explains builder centralization: only a small set of builders can access these profitable order flows.
-We proposed a metric to identify important order flows, and we investigated their accessibility by different builders and the underlying reasons.
+The difference in the builder’s ability to access important private order flows can drive the inequality in block-building capacity.
+Not every provider is equally important, and some providers who generate profitable order flows can significantly impact builders.
+Thus, we propose a metric to identify important providers and investigate their implications.
 Identifying pivotal private order flow providers.
@@
 The pivotal level of a provider $P$ is the percentage of MEV-Boost auctions in which $P$ is pivotal.
+Other metrics, calculated based on each provider’s contribution to the builders’ blocks [18, 19], may include temporarily influential providers, whereas this metric identifies sustainedly influential providers, resulting in fewer false positives.
 Using the auction dataset, we identified five pivotal providers whose pivotal level exceeds 50% over a period longer than two weeks.
 That is, if a builder cannot access the flows from any of these providers, it will lose the majority of the auctions during our study period.
-These order flows are not equally inaccessible to all builders for two reasons we now introduce.
-Inequality due to trust barriers.
-Order flow providers do not openly share private order flows with any builder because a malicious builder can harm the provider by, e.g., sandwich attacks [94] and imitation attacks [71].
-Providers set up different access barriers for self-protection.
-Most MEV searchers (a specific type of order flow providers) resort to “obscurity for security” in that they are not publicly reachable; thus, we could not learn about their access barriers and how to overcome them.
-On the other hand, MEV channels (another type of provider) typically publish a clear access policy, which allowed us to gain some insight into the nature of their access barrier.
-For example, MEV-Blocker only shares order flows with builders with a market share of at least 1%, using market share as a proxy for reputation and trustworthiness.
-This requirement introduces a chicken-and-egg problem for new builders: to gain market share, they need access to private order flows to win auctions, but as newcomers, they lack the required market share to access these flows.
-As a result, new builders often resort to subsidization: bidding higher than their actual profits in MEV-Boost auctions to establish the required market shares.
-Therefore, a new builder needs to keep subsidizing until reaching a certain market share to overcome the access barriers set by MEV channels.
-The cost to do so continues to increase over time and can reach up to 7.4 ETH, according to our quantification.
-In summary, with a moderate cost, builders can overcome the access barrier of one type of order flow provider (MEV channels).
-However, how to access order flows from other providers, particularly important searchers, remains unclear.
-Inequality due to integration.   A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders, causing an inequality in block-building capacities.
-To identify such provider-builder pairs, we introduced builder-specific pivotal level of a provider $P$ for a builder $B$ as the pivotal level of $P$ in auctions won by $B$.
-With this metric, we rediscovered known exclusive relationships (called “integration” [40]), but also found unreported integrated providers for top builders (Titan and Beaver respectively).
-This finding is interesting because Titan has been the only top builder without a well-known integrated provider.
-Our findings show that all three top builders, accounting for 90% of market share, have integrated providers.
+Order flows from these pivotal providers have two levels of accessibility, which we introduce below.
+Accessibility by multiple builders.
+Among the top 5 pivotal providers, MEV-Share and MEV Blocker send their order flows to all builders they consider reputable.
+They do not cause inequality loss since the competition among these builders allows proposers to capture this value.
+The reputability of a builder is accessed per their specific rules. While top builders usually pass the assessments, it is harder for new builders to enter the market.
+Exclusive pivotal providers.
+A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders.
+Besides the well-known exclusive relationships (known as “integration” [20]), we also identified that three of the top five pivotal providers, previously not known to have integration, began integrating with top builders in 2024.
+Our finding shows that the integration of these three pivotal providers accounted for at least 45% of the total proposer losses in 2024.
 Integration poses a market entry barrier as new builders likely lack integrated providers.
@@
 Integration is detrimental to proposers, as it will reduce their profits. The strong incentive for integration makes it increasingly difficult to address builder centralization.
-1.3 Mitigation Ideas
-To summarize, we found that trust barriers and incentives for integration contribute to builder centralization, leading proposers to incur significant losses.
-The consequences include the instability of PBS and the inaccuracy of MEV oracles.
-We propose a structural change to the existing MEV supply chain (Sec.
-2.2) and a solution based on the new supply chain structure.
-However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
-The new supply chain has only one builder to which the order flow providers send flows and from which the proposers accept blocks.
-In this setting, a recent work [57] suggests a truthful mechanism that incentivizes order flow providers to report the MEV in their order flows truthfully.
-Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15], hardware-based enclaves that guarantee the confidentiality and integrity of programs running inside.
-Order flow providers verify that the correct mechanism is running in the TEE using remote attestation, then send order flows to the TEE along with a bid indicating the MEV in their order flow.
-Then, the TEE computes a fair refund to order flow providers.
-The refund mechanism is designed so that bidding truthfully is a dominant strategy [57].
-(While oversimplifying, one can make a loose analogy to sealed-bid second-price auctions among order flow providers.) Then, the TEE can compute the total MEV in a block from the providers’ bids, solving the problem of inaccurate MEV oracles.
-This also addresses the instability of PBS because the builder program in the TEE can be programmed to pay the proposer fairly, leaving just enough for its operation.
-Since this builder is fair to order flow providers, it is conceivable that it will be preferred by order flow providers over other builders.
-However, as mentioned above, it is unclear if having a single builder is an equilibrium.
-Also, the proposers may be incentivized to become order flow providers, which may re-introduce the concern similar to PBS instability.
-We elaborate on the idea in Sec.
-7.1.
-Contribution
+1.3 Mitigation
+Our study reveals trust concerns and incentive issues within the MEV supply chain.
+They cannot be resolved by the proposed alternative PBS designs [21, 13, 22, 23], as these designs may retain the same underlying supply chain.
+To fully address these problems, we identify two essential properties: a security guarantee protecting providers from malicious builder attacks and a game-theoretical guarantee disincentivizing provider integration.
+While solving these problems remains a grand interdisciplinary challenge requiring further study, identifying these essential properties lays a foundation for future solutions.
+Summary
 To the best of our knowledge, we are the first to use the rich auction data to analyze the decentralization of Ethereum’s builder market, while most previous works are limited to on-chain data.
-We summarize our contribution below.
-  • We found that builder centralization has led to significant proposer losses, which could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
-We challenge the community’s belief that “builder centralization is okay".
+As a result, they cannot identify the primary cause of proposer losses or accurately identify important providers.
+We summarize our contributions below.
+  • We found that proposers incur significant losses in the centralized builder market, which could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
+We challenge the community’s belief that “builder centralization is okay”.
   • We attribute proposer losses to two factors: auctions are not competitive, builders’ block-building capacities are unequal, and inequality is the primary cause.
-  • We identify two reasons for inequality in block-building capacity: providers’ trust concerns and the incentive for integration.
+  • We identify a previously unrecognized integration of top providers, which causes inequality in block-building capacity and significant losses for proposers.
   • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest.
-Our dataset of 8.79 billion partial bids since 2022 is the largest scale known to us.
+Our dataset of 11 billion partial bids since 2022 is the largest scale known to us.
   • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions.
-Currently, there is only limited auction theoretic analysis that considers the complexity of real-world MEV auctions [40, 67, 90].
 The findings in this paper can serve as empirical evidence for future theoretical interrogation.
@@
 This results in 7,200 slots in Ethereum per day.
-During each slot, a single validator, referred to as the proposer, proposes a block, while other validators vote for the chain head in their local view.
+During each slot, a single validator, referred to as the proposer, proposes a block, while others vote for the chain head in their local view.
 Proposer-builder separation (PBS) and MEV-Boost.
@@
 A private channel, or channel for short, connects users directly with builders, bypassing the mempool.
-Channels receive transactions from users and then, forward them to the user-specified builders, offering privacy and atomicity guarantees.
+Channels receive user transactions and forward them to the user-specified builders, offering privacy and atomicity guarantees.
 E.g., a user can send a bundle (list) of transactions and request them not to be unbundled.
@@
 A searcher is said to integrate with a builder if it sends almost all of its orders to that builder exclusively.
-For example, searcher 0xA69b… is considered an integrated searcher of Beaver Builder because Beaver Builder includes the majority of its transactions.
+For example, SCP (addresses are listed in Tab.
+IV) is considered an integrated searcher of the builder Beaver [38].
 Some channels, such as MEV-Share and MEV Blocker, allow searchers to read pending transactions in the channel and extract MEV from them, provided that such extraction should not harm users (e.g., only backrunning is allowed).
@@
 We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
-Note that $EV(B)$ is a lower bound of the builder’s actual income from $B$ because the builder may own accounts out of $A_{p}$ in Equation 1, or even receive payments off-chain.
+Note that $EV(B)$ is an estimation of the builder’s actual income from $B$ because the builder may own accounts out of $A_{p}$ in Equation 1, or even receive payments off-chain.
 We discuss the implications in Sec.
-7.
+6.
 3 Dataset
-We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to July 2024, and a dataset of full bids from April to December 2023.
+We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to October 2024, and a dataset of full bids from April to December 2023.
 The full bids are obtained from the ultra sound relay [15].
@@
 Tab.
-1 summarizes the datasets.
+I summarizes the datasets.
 Dataset Content # of entries Source
-On-chain blocks 4,892,579 Reth node
-transactions 748,859,795
-Auction blocks produced by MEV-Boost 4,305,571 relay API
-partial bids 8,789,557,915 relay API
-full bids from ultra sound 326,574,676 ultra sound relay
-Misc. builder public keys 248 see Sec. 3.1
+On-chain blocks 5,551,675 Reth node
+transactions 852,482,466
+Auction blocks produced by MEV-Boost 4,909,683 relay API
+partial bids 11,966,097,921 relay API
+full bids from ultra sound 326,574,676 ultra sound
+Misc. builder public keys 273 see Sec. 3.1
 builder addresses 102
-searcher addresses 2,140
-private transactions 77,747,623
-transaction sources 60,345,066
-Table 1: Overview of collected datasets.
+searcher addresses 2,172
+private transactions 112,218,755
+transaction sources 74,922,499
+TABLE I: Overview of collected datasets.
 3.1 Data Collection and Validation
 On-chain data.
-We run an Ethereum execution client Reth [68] along with a consensus client Lighthouse [69] to collect blocks and transactions from September 15, 2022, to July 31, 2024.
-The on-chain dataset includes 4,892,579 blocks and 748,859,795 transactions.
+We run an Ethereum execution client Reth [47] along with a consensus client Lighthouse [48] to collect blocks and transactions from September 15, 2022, to October 31, 2024.
+The on-chain dataset includes 5,551,675 blocks and 852,482,466 transactions.
 Data from relay APIs.
@@
 Recall that partial bids do not include the block body.
-We collected 4,305,571 winning bids and 8,789,557,915 partial bids from September 15, 2022, to July 31, 2024.
-We use winning bids to identify blocks produced through MEV-Boost.
-We use partial bids to quantify the importance of private order flow providers in Sec.
+We collected 4,909,683 winning bids and 11,966,097,921 partial bids from September 15, 2022, to October 31, 2024.
+We use winning bids to identify blocks produced through MEV-Boost and partial bids to quantify the importance of private order flow providers in Sec.
 5.
 To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets [49, 29, 50].
-Our dataset includes 720 more winning bids than other datasets combined.
-Upon comparison with on-chain blocks, we find they are all included on Ethereum, thereby indicating a high level of completeness.
-Full bids from the ultra sound relay .
-From the ultra sound relay [62], we obtained the full bids in 301,479 historical auctions that took place between April and December 2023 (ultra sound relay stopped collecting bids at the end of 2023).
+Our dataset includes 720 more winning bids than other datasets.
+Upon comparison with on-chain blocks, we find they are all included on Ethereum, indicating a high level of completeness.
+Full bids from ultra sound relay .
+From ultra sound relay [15], we obtained the full bids in 301,479 historical auctions that took place between April and December 2023 (ultra sound relay stopped collecting bids at the end of 2023).
 Since the data volume is large (we record roughly 200 GB of data per day), our dataset samples one week from each month: April 9-15 and May 1-7 to December 1-7, 2023.
@@
 We use full bids in Sec.
-4 to quantify the competitiveness and efficiency of MEV-Boost auctions.
-The original data from relays includes bids submitted after the deadline because of network latency.
+4 to quantify the competitiveness and efficiency of MEV-Boost auctions, as well as the proposers’ losses.
+The original data from relays includes bids submitted after the deadline due to network latency.
 We removed them as we are only interested in the bids before the auction ends.
@@
 As shown in Fig.
-10, bids from the ultra sound relay covered more than 80% of the builders in over 75% of the MEV-Boost auctions during our study period.
-Moreover, the top-5 builders by market share (Flashbots, Beaver, Rsync, Builder0x69, and Titan) almost always submitted to ultra sound relay.
+10, bids from the ultra sound relay covered more than 80% of the builders in over 75% of the MEV-Boost auctions from April 2023 to December 2023.
+Moreover, the top-5 builders by market share in this period (Flashbots, Beaver, Rsync, Builder0x69, and Titan) almost always submitted to ultra sound relay.
 Miscellaneous data.
@@
   • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx.
-B for details).
-    To validate the completeness of our dataset, we manually cross-check it with public datasets [72, 65] and related works [42, 41], and the results show that our dataset includes 119 more public keys, so our dataset is the most complete one so far.
+A for details).
+    To validate the completeness of our dataset, we manually cross-check it with public datasets [52, 53] and related works [43, 38], and the results show that our dataset includes 154 more public keys, so our dataset is the most complete one so far.
   • Builder addresses: We collect the builder addresses from the blocks’ last transactions combined with the public dataset [54].
 Similarly, to validate their completeness, we manually cross-check them with related works [43, 38], and the results show that our dataset includes 86 more builder addresses, which is the most complete so far.
-  • Searcher addresses: We obtained known searcher addresses from Etherscan [25], libMEV [50], and related work [42].
+  • Searcher addresses: We obtained known searcher addresses from Etherscan [55], libMEV [56], and related works [38, 57].
   • Private transaction hashes: We use data from Mempool Guru [58] to label transactions that bypassed the public mempool as private.
@@
 We identified transaction sources using ground truth data (as described in Appx.
-B).
-4 Implications of Builder Centralization
+A).
+4 Proposer Loss in the MEV-Boost Auctions
 In this section, we focus on the status quo and study the implications of a centralized builder market as we have today.
-In particular, we are interested in the losses incurred by proposers, as significant losses could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
-We quantify proposers’ losses in historical MEV-Boost auctions considering two factors: the competitiveness of the auctions and the inequality in block-building capacity.
-Our study includes some MEV-Boost auctions where the highest bids were not included on Ethereum due to network latency.
-However, we do not omit these auctions since the auctions themselves remain meaningful.
+In particular, we are interested in the losses incurred by proposers in the MEV-Boost auctions, as significant losses could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
+In principle, builder centralization does not imply a significant proposer loss, as optimal outcomes require only a few non-colluding builders with similar block-building capacities to maintain competitive auctions.
+To evaluate these two conditions in the centralized builder market, we quantify proposers’ losses in historical MEV-Boost auctions by considering two factors: the competitiveness of the auctions and the inequality in block-building capacity.
 4.1 Competitiveness of MEV-Boost Auctions
@@
 For example, suppose executing $B$ yields a net profit of $0.1$ ETH after paying $0.9$ ETH in bid value in the auction, then the true value (the revenue) is $1.0$ ETH.
-To compute $TV(B)$ of a given block $B$, we follow the standard approach in open source MEV analysis tools [20, 31] and previous works [41, 80], though we built our own analysis toolchain and optimized its performance to handle the large volume of auction data.
-We compute $EV(B)$ using a modified Ethereum execution client (Reth [68]) to execute transactions in $B$ and compute the net ETH balance changes of the bidder’s addresses.
+To compute $TV(B)$ of a given block $B$, we follow the standard approach in MEV analysis tools [41, 42] and previous works [43, 44], though we built our own analysis toolchain and optimized its performance to handle the large volume of historical auction data (326.6 million full bids).
+We compute $EV(B)$ using a modified Ethereum execution client (Reth [47]) to execute transactions in $B$ and compute the net ETH balance changes of the builder’s addresses.
 Most builders’ addresses are well-known partly because they need to build up a reputation.
@@
 3.1 for details.) Then, with bid value $BV(B)$ from the auction dataset, the true value $TV(B)$ is the sum of $EV(B)$ and $BV(B)$.
-In every slot $s$, bidders can submit multiple bids throughout the duration of the slot.
-We use $BV(p,s)$ to denote the highest bid value from bidder $p$ in slot $s$, and $TV(p,s)$ the corresponding true value.
+In every slot $s$, builders can submit multiple bids throughout the duration of the slot.
+We use $BV(p,s)$ to denote the highest bid value from builder $p$ in slot $s$, and $TV(p,s)$ the corresponding true value.
 Competitive index.
@@
 Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value.
-$CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, satisfying the definition of a competitive auction, while $CI(s)<0$ means insufficient competition.
+$CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, thus satisfying the definition of a competitive auction, while $CI(s)<0$ means there is insufficient competition.
 Efficient index.   Same as above, the efficient index, $EI(s)$, is defined as:
@@
 Fig.
-4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the ultra sound relay dataset.
+4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the our dataset.
 Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
@@
 That is, under the slightly relaxed definition, approximately 95% auctions were competitive.
-In the interest of space, we leave the analysis of efficiency and the implication of MEV on efficiency to Appx. D.1.
 MEV’s implication on auction outcomes.
-To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we categorize every slot into one of three tiers according to the winning bid (as a proxy to the significance of the MEV opportunity in that slot): low MEV (below 0.0215 ETH, the 10th percentile of the winning bid values), high MEV (above 0.3933 ETH, 95th percentile) and medium MEV in between.
+To further understand how the magnitude of MEV in a slot affects the competitiveness and efficiency of the auction, we categorize every slot into one of three tiers based on the winning bid (as a proxy for MEV significance in that slot): low MEV (below 0.0215 ETH, the 10th percentile of the winning bid values), high MEV (above 0.3933 ETH, 95th percentile) and medium MEV in between.
 The ratio of competitive auctions in each tier is shown in Fig.
@@
 Finding I: MEV-Boost auctions are generally competitive, while the competitiveness of MEV-Boost auctions deteriorates as a slot’s MEV increases.
+In the interest of space, we leave the analysis of efficiency to Appx. D.1.
 Loss from uncompetiveness.
 For an uncompetitive auction, the loss from uncompetitiveness is the difference between the second-highest true value and the actual winning bid value, representing the loss a validator incurs compared to what it would have received in a competitive auction.
-We computed the loss from uncompetitiveness using our dataset, and the results are shown in Tab.
-2.
-The loss from uncompetitiveness ranges from 0.5% to 1.7% during the periods covered.
-Table 2: Losses from uncompetitiveness and inequality.
+We computed this loss using our dataset, and the results are shown in Tab.
+II.
+The loss from uncompetitiveness ranges from 0.5% to 1.7% during our study periods.
+TABLE II: Losses from uncompetitiveness and inequality.
 Time Slots Profits (ETH) Losses-un (ETH) (%) Losses-in (ETH) (%)
@@
 A larger QCD indicates greater dispersion (the intuition is that in a dataset with big dispersion, $Q_{3}$ is much larger than $Q_{1}$, thus $QCD$ approaches $1$).
-Since it uses quartiles, QCD is robust to outliers [88].
+Since QCD uses quartiles, this metric is robust to outliers [16].
 We now present the result given by QCD, noting that other metrics lead to a similar conclusion (see Sec. D.2).
@@
 First, within each MEV tier, there is a clear increase in inequality (a high QCD means high inequality) from the top to the middle and then to the tail.
-This suggests that while top builders have comparable block-building capacities, middle and tail builders have significantly different capacitiesx.
+This suggests that while top builders have comparable block-building capacities, middle and tail builders have significantly different capacities.
 This might be good news: top builders having similar abilities means they can meaningfully compete, which is positive for the decentralization of the builder market.
 Second, the inequality worsens as the MEV tier increases.
-One possible reason is that providers may give significant MEV opportunities to select builders, as we will discuss in Sec.
-5.3.
+One possible reason is that providers may give significant MEV opportunities to select builders, as we will discuss in the next section.
 Figure 6: Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders.
@@
 As shown in Tab.
-2, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset.
-Comparing the two types of loss, we find that inequality is the proposers’ primary source of loss.
-Note that the time range of the full bids is from April to December 2023, and the inequality in the current builder market may worsen (as discussed in Sec.
-5.3).
-5 What Caused Inequality in Block-building Capacities
-Given the significant losses proposers incur due to the inequality in builders’ block-building capacity, we further investigate the causes of this inequality in this section.
-Although the builder market is technically open to all, the barrier to accessing profitable private order flows (Sec. 2.2) forms an implicit access barrier.
-To illustrate the importance of private order flow to the success of builders, Fig.
-7 plots the fraction of builder income from private order flows.
-Currently, approximately 80% of the block value comes from private order flows.
-Figure 7: Fraction of builder income from private order flows.
-Finding III: The private order flow is crucial for builders to win the MEV-Boost auction because it contributes to more than 80% of the block value in over 50% of blocks.
-Where can new builders get private order flows?
-The entities who provide (or sell) private order flows to builders are called private order flow providers.
-Different providers impose different access barriers to their private order flows.
-Some providers, like MEV Blocker or MEV-Share, are open to new builders who satisfy certain requirements, while some (e.g., most searchers) are almost opaque to the public.
-To understand what causes the access barrier for new builders, we first identify important “pivotal” providers and then analyze the requirements to access private order flows from these representative pivotal providers.
+II, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset.
+Comparing the two types of loss, we find that inequality is the primary source of proposers’ loss.
+4.3 Proposer Losses Over Extended Periods
+Note that the time range of the full bids is from April to December 2023, raising the natural question of whether proposers may incur more significant losses in 2024.
+Recall that, in an MEV-Boost auction in the slot $s$, the proposer’s loss is calculated as $TV(p_{1},s)-BV_{w}(s)$.
+Unfortunately, without the full bid dataset, we cannot compute the highest true value $TV(p_{1},s)$ and, therefore, the proposer loss.
+However, since the winner’s true value cannot exceed the highest true value (otherwise, it would be the highest itself), we can compute a lower bound for the actual proposer losses using $TV_{w}(s)-BV_{w}(s)$, which we refer to as the estimated proposer loss.
+A significant estimated proposer loss suggests that proposers incur substantial losses.
+We calculate the daily estimated proposer loss by aggregating the estimated losses from all MEV-Boost auctions each day and comparing this total to the overall block value for that day.
+As shown in Fig.
+7, there is a clear upward trend in the daily estimated proposer loss in 2024—the median was 4.1% before, compared to 21.2% afterward!
+The daily estimated proposer loss is significant on August 6, 2024, primarily because, in an auction on that date (slot 9676257), the winner’s bid is 11 ETH while the true value is 803 ETH—resulting in a significant loss for the proposer.
+Figure 7: Daily estimated proposer loss over two years. Proposers suffered significant losses in 2024.
+Why are proposer losses more significant in 2024?
+One possible explanation is that worsening inequality in 2024 amplifies proposer losses.
+To explore this, we need to deepen our understanding of inequality in the builder market.
+5 Private Order Flows
+In the previous section, our study on historical auctions reveals that inequality is a primary source of proposers’ losses.
+Recall that the inequality loss comes from the disparity in builders’ true values—a bid’s true value is computed by the total value extracted from the transactions.
+As introduced in Sec.
+2.2, transactions in Ethereum have two ways to reach a builder: through the public peer-to-peer network or via a direct channel.
+The former is publicly available to all, which is less concerning, while the latter is restricted, providing access only to select builders.
+In this section, we further study the critical factor contributing to inequality—the difference in private order flows that builders receive.
+Figure 8: Fraction of block value from private order flows.
+Criticalness of private order flows.
+To illustrate the importance of private order flow to builders, we compute the ratio of block values derived from private transactions.
+We first identify all private order flows using data described in Sec.
+3.1.
+Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total priority fees [45] and direct transfers [46] to the block builder) to the block value.
+Fig.
+8 plots the fraction of block value from private order flows over two years.
+As of October 2024, approximately 80% of the block value comes from private order flows for the majority of the blocks, suggesting the criticalness of private order flows!
+Finding III: Private order flows are critical to builders’ block-building capacities since they contribute to more than 80% of the block value in over 50% of blocks.
 5.1 Pivotal Private Order Flow Providers
+Private order flows are critical to builders, and they can drive significant differences in builders’ block-building capacities, contributing to inequality.
+Therefore, analyzing them is essential for understanding such inequality, despite the challenges posed by the vast data volume and extended time span (112 million private transactions over two years).
+Our first step is to classify the large volume of private order flows based on their sources.
+The entities who provide (or sell) private order flows to builders are called private order flow providers, or providers for short.
 Prominent private order flow providers are listed in Tab.
-5.
+V.
 Based on their roles in the MEV supply chain (see Fig.
 3), we categorize them into three types: searcher, Telegram trading bot, and channel.
-We now present our method to quantify the importance of individual providers.
+The process of identifying the order flows from these providers, as well as the details of the finalized dataset, are described in Sec.
+3.1.
+Not every provider is equally important.
+Some providers generate more value than others and their order flows can significantly impact builders.
+Therefore, our next step is to identify these important providers.
 An existing approach [18, 19] is to compute the contribution of each provider to the builders’ income (Sec.
 2.3) during a specific period.
-However, this approach only considers the magnitude of impact but neglects the duration (in terms of time or the number of auctions).
-For example, a single private transaction, such as one3330xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip, could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot.
-Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that can affect half of the MEV-Boost auctions in over two weeks, thereby capturing their sustained impact.
+However, this approach can have false positives; for example, it could identify providers that were only influential temporarily.
+For example, a single private transaction3330xf0464b01d962f714eee9d4392b2494524d0e10ce3eb3723873afd1346b8b06e4 transferring 457.55 ETH to the builder as a tip could contribute significantly to the builder’s income, but this provider may only influence the MEV-Boost auction in a single slot.
+Thus, we assess the significance of a provider as the number of auctions for which they played an important role; in particular, we focus on the providers that have sustained impacts (e.g., they can affect half of the MEV-Boost auctions in over two weeks).
 We define a provider $P$ as pivotal for an instance of MEV-Boost auction if removing $P$’s transactions from the winning block causes the winner to lose the auction.
@@
 Note that there can be multiple pivotal providers for an auction.
-Figure 8: A provider $P$ is pivotal if the winner of an auction could have lost without order flows from $P$.
-The pivotal level of top-5 pivotal providers over time.
-A line may not start from the beginning because the provider did not exist at that point.
-Maestro became an independent provider around Dec 2023.
+Pivotal level.
 For each pivotal provider, we compute the fraction of auctions in which it is pivotal.
 We call this metric the pivotal level, a number from 0 to 1.
+A higher pivotal level indicates a provider’s greater significance in the MEV-Boost auctions.
 Fig.
 9 plots the pivotal level of the top-5 providers.
-A builder who lacks access to any of their order flows will lose the majority of auctions.
-However, we found that order flows from these providers are inaccessible to all builders, which requires further investigation.
-Other observations from Fig.
-8 are provided in Appx.
-F.2.
-5.2 Trust Barrier
-A direct reason for the access barrier is that private order flow providers are at a disadvantage when interacting with builders and must trust builders not to unbundle the transactions.
-However, in practice, builders occasionally unbundle searcher’s bundles.
-For example, the f1b builder unbundled a bundle from the searcher jaredfromsubway.eth at block height 18476515.
-Although the searcher was compensated afterward, this incident highlights the fragile security foundation of the MEV-Boost auction.
-Due to the lack of trustworthy fair exchange mechanisms, providers have to rely on an informal reputation system for self-protection, sending order flows only to builders who meet their requirements.
-To understand the access barrier caused by the trust crisis, we investigate different providers and their requirements for accessing order flows.
-Searchers.
-Since searchers are operated by anonymous entities without any public communication channels, there is no public statement about which builders may access their flows.
-Existing studies indicate that searchers tend to send their order flow to the top builders [81] or share it exclusively with vertically integrated builders [81, 80, 40].
-Telegram trading bots.
-The requirements for Telegram trading bots, in terms of sending order flow to builders, depend on whether they utilize channels like MEV Blocker [56, 48].
-If so, their requirements are aligned with those of these channels.
-Banana Gun does not explicitly state the requirements for builders but we observe its preference for Titan, which we will discuss later.
-Channels.
-Since channels are public-facing services, we were able to extract their access requirements from the documentation and/or through their communication channels.
-The requirements of each provider are summarized in Tab.
-5.
-The channels impose various requirements on a builder’s reputation.
-Market share—the percentage of blocks built by a given builder—is a metric most relay providers use to evaluate builders’ reputations.
-Intuitively, a builder with a high market share, seen as reputable, is less likely to compromise private order flows as doing so will forgo their future revenue.
-MeowRPC [58] sends the order flow to trusted builders with high market share.
-Similarly, MEV-Share [37] requires that a builder possesses a competitive market share.
-MEV Blocker [60] requires that a builder must maintain at least 1% market share over each week of one month.
-Subsidization.
-New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share.
-In practice, builders pay out of pocket to overbid in MEV-Boost auctions, a practice known as block subsidization [36], representing an entry barrier for new builders.
-We quantify the minimal cost for a new builder to meet the MEV Blocker’s requirements (Appx.
-F.5), which reflects the costs of building a reputation.
-The minimal cost continues to increase and can be up to 7.4 ETH.
-This cost represents a lower bound since we do not have the corresponding numbers for other providers.
-Note that although subsidization can help meet the requirements set by channels, it cannot ensure that searchers’ trust concerns are addressed, allowing builders to receive their order flows.
-5.3 Incentive for Integration
-Another access barrier to private order flows is that providers exclusively share order flows with specific builders, known as integration [40].
-We compute each provider’s builder-specific pivotal level to understand the relationship between providers and builders.
-For a given provider $P$, its pivotal level specific to builder $B$ is the pivotal level of the MEV-Boost auctions where $B$ is the winner.
-Figure 9: Weekly builder-specific pivotal level for top-3 builders. MEV-Blocker and MEV-Share are hidden for clarity.
-Fig.
-9 plots selected providers’ builder-specific pivotal levels for the top three builders, ranked based on their weekly market share in July 2024.
-We include providers whose weekly pivotal levels exceed 25% for any of the three builders over two months to illustrate these builder-specific pivotal levels.
-MEV Blocker and MEV-Share are excluded from this analysis, as all builders connected to them reportedly receive the same order flows.
-We note that some providers have a pivotal level exceeding 50% only with specific builders, suggesting a possible collaborative relationship.
-The searcher 0xA69b… consistently influences over 50% of the MEV-Boost auctions won by Beaver, while the searchers 0x0087…, 0x2800…, and 0x51c7… successively affect those won by Rsync.
-These pivotal providers are all integrated searchers of the corresponding builders that have been identified in previous works [42, 40].
-This observation corroborates findings from previous studies [42, 40, 80], confirming that the order flows from integrated searchers are crucial to those builders.
-Furthermore, we observe a clear variation in the importance of providers for different builders.
-For example, the significance of Banana Gun to Titan is greater than it is to other builders.
-We further investigate the relationship between Banana Gun and Titan and suppose that Banana Gun only sends valuable order flow to Titan when MEV is significant.
-A case in this point is slot 8614161, where only Titan received the valuable order flow from Banana Gun.
-The much lower bids from other builders suggest they didn’t receive this order flow; otherwise, they could have outbid Titan.
-We illustrate this new integration pattern in Appx.
-F.3.
-The results in Fig.
-15 indicate that this integration pattern also applies to jaredfromsubway.eth and Beaver.
-Both integrations emerged around January 2024.
-There is a clear incentive for integration between providers and builders.
-If MEV from a provider can determine the auction outcome, sending order flow to an integrated builder is better than sending it to multiple builders.
-The integrated builder only needs to bid slightly higher to win, and they can share the surplus MEV.
-However, this integration is detrimental to proposers, as it exacerbates the inequality in builders’ block-building capacities and can eventually lead to significant losses for them.
-Finding IV: Two providers (Banana Gun and jaredfromsubway.eth) exclusively share high-value order flows with specific builders (Titan and Beaver) respectively, suggesting a new integration pattern.
+A builder who lacks access to any of their order flows would lose the majority of auctions.
+However, order flows from the pivotal providers may not be accessible to all builders, which requires our further investigation.
+We also make three observations from Fig.
+9, which are presented in Appx.
+E.1 due to space constraints.
+Figure 9: The pivotal level of top-5 pivotal providers over time.
+A provider $P$ is pivotal if the winner of an auction could have lost without private order flows from $P$.
+A line may not start from the beginning because the provider did not exist at that point.
+Accessibility by multiple builders.
+Among the top five pivotal providers, MEV-Share and MEV Blocker are public-facing services with clear documentation and open communication channels.
+This enables us to analyze their statements regarding how their order flows are distributed to different builders—both claim that every connected builder receives the order flows [35, 36].
+Although their order flows are not accessible to all builders (as summarized in Tab.
+V), top builders such as Beaver, Titan, and Rsync have access to them.
+Ideally, competition among these top builders will ensure that the value from these order flows is ultimately captured by the proposers.
+Notably, while these order flows are distributed equally to all connected builders, inequality still exists.
+Order flow providers may exclude less reputable builders due to trust concerns.
+Although this exclusion may not harm proposers, it could exacerbate builder centralization, as new builders must expend significant effort to meet the requirements imposed by these pivotal providers.
+Due to page constraints, we provide a more detailed discussion of the trust barrier and how it prevents builder decentralization in Appx.
+E.2.
+5.2 Exclusive Pivotal Providers
+Pivotal providers like MEV-Share and MEV Blocker send their order flows to connected builders who satisfy their requirements.
+Many other providers (e.g., most searchers) do not publicly share their order flows; some only send order flows exclusively to specific builders.
+If multiple builders receive the same order flows, the proposer can, in principle, capture the value due to the competition.
+In contrast, a builder receiving exclusive order flows is detrimental to proposers, as it can exacerbate inequality and ultimately lead to losses for proposers.
+The exclusive private order flows from a provider to a builder indicate a special relationship called integration [20].
+In this section, we further study the integration between builders and providers in the current market and how much loss proposers incur due to the integration.
+Case study of block 21,000,813.
+To better understand exclusive order flows and their impact, we go through a case study of block 21,000,813, which has the highest estimated proposer loss in October 2024.
+In the auction for this block, Titan received a valuable order flow from Banana Gun (55.5 ETH) and bid 2.8 ETH, while the second-highest bid was only 0.028 ETH.
+The significantly lower bids from other builders suggest they did not receive this order flow; otherwise, they could have bid closer to Titan.
+In this case, the order from Banana Gun dominated the auction, as whoever had that order could win.
+Thus, there is a clear incentive for the provider when its order flows are highly profitable: sending order flow to a select builder is more advantageous than distributing it to multiple builders.
+The select builder only needs to bid slightly higher to win, allowing the provider and the builder to share the surplus MEV (52.7 ETH in this case).
+New integration pattern.
+The above case highlights a potential integration between Titan and Banana Gun.
+Previous studies [20, 38] analyze an integration pattern between builders and searchers engaged in arbitrage between centralized exchanges (CEXs) and decentralized exchanges (DEXs).
+These searchers may be run by the same party as their integrated builders and exclusively send order flows to them, making the integrated builders more likely to win auctions when MEV from CEX-DEX arbitrage is significant.
+Unlike known integrated searchers, Banana Gun is operated by a different party from Titan and typically sends order flows to multiple builders in most cases.
+For example, in October 2024, Banana Gun’s order flows were included in more blocks built by other builders than in those built by Titan.
+This suggests that Titan may only receive exclusive order flows from Banana Gun when those flows are particularly profitable.
+Similarly, other pivotal providers may tend to integrate with top builders when their order flows are highly profitable due to incentives.
+To confirm this conjecture and evaluate the extent to which proposer loss is related to integration, we need to investigate exclusive order flows from providers to builders.
+Challenges of identifying exclusive order flows.
+Identifying exclusive order flows is challenging, as we do not know the exact order flows each builder receives.
+The on-chain blocks only record the winning bids of MEV-Boost auctions, which reflect the order flows received by the winners.
+Even though we have the full bids dataset from 2023, it remains unclear whether the builders did not receive an order flow or chose to exclude it.
+Analyzing the number and value of transactions sent from providers to builders is also unreliable because providers might selectively send exclusive order flows, and the difference may not be very significant.
+Existing work [20] do not provide a methodology for identifying exclusive order flows, while [38] identifies integration by analyzing the correlation between the volume of searchers’ CEX-DEX arbitrage and the builders’ daily market share of blocks.
+This approach cannot be directly applied to identify generalized integration between providers and builders, as different providers may perform various MEV activities beyond CEX-DEX arbitrage.
+Most importantly, since we are interested in how many proposer losses are related to integration, identifying a builder-provider integration pair alone cannot accurately answer this question.
+As discussed earlier, a provider may send order flows exclusively to one builder under specific conditions.
+Therefore, attributing all proposer losses in a builder’s blocks to its integration might be inaccurate.
+before 2024 2024
+Provider MEV Activity Builder Blocks (#) Losses (ETH) (%) Provider MEV Activity Builder Blocks (#) Losses (ETH) (%)
+SCP CEX-DEX arbitrage Beaver 18,509 4,149.3  (21.0) Banana Gun Sniper Titan 33,748 19,401.3  (37.1)
+Searcher: 0x57c1… CEX-DEX arbitrage Addr: 0x4737… 7,166 1,037.3  (5.2) SCP CEX-DEX arbitrage Beaver 10,087 3,249.8  (6.2)
+Wintermute CEX-DEX arbitrage Rsync 6,555 553.1  (2.8) jaredfromsubway.eth Sandwich Beaver 20,213 2,575.4  (5.0)
+Searcher: 0xf8b7… CEX-DEX arbitrage Manta 4,606 428.2  (2.2) Searcher: 0x6980… Atomic arbitrage I can haz block? 513 2,484.9  (4.8)
+Searcher: 0x9dd8… Atomic arbitrage Addr: 0x3bee… 1,600 411.5  (2.6) Searcher: 0x6454… Liquidation Addr: 0x3bee… 793 1,229.6  (2.4)
+Searcher: 0x761d… Atomic arbitrage builder0x69 12 287.0  (1.5) c0ffeebabe.eth Liquidation Ty For The Block 367 1,003.6  (1.9)
+Searcher: 0x6454… Liquidation Addr: 0x3bee… 539 262.5  (1.5) Banana Gun Sniper Beaver 2,430 987.4  (1.9)
+Searcher: 0x5e51… Atomic arbitrage eth-builder 266 223.2  (1.1) WINTERMUTE CEX-DEX arbitrage Rsync 6,503 976.8  (1.9)
+Searcher: 0xeeaa… Atomic arbitrage eth-builder 51 194.1  (1.0) Maestro Sniper Beaver 10,103 520.3  (1.0)
+Searcher: 0xe20c… Atomic arbitrage Addr: 0x5638… 579 168.5  (0.9) Searcher: 0x77d5… Atomic arbitrage Rsync 782 345.3  (0.7)
+TABLE III: The top 10 providers sending exclusive order flows and their corresponding builders, contribute to the highest estimated proposer loss (before and in 2024).
+Our approach.
+Despite the challenges mentioned above, we are still able to infer whether an order flow is likely to be exclusively sent to a builder based on a reasonable assumption: all builders are rational and do not collude, so they will compete upon receiving the same order flow.
+We justify this assumption with our study in Sec.
+4, which shows that over 90% of MEV-Boost auctions are competitive.
+Based on this assumption, we design an algorithm to infer if an order flow is exclusively sent to a builder.
+For a given slot, we obtain the winning bid and compute the second-highest bid from our partial auction dataset.
+Then, for the winning bid in this slot, we identify the private order flows within this block and their associated providers using our private transaction dataset.
+For each provider, we evaluate whether the combined value of the private order flow from this provider and the public order flow exceeds the second-highest builder’s bid multiplied by a threshold.
+This threshold, set at 1.15, was derived from our analysis of historical MEV-Boost auctions, where less than 5% of builders withheld bids that exceeded this value.
+The threshold accounts for the possibility that builders may bid slightly below their true valuations.
+If this condition holds, it suggests that the provider exclusively sends order flows to the builder, thus giving the winner a competitive advantage.
+Since there is no ground truth for exclusive order flows, we validate our method using four builder-searcher integrations identified in previous work [38].
+Our algorithm cannot identify exclusive order flows when their value is insignificant, i.e., when they cannot determine the auction outcome.
+Thus, our approach will have false negatives.
+We focus on precision, specifically assessing, for each provider, the percentage of exclusive order flows correctly attributed to their well-known integrated builder.
+For the known builder-searcher integrations—Beaver with SCP, Rsync with Wintermute, Manta with searcher 0xf8b7…, and anonymous builder 0x4737… with searcher 0x57c1…—our methodology achieves accuracies of 99%, 96%, 99%, and 99%, respectively.
+Except for Rsync and its integrated searcher, our methodology achieves very high precision: less than 1% of order flows are inaccurately identified as exclusive (false positive).
+Further investigation into the identified exclusive order flows of Wintermute shows that about 70% of the false positives occurred in blocks built before Rsync began block-building.
+At this point, Rsync might send order flows to multiple builders.
+However, only one builder included the order flow before the auction concluded.
+Besides the known integrations, we also identified previously unknown integrations involving the top-5 pivotal providers—Banana Gun, Maestro, and jaredfromsubway.eth—as listed in Tab.
+III.
+Proposer loss due to integration.
+For each builder-provider integration pair identified by our algorithm, we count the number of blocks where the builder exclusively receives the order flows and calculate the total estimated proposer loss in these blocks.
+Tab.
+III lists the top 10 providers and their corresponding builders receiving exclusive order flows, ranked by the resulting losses before and in 2024 respectively.
+We classify each provider’s MEV activity using previous studies [20, 38] and MEV analysis tools [59, 56].
+For simplicity, we label each searcher with only its predominant MEV activity.
+Telegram bots, typical sniper bots [60], automatically assist users in quickly buying new tokens at launch by monitoring liquidity pools.
+We find that the top four exclusive providers causing the highest proposer loss before 2024 are integrated searchers engaged in CEX-DEX arbitrage.
+This observation aligns with previous studies on competitive advantages and the profits associated with integration between these searchers and their integrated builders [61, 38], highlighting our algorithm’s effectiveness in identifying exclusive order flows.
+Compared to the exclusive order flows before 2024, a more concerning problem is the formation of more exclusive relationships between top builders and top providers.
+This also confirms our conjecture.
+For instance, Banana Gun exclusively sends order flows to Titan in approximately 33K blocks, accounting for 37.1% of the estimated proposer losses covered in our study.
+Additionally, two other top-5 pivotal providers, jaredfromsubway.eth and Maestro, also exclusively send order flows to Beaver, contributing to 6% of the losses.
+In contrast, the total losses caused by the integrated CEX-DEX arbitrage searchers remain relatively similar before and during 2024.
+Notably, searcher 0x57c1… and searcher 0xf8b7… stopped operating in 2023.
+These results effectively explain that compared to the previous known integration, the emergence of new integration accounts for the majority of the proposer losses in 2024.
+Finding IV: Top pivotal providers Banana Gun, jaredfromsubway.eth, and Maestro integrate with the top two builders, Titan and Beaver. Their integration causes significant losses to proposers.
+Competing providers.
+In addition to noting that three of the top five pivotal providers have started integrating with Beaver or Titan, we also observe that providers engaging in the same MEV activities tend to avoid integrating with the same builder simultaneously (Banana Gun sent order flows exclusively on Beaver in January 2024, while its flows to Titan occurred after this period).
+A possible explanation is concern over trust: if a provider knows that a builder collaborates with its competitor, it may worry about the security of its order flow sent to that builder.
+The builder could potentially leak information about the provider’s order flow to the competitor.
+6 Discussion
+6.1 Implications of Integration
+Beyond proposer loss, the integration carries additional implications, which we discuss below.
+Inaccurate MEV analysis.
+Our study reveals a potential pattern between providers and builders: when the order flows from a provider may determine the auction outcome, the provider exclusively shares the order flow with a specific builder.
+In return, the provider may receive a refund from the builder.
+However, this refund could be off-chain, making it difficult to measure the providers’ actual costs.
+This limitation could lead to inaccuracies in MEV analysis tools.
+For example, the profit of a top MEV searcher, jaredfromsubway.eth, is only 1.3% of its total revenue according to [56]; its actual profit may be higher due to potential refunds.
+Suboptimal block-building.
+In principle, a builder with access to more profitable order flows can achieve better block-building.
+However, integration between top providers and builders can prevent builders from maximizing potential MEV, resulting in suboptimal block-building in the winners’ blocks and ultimately reducing profits for providers, builders, and proposers.
+For example, Beaver’s integrated searcher may identify a good opportunity for CEX-DEX arbitrage, while Banana Gun may also handle a large volume of private order flow exclusively sent to Titan.
+These two order flows are not in conflict; however, due to the builders’ competition, only one can be included in the winning block within the same slot.
+The excluded flow may not be included in the next block because the MEV opportunity could be time-sensitive (e.g., CEX-DEX arbitrages).
+Censorship.
+A security concern of builder centralization is transaction censorship.
+Ethereum’s current censorship resistance relies heavily on the choices of builders, particularly top builders.
+For example, before Beaver stopped censoring in July 2024, over 50% of blocks were built by censoring builders [62].
+Our findings suggest that builder centralization remains difficult to overcome due to integration, potentially exacerbating the censorship problem on Ethereum.
+6.2 Proposed Mitigations and Their Limitations
+We now discuss why the proposed mitigation fails to resolve the identified problems.
+This highlights that addressing the identified problems is a grand challenge that requires further, possibly interdisciplinary, study.
+Removing trust barriers.
+There is a lack of a trustworthy mechanism to facilitate fair exchanges between private order flow providers and builders.
+As a result, some providers require builders to have a minimal market share, and we also observe that competing providers typically do not integrate with the same builder.
+One promising direction to reduce this trust dependency is the use of encrypted mempools [63], where transactions remain encrypted until they are irreversibly included in a block, preventing builders from having complete knowledge of user transactions.
+For example, this can be implemented using Trusted Execution Environments (TEEs) [64] (e.g., SUAVE [65]), though standard challenges such as side channels and covert channels [66] still apply.
+Besides, requiring builders to achieve fair ordering [25, 26], is another viable solution to prevent transaction manipulation, though a first-come-first-served ordering policy may introduce other problems like off-chain latency wars [67].
+Proposed PBS designs.
+The Ethereum community has considered multiple alternative designs to mitigate the centralization and security problems in the out-of-protocol PBS.
+Enshrined-PBS (ePBS) [68] is proposed to decentralize PBS auctions by removing centralized relays from PBS.
+MEV burn, an add-on of ePBS, which helps smooth and redistribute MEV spikes, benefits all ETH holders by burning the winning bid value in the PBS auction [13].
+More recent proposals, including Execution Ticket [22] and Execution Auction [23], aim to further mitigate the effects of MEV on validator centralization by removing execution block proposing rights from validators.
+However, in both eBPS and MEV burn, blocks are still produced through auctions, and the PBS auction can continue to exist as a secondary market in both Execution Ticket and Execution Auction.
+This allows the re-sale of block proposing rights to builders who can build more profitable blocks [22].
+None of these proposals considers the incentives of providers, so the integration problem in PBS auctions remains unresolved.
+Future work.
+Based on the discussion, we identify that an effective mitigation must satisfy at least the following two properties: First, it must provide a security guarantee to providers’ order flows from potential attacks.
+Second, it should consider providers’ incentives, such as offering a game-theoretical guarantee to ensure they no longer have the motivation to continue such integration.
+Identifying these two properties is a step forward, but fully addressing them requires significant further study for future work.
+6.3 Limitations of Our Approach and Data
+We compute the true values of bids following a standard approach as [41, 42, 43, 44].
+This approach relies on identifying builder-controlled addresses, and missing addresses could cause inaccuracies.
+We cross-validate with multiple sources to ensure collected addresses are the most comprehensive to date (see Sec.
+3.1).
+It may underestimate values by ignoring off-chain payments, or overestimate profits if builders refund providers off-chain.
+A builder with large off-chain payments could bid higher than the computed true value, appearing as an extremely large subsidy.
+However, we observe that there are no excessively large subsidies, and most builders bid close to our computed true values in our study of competitiveness and inequality, suggesting that off-chain payments are not significant.
+The potential refunds to providers do not affect the computation of proposers’ losses because this portion is not allocated to proposers, regardless of the volume of refunds.
+Due to page constraints, we discuss the potential implications of the limitations in Appx.
+F.
 7 Related Work
@@
 Their observations were based on the market share and the distribution of total profits generated within the PBS ecosystem.
-Moreover, several online dashboards [87, 49, 7, 36] offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics.
+Moreover, several online dashboards [29, 70] offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics.
 However, these studies do not quantify the causes of centralization or the security implications of a centralized builder market.
@@
 He found that private order flows account for only 30% of the transactions yet represent 80% of the total value paid to builders.
-Our study highlights a similar result: in over 50% of blocks, private order flows account for more than 60% of the block value.
+Our study highlights a similar result: private order flows account for more than 80% of the block value in most blocks.
 Gupta et al.
-[40] confirmed that builders are more likely to win the MEV auction when their integrated searchers can provide high-value exclusive order flow by effectively exploiting arbitrage between a centralized exchange (CEX) and a decentralized exchange (DEX).
-The dashboard by Winnie [91] and the study by Titan [81] analyzed the vertical integration between searchers and builders through the quantity and value of transactions sent from searchers to builders.
+[20] confirmed that builders are more likely to win the MEV-Boost auctions when their integrated searchers can provide high-value exclusive order flow by effectively exploiting CEX-DEX arbitrage.
 Heimbach et al.
-[42] observed that during their study, Beaver Builder received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH.
+[38] observed that during their study, Beaver received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH.
 This large gap suggests that Beaver might have received significant off-chain profits from its integrated searchers.
-These studies show that three integrated builders (Beaver, Rsync, and Manta) gain a significant advantage in order flow from their integrated searchers.
-Our study made the same observation and further quantified the pivotal level.
+Öz et al.
+[57] employed Linear Discriminant Analysis to identify the integration between builders and providers.
+Our study uncovers previously unknown integration of top providers and quantifies the resulting proposer losses.
 Strategic behaviors in MEV-Boost auctions.
@@
 [72] further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it.
-A recent study [1] empirically investigated timing games in the wild and how block proposal delays affect the block fork rate.
-The results indicate that three validator sets–P2P, Kiln, and Attestant, that are known or suspected of playing timing games–did not delay their block proposals enough to cause their blocks to be forked.
 Wu et al.
@@
 Causes and implications of centralized builder market.
-A provider may exclusively share its order flow with a single builder.
-Previous works such as [46, 38, 40] pointed out that exclusive order flows have potential centralizing effects on the builder market, because builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers and in winning order flow auctions.
+Previous works such as [20, 74] theoretically analyzed the potential centralizing effects of exclusive order flows on the builder market: builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers.
+Capponi et al.
+identified this dynamic as leading to a “Proof-Of-MEV” paradigm [75], where the capacity to capture MEV becomes a critical factor for success in MEV-Boost auctions.
 The decentralization of Ethereum’s builder market not only benefits itself but also positively impacts the decentralization of validators.
 A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization [5].
-We complement these theoretical analyses with empirical studies to analyze potential problems caused by the current centralized builder market.
-Blockchain decentralization metrics.
-Nakamoto coefficient was proposed to describe the minimum number of entities required to disrupt the blockchain’s subsystem [77].
-Other metrics, such as the Gini coefficient [18], Shannon entropy [51], and Herfindahl–Hirschman index [73], have also been utilized to measure decentralization (or the lack of it) in staking distribution [39, 47], mining power [52, 89, 47], and block building [41].
-An SoK paper [93] comprehensively summarized the existing works on blockchain centralization; however, none explored decentralization from the perspective of entry barriers or auctions.
-7 Discussion
-7.1 Mitigation
-Our results suggest that proposers suffer from significant losses because of builder centralization, which can lead to two undesirable outcomes: First, PBS could become unstable.
-Second, MEV measured by the auctions may be inaccurate.
-In this section, we propose a structural change to the existing MEV supply chain Sec. 2.2 and a solution based on the new supply chain structure to address these two problems.
-In the new supply chain, only one builder receives order flows from providers, and proposers can only accept blocks from this builder.
-Protocol.   Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15] The pseudocode for the TEE builderis described in Fig. 17.
-The builder provides two functions.
-First, it opens RPC endpoints to receive the private order flows from the providers.
-Providers send their order flows along with a bid indicating the MEV in their order flows.
-Second, when it receives a signal for block building (e.g., a new slot starts in Ethereum), it will run an ordering algorithm to build a block with optimal MEV based on the received order flows (a reasonably performant ordering algorithm is described in Appx.
-E).
-The TEE builder runs the truthful mechanism in [57]: it refunds all providers for their marginal value contribution to the block.
-For each included order flow, the builder builds an alternative block without that specific order flow to calculate its marginal value contribution by comparing the MEV extracted between these two blocks.
-Next, we outline the key advantages of this solution.
-Security guarantees.
-TEEs offer a secure and isolated environment, ensuring that private order flows remain confidential and protected from external tampering.
-Moreover, order flow providers can utilize remote attestation to verify that the builder running within the TEE is indeed the intended one, thereby effectively mitigating trust concerns.
-Incentive compatibility.
-[57] shows that for providers, truthfully bidding is a dominant strategy when they are refunded based on their marginal value contribution.
-The formal proof and the required assumptions are in [57].
-Here, we offer an intuitive explanation: a provider effectively pays the second price for their order flows when refunding their marginal value contribution.
-Analysis.
-Since truthfully bidding is an incentive-compatible dominant strategy for all providers, the TEE builder can compute the total MEV in a block from the providers’ bids, achieving an accurate MEV oracle.
-Moreover, because the builder program in the TEE can be programmed to pay the proposer fairly, the instability of PBS is also addressed.
-Future works.
-We leave the analysis of whether the new supply chain structure is sustainable in the long term for future work.
-Intuitively, since providers are incentivized to send order flows only to the builder due to refunds, other builders will naturally be driven out of the market, leaving proposers to only accept blocks from the TEE builder.
-Furthermore, since real-world proposers might behave rationally, exploring the new equilibrium that emerges among a rational provider, builder, and proposer is another intriguing question to be addressed.
-Finally, this solution provides an accurate MEV oracle and transparent MEV information.
-These insights open up new possibilities for future MEV mitigation directions.
-7.2 Limitations of Our Approach and Data
-True value.
-We compute the true values of bids using a similar approach as that in MEV analysis tools [20, 31] and previous works [41, 80].
-This approach has a few limitations.
-First, it relies on the knowledge of the set of addresses controlled by each builder.
-Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses.
-We cross-validated our builder address dataset with other sources to ensure completeness (see Sec.
-3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate.
-Secondly, this approach neglects off-chain profits, such as off-chain payments received from searchers.
-Off-chain payments are more likely to happen between integrated builder-searcher pairs.
-We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization.
-However, in our study, we did not observe any excessively large subsidies.
-Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness.
-We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value.
-However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value.
-It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
-Missing pivotal providers.
-Our results in Sec.
-5.1 might omit some pivotal providers since the providers we can identify are a subset of all providers.
-For example, BloXroute (a widely used channel) does not publish the list of transactions they processed.
-Thus, we can’t evaluate its pivotal level.
-This limitation is common in existing studies [53, 54].
-However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec.
-3).
-Thus, the significance of each provider will only be underestimated rather than overestimated.
-Our auction dataset might miss partial bids since they are self-reported by relays.
-Recall that we used the second-highest bid value in an MEV-Boost auction as the threshold to determine a pivotal provider.
-Therefore, underestimating the second-highest bid value might also lead us to underestimate the significance of each provider.
-7.3 Security Problems in PBS
-As we mentioned in Sec.
-5.2, one reason for the access barrier to private order flows is the difficulty of establishing trust.
-A malicious builder, relay, or proposer may attack providers.
-A real-world incident involved a malicious proposer who double-signed a header to unbundle a builder block and further stole MEV from searchers by proposing a different block [61].
-Another security concern is censorship.
-Ethereum’s censorship resistance currently depends heavily on whether builders choose to censor, particularly the actions of the top builders.
-For example, before the top builder, Beaver, stopped censoring in July 2024, over 50% of blocks were censoring [86].
-Our study reveals that builder centralization remains difficult to overcome, potentially exacerbating the censorship problem in Ethereum.
+We complement these theoretical analyses with empirical studies to analyze the causes and potential problems of the current centralized builder market.
 8 Conclusions
-In this paper, we interrogate the causes and implications of builder centralization in Ethereum and challenge the community’s belief that it is acceptable.
-Our empirical study shows that the current blocker market has led to significant profit losses for validators.
-These significant losses could lead to PBS’s instability and MEV oracles’ inaccuracy and, if left uncontrolled, can lead to validator centralization.
+In this paper, we challenge the community’s belief that builder centralization is acceptable by empirically quantifying the significant profit losses incurred by validators in the current builder market.
+These significant losses could lead to PBS’s instability and MEV oracles’ inaccuracy and, if left uncontrolled, may lead to validator centralization.
 We attribute the loss to two factors: insufficient competition and inequality of block-building capacities, with the latter being the dominant reason.
-We further investigated the trust and incentive issues of the current MEV supply chain that contribute to builder centralization.
-To address these two problems, we propose a structural change to the current MEV supply chain and a solution based on this new structure.
+We further investigated the incentive issues of the current MEV supply centralization that contribute to the builder centralization and significant proposer loss.
+Finally, we identify the properties required to address these issues and hope these insights will benefit future work.
 Acknowledgements
@@
 Builder public keys.
-We collect builder public keys from two data sources.
+We collect builder public keys from three data sources.
 First, builders often disclose their public keys in official documents.
@@
 For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’.
-Given that such marks are unauthenticated and susceptible to impersonation, we ignore them unless they are used more than 100 times in the bids by the same builder public key.
+We manually group builder public keys based on these marks and exclude the public keys which are known as impersonation [77].
 Transaction sources.
-We identified transactions originated from Telegram trading bots, including Maestro [55], Banana Gun [5], and Unibot [82], by searching for transactions that interacted with their official smart contracts.
+We identified transactions originated from top Telegram trading bots, including Maestro [78], Banana Gun [79], and Unibot [80], by searching for transactions that interacted with their official smart contracts.
 Two channels, MEV Blocker and MEV-Share publish data [81, 82] about transactions submitted to them.
-Specifically, transactions submitted to MEV Blocker are publicly available on Dune [70], from which we extracted 22,819,437 transaction hashes that fall within our study’s range.
-We obtained 20,334,513 transactions submitted to MEV-Share from the public Flashbots dataset [33].
+Specifically, transactions submitted to MEV Blocker are publicly available on Dune [81], from which we extracted 28,233,223 transaction hashes.
+We obtained 24,449,086 transactions submitted to MEV-Share from the public Flashbots dataset [82].
 These transaction labels are sourced from their official datasets and we view them as ground truth.
-Appendix C List of Searchers and Builders
-In Tab.
-3, we provide the list of searchers and their addresses mentioned in our study.
-In Tab.
-4, we have counted the top 25 builders in the market, ranked by their market share from April 2023 to December 2023.
-Table 3: The list of searchers
+Appendix C List of Searchers
+In Tab. IV, we provide the list of searchers and their addresses mentioned in our study.
+TABLE IV: The list of searchers
 Searcher Address
 jaredfromsubway.eth 0xae2Fc483527B8EF99EB5D9B44875F005ba1FaE13
-Searcher: 0xA69b… 0xA69babEF1cA67A37Ffaf7a485DfFF3382056e78C
-Searcher: 0x0087… 0x0087BB802D9C0e343F00510000729031ce00bf27
-Searcher: $0$x280027dd00eE0050d3F9d168EFD6B40090009246
-Searcher: 0x51c7… 0x51C72848c68a965f66FA7a88855F9f7784502a7F
-Table 4: The percentages of blocks and MEV-Boost blocks built by each builder from April to December 2023.
-Builder % of total blocks % of MEV-Boost blocks
-Beaver 22.53 24.69
-Rsync 20.75 22.74
-builder0x69 12.65 13.86
-Titan 12.35 13.53
-Flashbots 12.24 13.42
-BloXroute 2.14 2.34
-f1b.io 1.66 1.81
-Blocknative 1.40 1.53
-eth-builder 0.77 0.84
-BuildAI 0.73 0.80
-Gambit Labs 0.64 0.70
-Jetbldr 0.47 0.51
-Manta 0.46 0.51
-payload.de 0.39 0.42
-boba-builder.com 0.27 0.30
-Penguinbuild 0.23 0.25
-Eden 0.22 0.24
-lightspeedbuilder 0.21 0.23
-0x3be…436 0.16 0.18
-lokibuilder.xyz 0.12 0.13
-bobTheBuilder.xyz 0.11 0.13
-blockbeelder.com 0.11 0.12
-finest artisanal blocks 0.09 0.10
-It’s Free Real Estate 0.08 0.09
-nfactorial.xyz 0.06 0.07
-Appendix D Implications of Builder Centralization
+SCP 0xA69babEF1cA67A37Ffaf7a485DfFF3382056e78C
+0x56178a0d5F301bAf6CF3e1Cd53d9863437345Bf9
+0xa57Bd00134B2850B2a1c55860c9e9ea100fDd6CF
+0x4Cb18386e5d1F34dC6EEA834bf3534A970a3f8e7
+0x4Cb18386e5d1F34dC6EEA834bf3534A970a3f8e7
+0x5050e08626c499411B5D0E0b5AF0E83d3fD82EDF
+0xFA103c21ea2DF71DFb92B0652F8B1D795e51cdEf
+0x0DA9d9eCeA7235c999764e34F08499cA424c0177
+WINTERMUTE 0x0087BB802D9C0e343F00510000729031ce00bf27
+0x280027dd00eE0050d3F9d168EFD6B40090009246
+0x51C72848c68a965f66FA7a88855F9f7784502a7F
+Appendix D Efficiency and Block-building Capacities
 D.1 Efficiency of MEV-Boost Auctions
@@
 4 shows that 24.9% of MEV-Boost auctions were not efficient.
-In an inefficient auction, the builder with the highest true value does not win the auction.
+An auction is inefficient when the builder with the highest true value does not win.
 There are two possible reasons.
-First, a builder can win the auction without possessing the highest true value by overbidding, known as block subsidization [36].
+First, a builder can win the auction without possessing the highest true value by overbidding, known as subsidization [70].
 Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
@@
 Thus, the cumulative percentage exceeds 100%.
-It is important to note that inefficiency does not necessarily disadvantage the validators. In fact, validators may receive more revenues in inefficient scenarios involving subsidization.
+It is important to note that inefficiency does not necessarily harm validators. In fact, validators may receive more revenues in inefficient scenarios involving subsidization.
 MEV’s implication on efficiency.
@@
 The CV of slot $s$ is computed as $CV(\vec{tv}_{s})$.
-Figure 12: Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of block-build capacity across builders.
+Figure 12: Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of builders’ block-build capacity.
 As shown in Fig.
@@
 Fig. 13 shows the distribution of the competitive index (CI) of the MEV-Boost auctions in our study.
-Figure 13: The distribution of the CIs of MEV-Boost auctions
-Appendix E Reference Builder Algorithm
-Flashbots has open-sourced a Rust implementation of the MEV-Boost builder [34], which serves as a reasonably performant reference for builder algorithms.
-We will briefly introduce their approach.
-The builder processes each bundle (where a single transaction is considered a bundle containing one transaction) by computing its effective gas price and total profit.
-Subsequently, the builder sorts all received bundles in descending order based on either effective gas price or total profit.
-Employing a greedy algorithm, the builder then appends bundles from this sorted list.
-1 presents the pseudocode of their algorithm.
-Listing 1: Flashbots block-building algorithm
-# Compute the effective gas price and profit for a given bundle
-# A single transaction is recognized as a bundle of size 1
-def receive(bundle):
-effective_gas_price, profit = compute_price_and_profit(bundle)
-bundle.egp = effective_gas_price
-bundle.profit = profit
-return bundle
-# Building a block
-def build(bundles, sorting_key):
-# Sort bundles by price or profit
-if sorting_key == ’egp’:
-bundles.sort(key=lambda b: b.egp)
-else:
-bundles.sort(key=lambda b: b.profit)
-# Initialize
-B = []
-# Get the current state
-state = get_current_state()
-# Greedy algorithm
-while not bundles.is_empty():
-# Get the first bundle from the queue
-bundle = bundles.pop(0)
-# Simulate the execution of the bundle
-success = simulate_bundle(state, bundle)
-# Update with the successful bundle
-if success:
-B.append(bundle)
-state.update(bundle)
-# Compute payment based on the final state
-payment = compute_payment(state)
-# Append the payment to the block
-B.append(payment)
-# Return the valid block
-return B
-Appendix F What Caused Inequality in Block-building Capacities
-F.1 Importance of Private Order Flows
-To quantify the importance of private order flows for building valuable blocks, we compute the ratio of block values derived from private transactions.
-We first identify all private order flows using methods described in Sec.
-3.1.
-Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total transaction priority fees [21] and direct transfers [32] to the block builder) to the total values extracted from all transactions in the block.
-F.2 Pivotal Private Order Flow Providers
+Figure 13: Distribution of the CIs of MEV-Boost auctions
+Appendix E Private Order Flows
+E.1 Pivotal Private Order Flow Providers
 We can make a few observations from Fig.
-8.
+9.
 First, a particular searcher, jaredfromsubway.eth, has been gaining prominence since February 2023 and reached its peak significance in May 2023.
@@
 9 that MEV Share and MEV Blocker have been increasingly pivotal since May 2023.
-These two providers positively impact decentralization because they have relatively clear requirements for accessing their order flows (the requirements of different providers will be discussed in Sec.
-5.2).
+These two providers positively impact decentralization because they have relatively clear requirements for accessing their order flows (the requirements of different providers will be discussed next.)
 Third, we observe a notable downtrend in the pivotal level of MEV-Share and MEV Blocker starting from November 2023, accompanied by an uptrend in Maestro, a popular Telegram trading bot [78], beginning in December 2023.
@@
 Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned.
-For example, Mastro may prefer to send flows to builders directly for a better price than what MEV-Share and MEV Blocker can offer.
-It is surprising to see that upstream providers can significantly swing the significance of platforms.
-Finding V: We identified five pivotal providers (MEV-Share, MEV Blocker, jaredfromsubway.eth, Banana Gun, Maestro) who had a sustained influence on the outcome of more than half of MEV-Boost auctions.
-A positive trend towards decentralization is observed as the pivotal level of the searcher jaredfromsubway.eth decreased, while the pivotal level of MEV-Share and MEV Blocker increased.
-However, their reliance on upstream providers like Maestro is alarming.
-Case study for maestro.
-To delve deeper into the relationship between these two trends, we plot Maestro’s contribution to the total daily value of transactions sent through MEV-Share and MEV Blocker during our study in Fig.
-14.
-Surprisingly, during the time periods when MEV-Share and MEV Blocker were highly pivotal in Fig.
-8, i.e., from May to November, Maestro’s contribution accounts for approximately 80%!
-This means Maestro used to be an important upstream of MEV-Share and MEV Blocker.
-In November 2023, Maestro stopped sending transactions to MEV-Share and MEV Blocker, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
-Figure 14: Daily percentage of the value of the transactions originating from Maestro for both MEV-Share and MEV Blocker.
-Figure 15: Weekly builder-specific pivotal levels for the top-3 builders for slots with winning bid values of no less than 0.5 ETH.
-F.3 New Integration Pattern
-To illustrate the new integration pattern between providers and builders, we consider the MEV significant in a slot if its winning bid value is no less than 0.5 ETH.
-Additionally, we focus only on the provider that is the sole pivotal provider in a slot, meaning this provider provides significant MEV to the builder, enabling the builder to win the auction.
-We apply the above adjustments to builder-specific pivotal levels, and the results are shown in Fig.
-15.
-Compared to Fig.
-9, we observe that Banana Gun and Titan, as well as jaredfromsubway.eth and Beaver, follow this integration pattern: a provider selectively shares valuable order flows with specific builders when MEV is significant.
-F.4 Requirements of Providers
-Tab. 5 summarizes the providers we identify and their requirements for accessing their private order flows.
+For example, we observe that Maestro formed an integration with Beaver after ceasing to use these two platforms.
+E.2 Requirements of Providers
+Due to the lack of trustworthy fair exchange mechanisms, providers have to rely on an informal reputation system for self-protection, sending order flows only to builders who meet their requirements.
+To understand the access barrier caused by the trust crisis, we investigate different providers and their requirements for accessing order flows.
+Tab.
+V summarizes the providers we identify and their requirements for accessing their private order flows.
 Type Provider Requirements
@@
     We do not claim that the list of the private order flow providers is complete since there is no ground truth.
-Table 5: Collections of the private order flow providers and their requirements.
-F.5 Minimal Cost for 1% Market Share
+TABLE V: Collections of the private order flow providers and their requirements.
+Subsidization.
+New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share.
+In practice, builders pay out of pocket to overbid in MEV-Boost auctions, a practice known as block subsidization [70], representing an entry barrier for new builders.
+We quantify the minimal cost for a new builder to meet the MEV Blocker’s requirements (Appx.
+E.3), which reflects the costs of building a reputation.
+The minimal cost continues to increase and can be up to 18 ETH.
+This cost represents a lower bound since we do not have the corresponding numbers for other providers.
+Note that although subsidization can help meet the requirements set by channels, it cannot ensure that searchers’ trust concerns are addressed, allowing builders to receive their order flows.
+E.3 Minimal Cost for 1% Market Share
 We compute the minimal subsidy needed to establish a 1% market share for one month (which is the requirement of MEV Blocker [36]) by replaying historical auctions with an added new builder who can only access the public order flow.
@@
 The result is plotted in Fig.
-16.
-Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 7.4 ETH by July 2024.It is important to recognize that the costs derived from this method only represent the lower bound.
+14.
+Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 18 ETH by October 2024.
+It is important to recognize that the costs derived from this method only represent the lower bound.
 Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory.
 Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share.
-In reality, the subsidies new builders need to provide are substantially higher.
-For example, tbuilder [78], a new builder in the builder market, subsidized 6.8 ETH in total for 645 blocks in the week before March 21, 2024 [36].
 Figure 14: Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
-Appendix G Pseudocode of TEE Builder
-The pseudocode for the builder running in TEEs with a refund mechanism is shown in Fig. 17.
-[…]
-$orders\leftarrow[]$ // Pending transactions.
-$sources\leftarrow\{\}$ // Record the order flow providers.
-// Receive order flows from providers.
-[…]
-$orders\leftarrow orders+[order]$
-$sources[order]\leftarrow provider$
-// Build best block and refund providers.
-On building:
-$v,B\leftarrow\textsc{build}(orders)$
-$payments\leftarrow[]$
-// For each provider, refund its marginal contribution,
- For every $order$ in $B$:
-  $provider\leftarrow sources[order]$
-  $v^{\prime},\_\leftarrow\textsc{build}(orders-[order])$
-  $payment\leftarrow\textsc{transfer}(provider,v-v^{\prime})$
-  $payments\leftarrow payments+[payment]$
-// Remove all transactions included on chain.
- For every $order$ in $B$:
-  $orders\leftarrow orders-[order]$
-  Delete $sources[order]$
-$B\leftarrow B+payments$
-[…]
-Figure 17: The pseudocode of TEE builder.
+Appendix F Limitations of Our Approach and Data
+True value.
+We compute the true values of bids using a similar approach as that in MEV analysis tools [41, 42] and previous works [43, 44].
+This approach has a few limitations.
+First, it relies on the knowledge of the set of addresses controlled by each builder.
+Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses.
+We cross-validated our builder address dataset with other sources to ensure completeness (see Sec.
+3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate.
+Second, this approach may neglect off-chain profits, such as off-chain payments received from searchers.
+Off-chain payments are more likely to happen between integrated builder-searcher pairs.
+We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization.
+However, in our study of evaluating competitiveness and efficiency, we did not observe any excessively large subsidies.
+Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness.
+We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value.
+However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value.
+It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
+Third, this approach may overestimate builders’ profits because builders might refund providers for their private order flows.
+Such refunds may occur through private, off-chain deals and are therefore not captured.
+Overestimating the winner’s true value does not impact the competitiveness calculation; however, overestimating the second-highest true value could lead us to misclassify a competitive auction as uncompetitive.
+However, our study indicates that this kind of integration between top builders and providers began on a large scale starting in 2024, after the time range of our evaluation of competitiveness and efficiency.
+Missing pivotal providers.
+Our results in Sec.
+5.1 might omit some pivotal providers since the providers we can identify are a subset of all providers.
+For example, BloXroute (a widely used channel) does not publish the list of transactions they processed.
+Thus, we can’t evaluate its pivotal level.
+This limitation is common in existing studies [18, 19].
+However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec.
+3).
+Thus, the significance of providers will only be underestimated.
+Our auction dataset might miss partial bids since they are self-reported by relays.
+Recall that we used the second-highest bid value in an MEV-Boost auction as the threshold to determine a pivotal provider.
+Therefore, underestimating the second-highest bid value might also lead us to underestimate the significance of each provider.

### v5

@@
 Only the winning bids are recorded in the blockchain, but to compute the losses, we also need losing bids.
-The key ingredient that enabled our study is a large-scale auction dataset we curated since 2022: we collected 8.79 billion partial bids [^1] Partial and full bids are to be defined in Sec.
-3.1 of auctions from September 2022 to July 2024; in addition, in collaboration with ultra sound relay [15], we obtained full bids for a total of 301,479 auctions from April 2023 to December 2023.
+The key ingredient that enabled our study is a large-scale auction dataset we curated since 2022: we collected approximately 12 billion partial bids [^1] Partial and full bids are to be defined in Sec.
+3.1 of auctions from September 2022 to October 2024; in addition, in collaboration with ultra sound relay [15], we obtained full bids for a total of 301,479 auctions from April 2023 to December 2023.
 From the auction dataset, we can compute the true value of all bids.
@@
 The difference in the builder’s ability to access important private order flows can drive the inequality in block-building capacity.
-Not every provider is equally important, and some providers who generate profitable order flows can significantly impact builders.
+Not every provider is equally important, and some providers that generate profitable order flows can significantly impact builders.
 Thus, we propose a metric to identify important providers and investigate their implications.
@@
 To formally capture the intuition of “important providers”, we defined pivotal level to quantify the sustained impact of a given provider.
-A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ were not available.
+A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ not been available.
 In other words, pivotal providers are necessary for the winner to win.
@@
 They do not cause inequality loss since the competition among these builders allows proposers to capture this value.
-The reputability of a builder is accessed per their specific rules. While top builders usually pass the assessments, it is harder for new builders to enter the market.
+The reputability of a builder is assessed per their specific rules. While top builders usually pass the assessments, it is harder for new builders to enter the market.
 Exclusive pivotal providers.
@@
 Besides the well-known exclusive relationships (known as “integration” [20]), we also identified that three of the top five pivotal providers, previously not known to have integration, began integrating with top builders in 2024.
-Our finding shows that the integration of these three pivotal providers accounted for at least 45% of the total proposer losses in 2024.
+Our findings show that the integration of these three pivotal providers accounted for significant proposer losses in 2024.
 Integration poses a market entry barrier as new builders likely lack integrated providers.
@@
   • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest.
-Our dataset of 11 billion partial bids since 2022 is the largest scale known to us.
+Our dataset of approximately 12 billion partial bids since 2022 is the largest scale known to us.
+We have open-sourced the code and datasets at [24] and [25].
   • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions.
@@
   • Builders submit bids to relays throughout the slot.
-For this paper, we denote a bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key.
-Bids are signed by the builder but we omit the signature for clarity.
+For this paper, we denote a full bid with $(B,BV,PK_{\text{builder}})$ that consists of a block $B$, a bid value $BV$ to be paid to the proposed if the block is proposed, and the builder’s public key.
+Bids are signed by the builder, but we omit the signature for clarity.
 Relays verify that $B$ indeed pays the proposer at least $BV$.
@@
 We assume a builder $p$’s income from building a block $B$ is $EV_{p}(B)$, or $EV(B)$ for short since the builder identity is usually encoded in the block.
-We compute EV following a similar approach as that in MEV analysis tools [41, 42] and previous works [43, 44].
-Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees [45] and direct transfers [46] to the builder’s addresses minus the refund to other addresses.
+We compute EV following a similar approach to that in MEV analysis tools [44, 45] and previous works [46, 47].
+Specifically, $EV(B)$ can be computed as the sum of the transaction priority fees [48] and direct transfers [49] to the builder’s addresses, minus the payment to the proposer and refund to other addresses.
 We also refer to the builder’s income from a block (or a set of transactions in a block) as the value of the block (or the set of transactions).
@@
 Recall from Sec.
-2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc.
+2.1 that a bid consists of a block $B$ and a bid value $BV$, while a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$, where $metadata$ includes parent hash, gas limit, gas used, etc.
 To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below.
@@
   • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx.
-A for details).
+B for details).
     To validate the completeness of our dataset, we manually cross-check it with public datasets [55, 56] and related works [46, 41], and the results show that our dataset includes 154 more public keys, so our dataset is the most complete one so far.
@@
   • Private transaction hashes: We use data from Mempool Guru [61] to label transactions that bypassed the public mempool as private.
-We believe these labels have already been cross-validated since the private transactions are collected by seven distributed nodes.
+We believe these labels have already been cross-validated, as the private transactions were not collected by any of the seven distributed nodes.
   • Transaction sources: We are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through.
 We identified transaction sources using ground truth data (as described in Appx.
-A).
+B).
 4 Proposer Loss in the MEV-Boost Auctions
@@
 For example, suppose executing $B$ yields a net profit of $0.1$ ETH after paying $0.9$ ETH in bid value in the auction, then the true value (the revenue) is $1.0$ ETH.
+Thanks to the full bids dataset, we can compute not only the transactions that appear in on-chain blocks but also those included in other bids submitted in the auction.
 To compute $TV(B)$ of a given block $B$, we follow the standard approach in MEV analysis tools [44, 45] and previous works [46, 47], though we built our own analysis toolchain and optimized its performance to handle the large volume of historical auction data (326.6 million full bids).
@@
 Namely, $CI(s)$ measures the relative difference between the winning bid and the second-highest true value.
-$CI(s)\geq 0$ indicates that the winning bid value is not less than the second-highest true value, thus satisfying the definition of a competitive auction, while $CI(s)<0$ means there is insufficient competition.
+We define an auction as competitive if $CI(s)\geq 0$, meaning the winning bid is at least as high as the second-highest true value; if $CI(s)<0$, the auction is uncompetitive.
 Efficient index.   Same as above, the efficient index, $EI(s)$, is defined as:
@@
 Fig.
-4 shows the calculated CI and EI distribution across all MEV-Boost auctions in the our dataset.
+4 shows the calculated CI and EI distribution across all MEV-Boost auctions in our dataset.
 Notably, 91.8% (sum of the areas with $CI\geq 0$) of the auctions are competitive, yet only 75.1% (sum of the areas with $EI>0$) are efficient.
@@
 Loss from uncompetiveness.
-For an uncompetitive auction, the loss from uncompetitiveness is the difference between the second-highest true value and the actual winning bid value, representing the loss a validator incurs compared to what it would have received in a competitive auction.
+For an uncompetitive MEV-Boost auction in slot $s$, let $TV(p_{2},s)$ denote the second-highest true value and $BV_{w}(s)$ denote the winning bid.
+The loss from uncompetitiveness is defined as
+$L_{\text{uncomp.}}(s)=TV(p_{2},s)-BV_{w}(s),$
+representing the loss a proposer incurs compared to what it would have received in a competitive auction.
 We computed this loss using our dataset, and the results are shown in Tab.
@@
 TABLE II: Losses from uncompetitiveness and inequality.
-Time Slots Profits (ETH) Losses-un (ETH) (%) Losses-in (ETH) (%)
+Time Slots Profits (ETH) $L_{\text{uncomp.}}$ (ETH) (%) $L_{\text{inequal.}}$ (ETH) (%)
 April 9-15 28,385 2,704.4 46.9 (1.7) 312.1 (11.5)
@@
 However, a proposer may still incur loss due to a significant block-building capacity gap between builders.
-In MEV Boost auctions, this quantity is captured by the true values of the builder’s bids.
+In MEV-Boost auctions, this quantity is captured by the true values of the builder’s bids.
 Quantifying the inequality of block-building capacities.
@@
 Figure 6: Violin graphs for the distribution of the QCD across three MEV tiers, for top, middle, and tail builders.
-The QCD measures the inequality of block-build capacity across builders.
+The QCD measures the inequality of block-building capacity across builders.
 This figure reveals that such inequality worsens as MEV increases.
@@
 Finding II: The inequality in block building is lowest among top builders, higher among middle builders, and most pronounced among tail builders, and the inequality worsens as a slot’s MEV increases.
-Loss from inequality.
-For each auction, the loss from inequality is the difference between the highest and second-highest true values, reflecting the potential loss a validator may face due to unequal block-building capacities, as a competitive auction only ensures the winning bid slightly exceeds the second-highest true value.
+Loss from inequality.   For the auction in slot $s$, let $TV(p_{1},s)$ and $TV(p_{2},s)$ denote the highest and second-highest true values, respectively. The loss from inequality is defined as
+$L_{\text{inequal.}}(s)=TV(p_{1},s)-TV(p_{2},s),$
+reflecting the potential loss a proposer may face due to unequal block-building capacities, since a competitive auction only ensures the winning bid slightly exceeds the second-highest true value.
 As shown in Tab.
-II, the loss from inequality ranges from 5.6% to 11.5% during the periods covered by our dataset.
+II, the loss from inequality ranges from 5.6% to 11.5% over the periods in our dataset.
 Comparing the two types of loss, we find that inequality is the primary source of proposers’ loss.
@@
 A significant estimated proposer loss suggests that proposers incur substantial losses.
+Figure 7: Daily estimated proposer loss over two years. Proposers suffered significant losses in 2024.
 We calculate the daily estimated proposer loss by aggregating the estimated losses from all MEV-Boost auctions each day and comparing this total to the overall block value for that day.
@@
 The daily estimated proposer loss is significant on August 6, 2024, primarily because, in an auction on that date (slot 9676257), the winner’s bid is 11 ETH while the true value is 803 ETH—resulting in a significant loss for the proposer.
-Figure 7: Daily estimated proposer loss over two years. Proposers suffered significant losses in 2024.
 Why are proposer losses more significant in 2024?
@@
 This enables us to analyze their statements regarding how their order flows are distributed to different builders—both claim that every connected builder receives the order flows [38, 39].
-Although their order flows are not accessible to all builders (as summarized in Tab.
-V), top builders such as Beaver, Titan, and Rsync have access to them.
+Although the order flows are not accessible to all builders (as summarized in Tab.
+V, presented in Sec.
+E.2), top builders such as Beaver, Titan, and Rsync have access to them.
 Ideally, competition among these top builders will ensure that the value from these order flows is ultimately captured by the proposers.
 Notably, while these order flows are distributed equally to all connected builders, inequality still exists.
-Order flow providers may exclude less reputable builders due to trust concerns.
+Order flow providers may exclude less reputable builders due to trust concerns because builders are at an advantage and malicious builders can harm providers by, e.g., sandwich attacks [62] and imitation attacks [63].
 Although this exclusion may not harm proposers, it could exacerbate builder centralization, as new builders must expend significant effort to meet the requirements imposed by these pivotal providers.
@@
 To confirm this conjecture and evaluate the extent to which proposer loss is related to integration, we need to investigate exclusive order flows from providers to builders.
+before 2024 2024
+Provider MEV Activity Builder Blocks (#) Losses (ETH) (%) Provider MEV Activity Builder Blocks (#) Losses (ETH) (%)
+SCP CEX-DEX arbitrage Beaver 42,639 4,566.0  (22.7) Banana Gun Sniper Titan 34,442 17,488.1  (32.6)
+Searcher: 0x57c1… CEX-DEX arbitrage Addr: 0x4737… 12,220 1,088.7  (5.4) SCP CEX-DEX arbitrage Beaver 13,148 3,077.8  (5.7)
+Wintermute CEX-DEX arbitrage Rsync 13,823 642.6  (3.2) Searcher: 0x6980… Atomic arbitrage I can haz block? 516 2,495.7  (4.7)
+Searcher: 0xf8b7… CEX-DEX arbitrage Manta 9,699 473.9  (2.4) jaredfromsubway.eth Sandwich Beaver 22,230 2,480.5  (4.6)
+Searcher: 0x9dd8… Atomic arbitrage Addr: 0x3bee… 1,826 424.9  (2.1) Searcher: 0x6454… Liquidation Addr: 0x3bee… 809 1,231.8  (2.3)
+Searcher: 0x761d… Atomic arbitrage builder0x69 19 287.5  (1.4) WINTERMUTE CEX-DEX arbitrage Rsync 8,586 1,038.3  (1.9)
+Searcher: 0x6454… Liquidation Addr: 0x3bee… 456 246.6  (1.2) c0ffeebabe.eth Liquidation Ty For The Block 368 1,003.6  (1.9)
+Searcher: 0x5e51… Atomic arbitrage eth-builder 370 226.5  (1.1) Banana Gun Sniper Beaver 2,424 920.5  (1.7)
+Searcher: 0xeeaa… Atomic arbitrage eth-builder 67 195.6  (1.0) Maestro Sniper Beaver 10,587 439.7  (0.8)
+Searcher: 0xe20c… Atomic arbitrage Addr: 0x5638… 592 168.9  (0.8) Searcher: 0x77d5… Atomic arbitrage Rsync 797 333.3  (0.6)
+TABLE III: The top 10 providers sending exclusive order flows and their corresponding builders, contribute to the highest estimated proposer loss (before and in 2024).
 Challenges of identifying exclusive order flows.
@@
 Therefore, attributing all proposer losses in a builder’s blocks to its integration might be inaccurate.
-before 2024 2024
-Provider MEV Activity Builder Blocks (#) Losses (ETH) (%) Provider MEV Activity Builder Blocks (#) Losses (ETH) (%)
-SCP CEX-DEX arbitrage Beaver 18,509 4,149.3  (21.0) Banana Gun Sniper Titan 33,748 19,401.3  (37.1)
-Searcher: 0x57c1… CEX-DEX arbitrage Addr: 0x4737… 7,166 1,037.3  (5.2) SCP CEX-DEX arbitrage Beaver 10,087 3,249.8  (6.2)
-Wintermute CEX-DEX arbitrage Rsync 6,555 553.1  (2.8) jaredfromsubway.eth Sandwich Beaver 20,213 2,575.4  (5.0)
-Searcher: 0xf8b7… CEX-DEX arbitrage Manta 4,606 428.2  (2.2) Searcher: 0x6980… Atomic arbitrage I can haz block? 513 2,484.9  (4.8)
-Searcher: 0x9dd8… Atomic arbitrage Addr: 0x3bee… 1,600 411.5  (2.6) Searcher: 0x6454… Liquidation Addr: 0x3bee… 793 1,229.6  (2.4)
-Searcher: 0x761d… Atomic arbitrage builder0x69 12 287.0  (1.5) c0ffeebabe.eth Liquidation Ty For The Block 367 1,003.6  (1.9)
-Searcher: 0x6454… Liquidation Addr: 0x3bee… 539 262.5  (1.5) Banana Gun Sniper Beaver 2,430 987.4  (1.9)
-Searcher: 0x5e51… Atomic arbitrage eth-builder 266 223.2  (1.1) WINTERMUTE CEX-DEX arbitrage Rsync 6,503 976.8  (1.9)
-Searcher: 0xeeaa… Atomic arbitrage eth-builder 51 194.1  (1.0) Maestro Sniper Beaver 10,103 520.3  (1.0)
-Searcher: 0xe20c… Atomic arbitrage Addr: 0x5638… 579 168.5  (0.9) Searcher: 0x77d5… Atomic arbitrage Rsync 782 345.3  (0.7)
-TABLE III: The top 10 providers sending exclusive order flows and their corresponding builders, contribute to the highest estimated proposer loss (before and in 2024).
 Our approach.
@@
 This also confirms our conjecture.
-For instance, Banana Gun exclusively sends order flows to Titan in approximately 33K blocks, accounting for 37.1% of the estimated proposer losses covered in our study.
-Additionally, two other top-5 pivotal providers, jaredfromsubway.eth and Maestro, also exclusively send order flows to Beaver, contributing to 6% of the losses.
+For instance, Banana Gun exclusively sends order flows to Titan in approximately 34K blocks, accounting for 32.6% of the estimated proposer losses covered in our study.
+Additionally, two other top-5 pivotal providers, jaredfromsubway.eth and Maestro, also exclusively send order flows to Beaver, contributing to 5.4% of the losses.
 In contrast, the total losses caused by the integrated CEX-DEX arbitrage searchers remain relatively similar before and during 2024.
@@
 A builder with large off-chain payments could bid higher than the computed true value, appearing as an extremely large subsidy.
+This could lead to the misclassification of an efficient auction as inefficient due to subsidization.
 However, we observe that there are no excessively large subsidies, and most builders bid close to our computed true values in our study of competitiveness and inequality, suggesting that off-chain payments are not significant.
+While underestimating the winner’s true value does not affect our competitiveness assessment, underestimating the second-highest true value could cause an uncompetitive auction to be misclassified as competitive.
+Nevertheless, this scenario is rare.
+In practice, builders with higher true values would have an incentive to bid higher to win, making such misclassifications unlikely.
 The potential refunds to providers do not affect the computation of proposers’ losses because this portion is not allocated to proposers, regardless of the volume of refunds.
-Due to page constraints, we discuss the potential implications of the limitations in Appx.
-F.
+However, overestimating the second-highest true value could lead us to misclassify a competitive auction as uncompetitive.
+Our study shows that large-scale integration between top builders and providers—where such refunds are more likely—began in 2024, after the period covered by our evaluation of competitiveness and efficiency.
+We might also miss pivotal providers due to incomplete coverage of transaction sources.
+For example, BloXroute (a widely used channel) does not publish the list of transactions they processed.
+As a result, some providers may be missing from our analysis.
+However, there will be no false positives because each provider attribution is either directly from official sources or through known searcher addresses (see Sec.
+3).
+This limitation leads to an underestimation of the significance of providers, but not an overestimation.
+Similarly, because partial bids are self-reported by relays, underestimating the second-highest bid could cause us to underestimate a provider’s pivotal level.
 7 Related Work
@@
 This work is partially supported by an Ethereum Foundation Grant.
-We thank ultra sound relay for sharing the full bid dataset.
-We also thank the teams of BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions about access requirements.
-We are grateful for the insightful comments by Prof.
+We thank our shepherd and the reviewers for their helpful comments, which significantly improved the paper.
+We are grateful to ultra sound relay for sharing the full bid dataset.
+We also thank the teams at BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions regarding access requirements.
+Finally, we appreciate the insightful comments from Prof.
 Campbell Harvey and the valuable discussions with Thomas Thiery, Barnabé Monnot, Danning Sui, Burak Öz, Data Always, and Georgios Konstantopoulos.
@@
 Appendix E Private Order Flows
+Type Provider Requirements
+Searcher Individual searchers Requirements are unclear.
+Telegram trading bot Banana Gun [84], Maestro [83], Unibot [85] The requirements are the same as the channels they use or unclear.
+Channel BloXroute [40] Builders are granted access to different tiers, each providing varying privileges and permissions, based on an assessment.
+Merkle [88] Builders must have a proven track record of accurately and honestly computing bundles, or they should consult with the Merkle team for consideration.
+MeowRPC [89] The transactions are sent to the top builders which MeowRPC’s team trusts.
+MEV Blocker [39] Builders must maintain at least 1% market share each week of the previous month.
+MEV-Share [38] Builders must demonstrate a competitive market share for consideration.
+Transaction Boost [90] Consideration for partnership is based on the Blocknative team’s evaluation of the builder’s reputation.
+  • $\star$
+    We do not claim that the list of the private order flow providers is complete since there is no ground truth.
+TABLE V: Collections of the private order flow providers and their requirements.
 E.1 Pivotal Private Order Flow Providers
@@
 V summarizes the providers we identify and their requirements for accessing their private order flows.
-Type Provider Requirements
-Searcher Individual searchers Requirements are unclear.
-Telegram trading bot Banana Gun [79], Maestro [78], Unibot [80] The requirements are the same as the channels they use or unclear.
-Channel BloXroute [37] Builders are granted access to different tiers, each providing varying privileges and permissions, based on an assessment.
-Merkle [83] Builders must have a proven track record of accurately and honestly computing bundles, or they should consult with the Merkle team for consideration.
-MeowRPC [84] The transactions are sent to the top builders which MeowRPC’s team trusts.
-MEV Blocker [36] Builders must maintain at least 1% market share each week of the previous month.
-MEV-Share [35] Builders must demonstrate a competitive market share for consideration.
-Transaction Boost [85] Consideration for partnership is based on the Blocknative team’s evaluation of the builder’s reputation.
-  • $\star$
-    We do not claim that the list of the private order flow providers is complete since there is no ground truth.
-TABLE V: Collections of the private order flow providers and their requirements.
 Subsidization.
@@
 E.3 Minimal Cost for 1% Market Share
+Figure 14: Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
 We compute the minimal subsidy needed to establish a 1% market share for one month (which is the requirement of MEV Blocker [39]) by replaying historical auctions with an added new builder who can only access the public order flow.
@@
 Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share.
-Figure 14: Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
-Appendix F Limitations of Our Approach and Data
-True value.
-We compute the true values of bids using a similar approach as that in MEV analysis tools [41, 42] and previous works [43, 44].
-This approach has a few limitations.
-First, it relies on the knowledge of the set of addresses controlled by each builder.
-Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses.
-We cross-validated our builder address dataset with other sources to ensure completeness (see Sec.
-3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate.
-Second, this approach may neglect off-chain profits, such as off-chain payments received from searchers.
-Off-chain payments are more likely to happen between integrated builder-searcher pairs.
-We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization.
-However, in our study of evaluating competitiveness and efficiency, we did not observe any excessively large subsidies.
-Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness.
-We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value.
-However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value.
-It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
-Third, this approach may overestimate builders’ profits because builders might refund providers for their private order flows.
-Such refunds may occur through private, off-chain deals and are therefore not captured.
-Overestimating the winner’s true value does not impact the competitiveness calculation; however, overestimating the second-highest true value could lead us to misclassify a competitive auction as uncompetitive.
-However, our study indicates that this kind of integration between top builders and providers began on a large scale starting in 2024, after the time range of our evaluation of competitiveness and efficiency.
-Missing pivotal providers.
-Our results in Sec.
-5.1 might omit some pivotal providers since the providers we can identify are a subset of all providers.
-For example, BloXroute (a widely used channel) does not publish the list of transactions they processed.
-Thus, we can’t evaluate its pivotal level.
-This limitation is common in existing studies [18, 19].
-However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec.
-3).
-Thus, the significance of providers will only be underestimated.
-Our auction dataset might miss partial bids since they are self-reported by relays.
-Recall that we used the second-highest bid value in an MEV-Boost auction as the threshold to determine a pivotal provider.
-Therefore, underestimating the second-highest bid value might also lead us to underestimate the significance of each provider.